### v2

# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

Abstract
Autonomous mobile robots must maintain safety, but should not sacrifice performance, leading to the classical reach-avoid problem. This paper seeks to compute trajectory plans for which a robot is guaranteed to reach a goal and to avoid obstacles in the specific near danger case that the obstacles and goal are near each other. The proposed method builds off of a common approach of using a simplified planning model to generate plans, which are then tracked using a high-fidelity tracking model and controller. Existing safe planning approaches use reachability analysis to overapproximate the error between these models, but this introduces additional numerical approximation error and thereby conservativeness that prevents goal reaching. The present work instead proposes a Piecewise Affine Reach-avoid Computation (PARC) method to tightly approximate the reachable set of the planning model. With PARC, the main source of conservativeness is the model mismatch, which can be mitigated by careful controller and planning model design. The utility of this method is demonstrated through extensive numerical experiments in which PARC outperforms state-of-the-art reach avoid methods in near-danger goal reaching. Furthermore, in a simulated demonstration, PARC enables generation of provably-safe extreme vehicle dynamics drift parking maneuvers.
I Introduction
The increasing deployment of autonomous robots in daily tasks necessitates addressing complex operational challenges. Many of these tasks can be conceptualized as reach-avoid problems, wherein a robot must steer its states into a set of goal states without violating safety-critical constraints [1]. Achieving both liveness – the assurance of goal fulfillment [2] – and adherence to safety constraints is pivotal in these scenarios. In this paper, we are particularly interested in goal reaching near danger, meaning that an obstacle is near enough to a desired goal location to prevent typical robust planning or control approaches from reaching the goal. This often occurs when a goal and obstacles are closer to each other than one width of a robot’s body, such as for a quadrotor drone flying through a narrow gap or an autonomous car parallel parking.
The reach-avoid problem is typically solved with one of two approaches, both based on reachability analysis or computing reachable sets of a system’s trajectories. Note that an extended literature review is presented in Appendix -A.
Figure 1: Our overall goal is to find safe goal-reaching motion plans near danger. In (a), we show the components of our proposed Piecewise Affine Reach-avoid Computation (PARC) method, with the paper section number on the top left, and the corresponding symbol on the bottom. This method enables a new state-of-the-art for safe reach-avoid problems, as evidenced by (b), which shows examples of extreme vehicle maneuvers. These planned drift parking trajectories are sampled from the Backward Reach-Avoid Set (BRAS, green triangle on the top left) using PARC. Our vehicle is in a blue time lapse. The two red cars are obstacles. The goal set for the center of mass is the green rectangle. The small size of the BRAS indicates the challenge of finding safe trajectories.
The first approach is control intervention, wherein a motion planner proposes actions, while a safety filter [3, 4] prevents the robot from executing actions for which its reachable set would intersect obstacles. This approach often requires compensating for the worst-case behavior of the motion planner [5, 6], which we find leads to conservativeness (i.e., sacrificing liveness) in our experiments (Section V). The second approach is a planner-tracker framework, wherein a motion planner proposes plans that are safe by leveraging the knowledge of a low-level controller’s tracking ability. This controller can be synthesized [7, 8] or hand-engineered [9, 10]; the key difference is that the planner need not be treated as adversarial, which reduces conservativeness. However, this approach typically limits the space of possible plans due to the need to compute the robot’s tracking ability. Furthermore, one must compensate for tracking error between the simplified planning model and the high-fidelity tracking model (i.e., the robot cannot perfectly track plans).
Both control intervention and planner-tracker approaches suffer from challenges in numerical representation due to the need to underapproximate the set of states that reach the target set and overapproximate those that avoids obstacles when guaranteeing safety and liveness in implementation. Implementations such as sums-of-squares polynomials [7, 9, 11], grid-based PDE solutions [5], or polytope methods [10, 12] all introduce additional conservativeness purely due to compensating for numerical representation error. This challenge is exacerbated for systems with high-dimensional, nonlinear dynamics or observations; machine learning can address some of these concerns [13, 14, 15, 16, 17], at the expense of losing guarantees.
In this work, we leverage two key insights to perform the goal-reaching near danger. First, as mentioned before, planner-tracker methods can reduce conservativeness by treating the planner and tracking controller as cooperative, as opposed to adversarial. Second, by representing our motion planner as a piecewise affine (PWA) system [18], we can tightly approximate its reachable set, with conservativeness mainly due to tracking controller and planning model design. We build off of a long history of discrete-time PWA system analysis in optimal control (e.g., [19, 20, 21]), which we extend to the continuous-time reach-avoid problem for robot motion planning. We demonstrate that this unlocks a faster and less conservative approach to a variety of reach-avoid problems from the robotics literature, and provides the solution to a reachable set computation for drift parking vehicle dynamics. An overview and example of our approach is shown in Fig. 1.
Contributions:
  1. 1.
    We propose an efficient, parallelizable method for Piecewise Affine Reach-avoid Computation (PARC) to underapproximate the Backwards Reach-Avoid Set (BRAS) of a planning model represented as a low-dimensional discrete-time, time-variant PWA system. Our method extends to discrete-time, time-variant affine systems as a particular case.
  2. 2.
    We propose a scheme to incorporate tracking error into PARC to underapproximate the reach set and overapproximate the avoid set for a trajectory planning model. This allows us to represent the BRAS that compactly represents all trajectory plans that reach the goal and avoid obstacles when tracked by the robot, represented by a high-fidelity dynamical model.
  3. 3.
    We stress-test PARC on examples drawn from the safe control and planning literature: a 4-D TurtleBot, a 10-D near-hover quadrotor, and a 13-D quadrotor. We show that PARC outperforms state-of-the-art methods such as FaSTrack [5], Neural CLBF [15], and RTD [9, 10]. We also demonstrate PARC on a simulated 6-D vehicle model to plan safe, extreme drift parallel parking maneuvers.
Paper Organization: Next, we introduce mathematical objects used throughout our formulation (Section II). We then present our problem formulation (Section III) and proposed approach (Section IV). We assess PARC with numerical experiments (Section V), then demonstrate a drift parking extreme vehicle dynamics example in simulation, extending beyond the state of the art in reachability-based safe motion-planning and control methods (Section VI). We discuss takeaways and limitations in Section VII. Note, we have extensive appendices, and call out key results in the text.
II Preliminaries
We now introduce our notation conventions and define H-polytopes and PWA functions, which we use to represent sets and dynamics of planning model throughout the paper. We then present a one-step Backward Reachable Set (BRS) computation that is the core of our proposed PARC method.
II-A Notation
The real numbers are $\mathbb{R}$ and the natural numbers are $\mathbb{N}$. Scalars are lowercase italic and sets are in uppercase italic (e.g., $a\in A\subset\mathbb{R}^{n}$). Vectors are lowercase bold, and matrices are uppercase bold (e.g., $\mathbf{y}=\mathbf{A}\mathbf{x}$). An $n\times m$ matrix of zeros is $\mathbf{0}_{n\times m}$, and an $n\times n$ identity matrix is $\mathbf{I}_{n}$. The $p^{\mathrm{\textnormal{th}}}$ to $q^{\mathrm{\textnormal{th}}}$ dimensions of vector $\mathbf{x}$ and the submatrix of $\mathbf{A}$ defined by row dimensions $p_{1}$:$q_{1}$ and column dimensions $p_{2}$:$q_{2}$ are denoted as $(\mathbf{x})_{p:q}$ and $(\mathbf{A})_{p_{1}:q_{1},p_{2}:q_{2}}$, respectively.
II-B H-Polytopes
In this work, we represent all sets as H-polytopes or unions of H-polytopes, which have extensive algorithm and toolbox support [22, 23]. An ${n}$-dimensional H-polytope $\mathcal{P}(\mathbf{A},\mathbf{b})\subset\mathbb{R}^{{n}}$ is a convex set parameterized by ${n}_{\mathrm{\textnormal{h}}}$ linear constraints $\mathbf{A}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}}$ and $\mathbf{b}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}}$:
$\mathcal{P}(\mathbf{A},\mathbf{b})$ =$=\left\{\mathbf{x}\ |\ \mathbf{A}\mathbf{x}\leq\mathbf{b}\right\}.$ (1)
Unless otherwise stated, we assume all H-polytopes are compact.
Consider a pair of H-polytopes $P_{1}=\mathcal{P}(\mathbf{A}_{1},\mathbf{b}_{1})$, $P_{2}=\mathcal{P}(\mathbf{A}_{2},\mathbf{b}_{2})$. We make use of intersections $P_{1}\cap P_{2}$, Minkowski sums ($P_{1}\oplus P_{2}=\left\{\mathbf{x}+\mathbf{y}\ |\ \mathbf{x}\in P_{1},\mathbf{y}\in P_{2}\right\}$), Pontryagin differences ($P_{1}\ominus P_{2}=\left\{\mathbf{x}\in P_{1}\ |\ \mathbf{x}+\mathbf{y}\in P_{1}\ \forall\ \mathbf{y}\in P_{2}\right\}$), set differences $P_{1}\setminus P_{2}$, Cartesian products $P_{1}\times P_{2}$, and convex hulls $\mathrm{\textnormal{conv}}\!\left(P_{1},P_{2}\right)$. We also project an $n$-dimensional polytope $P$ into dimensions $p:q$ as $\mathrm{\textnormal{proj}}_{{p}:{q}}\!\left(P\right)$. These operations are detailed in Appendix -B.
II-C PWA Systems
We consider discrete-time, time-variant PWA functions to enable our reach-avoid analysis. We denote a PWA function as $\psi:X_{0}\to X_{t_{\mathrm{\textnormal{f}}}}\subseteq\mathbb{R}^{{n}}$ with output $\mathbf{x}(t_{\mathrm{\textnormal{f}}})=\mathbf{\psi}\!\left(\mathbf{x}(0)\right)$ given an input $\mathbf{x}(0)=\mathbf{x}_{0}\in X_{0}\subset\mathbb{R}^{{n}}$. The system has state $\mathbf{x}(t)\in X_{t}\subset\mathbb{R}^{{n}}$ at time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which is the solution of difference equation of dynamics:
$\mathbf{x}(t+\Delta t)=\mathbf{C}_{\ell,t}\mathbf{x}(t)+\mathbf{d}_{\ell,t},$ (2)
where $\mathbf{C}_{\ell,t}\in\mathbb{R}^{{n}\times{n}}$ and $\mathbf{d}_{\ell,t}\in\mathbb{R}^{{n}}$ define the dynamics of the $\ell^{\mathrm{\textnormal{th}}}$ PWA region. The index $\ell$ is given by
$\ell$ $=\min\{i\ |\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\},$ (3)
where $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$ is the $i^{\mathrm{\textnormal{th}}}$ of the ${n}_{\mathrm{\textnormal{PWA}},t}$ PWA regions at time $t$. We say that $\mathbf{x}(t)$ has the mode $s_{t}=\ell$ at time $t$.
In addition, if the mode of $\mathbf{x}(0),\mathbf{x}(\Delta t),\cdots,\mathbf{x}(t_{\mathrm{\textnormal{f}}}-\Delta t)$ is $s_{0},s_{\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t}$ for some input $\mathbf{x}(0)\in X_{0}$, we denote the mode sequence for $\mathbf{x}(0)$ as:
$S=(s_{0},s_{\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t}).$ (4)
For all $\mathbf{x}(0)$ that share the same mode sequence, the dynamics reduce from time-variant PWA to time-variant affine up to time $t_{\mathrm{\textnormal{f}}}$. This is the core idea that enables the efficient underapproximating BRS computation of PARC. Note, taking advantage of fixed mode sequences is well-established in robotics and controls (e.g., [19, 24, 20, 25]), but we believe we are the first to show its utility in safe reach-avoid computation of robot planning models.
We require that the interior of each PWA region does not intersect with other regions, such that any two regions are either disjoint or intersect only at their boundaries:
$\left\{\mathbf{x}\ |\ \mathbf{A}_{i}(t)\mathbf{x}<\mathbf{b}_{i}(t)\right\}\cap\mathcal{P}(\mathbf{A}_{j}(t),\mathbf{b}_{j}(t))=\emptyset,$ (5)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and $i,j\in\{1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\}$, $i\neq j$. Thus, the $\min$ in (3) dictates the dynamics’ behavior at the intersecting subsets of boundaries.
Note that, for $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define:
$X_{t}=\bigcup_{i=1}^{{n}_{\mathrm{\textnormal{PWA}},t}}\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),$ (6)
such that the input to $\psi$ is restricted to $\{\mathbf{x}_{0}\ |\ \mathbf{x}_{t}\in X_{t},t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$. Thus, a PWA function $\psi$ can be fully defined and parameterized by some given $\mathbf{C}_{i,t}$, $\mathbf{d}_{i,t}$, $\mathbf{A}_{i,t}$, and $\mathbf{b}_{i,t}$ for each $i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}$ and $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
When ${n}_{\mathrm{\textnormal{PWA}},t}=1$ for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, one obtains a discrete-time, time-variant affine system as a particular case of this problem formulation.
Finally, to ensure safety for realistic systems, we must extend our system definition from discrete to continuous time. We do so with linear interpolation: at each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, $\mathbf{x}(t^{\prime})$ for $t<t^{\prime}<t+\Delta t$ is defined by
$\mathbf{x}(t^{\prime})=\mathbf{x}(t)+(\mathbf{x}(t+\Delta t)-\mathbf{x}(t))(\tfrac{t^{\prime}-t}{\Delta t}).$ (7)
II-D One Step BRS via Inverse Affine Map
Finally, we take advantage of the H-polytope analytic solution for inverse affine maps to enable BRS computation using the following well-known lemma (c.f., [19, 26, 22]).
Lemma 1.
Consider an affine system $\mathbf{x}(t+\Delta t)=\mathbf{C}\mathbf{x}(t)+\mathbf{d}$. Define the one-step BRS as the map $\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)=\left\{\mathbf{x}|\ \mathbf{A}\mathbf{x}_{next}\leq\mathbf{b},\ \mathbf{x}_{next}=\mathbf{C}\mathbf{x}+\mathbf{d}\right\}$. This one-step BRS is in fact an H-polytope:
$\begin{split}\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)=\mathcal{P}(\mathbf{A}\mathbf{C},\mathbf{b}-\mathbf{A}\mathbf{d}).\end{split}$ (8)
Proof.
$\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)$ =$=\left\{\mathbf{x}|\ \mathbf{A}(\mathbf{C}\mathbf{x}+\mathbf{d})\leq\mathbf{b}\right\},$ (9)
$=\mathcal{P}(\mathbf{A}\mathbf{C},\mathbf{b}-\mathbf{A}\mathbf{d}).$ (10)
We take advantage of the fact that repeated applications of Lemma 1 to an H-polytope still produce an H-polytope. Notice that (8) does not require invertibility of $\mathbf{C}$, conferring an advantage over other set polytope representations such as constrained zonotopes [27]. This makes H-polytopes an excellent choice for representing the BRS.
III Problem Statement
Our overall goal is to design a reach-avoid trajectory using a simple planning model, which enables computational efficiency while also accounting for the tracking error when the plan is tracked by a robot, represented by a tracking model. We now describe each of these models and the problem we seek to solve.
III-A Planning Model
A planning model provides a low-fidelity but fast-to-compute model of the robot’s dynamics. In this work, we use parameterized reference trajectories, which can allow quick computation and strong safety guarantees [9, 10, 28].
Before defining the model, we define a workspace state $\mathbf{w}\in W\subset\mathbb{R}^{{n}_{w}}$ (i.e., $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$) that represents the position of the robot. The planning state which we denote as $\mathbf{p}\in P\subset\mathbb{R}^{{n}_{p}}$ is chosen to include the workspace state, such that $\mathbf{p}=[\mathbf{w}^{\intercal},\mathbf{p}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$ where $\mathbf{p}_{\mathrm{\textnormal{other}}}\in P_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{{n}_{p}}_{\mathrm{\textnormal{other}}}}$ represents states such as heading, velocity, etc.
We then define the planning model as
$\dot{\mathbf{p}}(t)=\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}),\quad\mathbf{p}(0)=\mathbf{p}_{0}$ (11)
where $\mathbf{p}_{0}$in $P\subset\mathbb{R}^{{n}_{p}}$ defines the initial planning state, $\mathbf{k}\in K\subset\mathbb{R}^{{n}_{k}}$ represents a trajectory parameter, and $t\in[0,t_{\mathrm{\textnormal{f}}}]$ is the time horizon. A solution $\mathbf{p}:[0,t_{\mathrm{\textnormal{f}}}]\rightarrow P$ is referred to as a plan, or equivalently, reference trajectory. For a given initial planning state and trajectory parameter, the plan at time $t$ is $\mathbf{p}(t;\mathbf{p}_{0},\mathbf{k})$. We restrict the class of planning models to enable computing reachable sets for parameterized planning models:
Assumption 2 (Extended Translation Invariance (ETI) in Workspace).
Changing the workspace state either does not affect the planning dynamics $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ in the workspace or else the dynamics are constant:
$\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\mathbf{f}_{\mathbf{w}}\right\|_{2}=0\quad\mathrm{\textnormal{and}}\quad\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{p}}\right\|_{\mathrm{\textnormal{F}}}=0,$ (12)
where $\mathbf{f}_{\mathbf{w}}=({\mathbf{f}_{\mathrm{\textnormal{plan}}}})_{1:{n}_{w}}$ and $\left\|\cdot\right\|_{\mathrm{\textnormal{F}}}$ indicates the Frobenius norm.
This is not overly restrictive; a variety of ETI planning models are shown in Section V. The need for the dynamics to be constant in some states is because we treat $\mathbf{k}$ as states with 00 dynamics to formulate our reachability method.
III-B Tracking Model
The tracking model is a high-fidelity model of robot dynamics, denoted
$\dot{\mathbf{z}}(t)=\mathbf{f}_{\mathrm{\textnormal{track}}}({\mathbf{z}(t)},{\mathbf{u}_{\mathrm{\textnormal{fb}}}(t)}),~{}~{}~{}\mathbf{z}(0)=\mathbf{z}_{0},$ (13)
where $\mathbf{z}_{0}\in Z\subset\mathbb{R}^{{n}_{z}}$ is the initial state. The tracking state $\mathbf{z}$ includes the planning state $\mathbf{p}$, i.e., $\mathbf{z}=[\mathbf{p}^{\intercal},\mathbf{z}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$. Given a plan $\mathbf{p}$ parameterized by $\mathbf{k}$, we use a feedback tracking controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}(t,\mathbf{z}(t);\mathbf{k})\in U$ (e.g., PID or MPC) to control the tracking model. We assume $Z$ and $U$ are compact and the dynamics are Lipschitz such that trajectories exist [29]. For a particular $\mathbf{z}_{0}$, we denote the resulting, unique realized trajectory (i.e., solution to (13)) as $\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})$ at time $t$.
Remark 3.
Designing the feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ is not the focus of this work. Conveniently, specifying a parameterized (i.e., restricted) space of plans $\mathbf{p}$ simplifies controller design. Our overall goal is to compensate for the fact that the controller is imperfect.
III-C Reach-Avoid Problem
We now define our reach-avoid problem: find all plans for which the tracking model safely reaches a goal set at time $t_{\mathrm{\textnormal{f}}}$. We denote the goal $\overline{P_{\mathrm{\textnormal{G}}}}\subset W$ and obstacles $\overline{\mathcal{O}}\subset W$; the overline indicates these are restricted to the workspace, which we modify later to enable the proposed method.
Note, we only consider static obstacles, because our goal is to plan a robot’s motion near danger. The danger resulting from moving or adversarial obstacles depends upon accurately predicting obstacle motion, which introduces an independent variable unrelated to our purpose. We leave extension to dynamic scenes for future work; we anticipate we can adapt reachability-based methods such as [30].
Problem 4 (Backward Reach-Avoid Set (BRAS)).
Given the the planning model (11), tracking model (13), goal set $\overline{P_{\mathrm{\textnormal{G}}}}$, obstacles $\overline{\mathcal{O}}$, and a fixed time $t_{\mathrm{\textnormal{f}}}$, find the set of planning models states and trajectory parameters for which the robot reaches $\overline{P_{\mathrm{\textnormal{G}}}}$ exactly at time $t_{\mathrm{\textnormal{f}}}$ while avoiding collisions with obstacles $\overline{\mathcal{O}}$ :
$\begin{split}\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})=\{&(\mathbf{p}_{0},\mathbf{k})\in P\times K\ \mid\forall~{}\mathbf{z}_{0}~{}\mathrm{\textnormal{s.t.}}\\ &(\mathbf{z}_{0})_{1:{n}_{w}}=\mathbf{p}_{0},\\ &(\mathbf{z}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\in\overline{P_{\mathrm{\textnormal{G}}}},~{}\mathrm{\textnormal{and}}\\ &(\mathbf{z}(\tau;\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\notin\overline{\mathcal{O}}\quad\forall\tau\in[0,t_{\mathrm{\textnormal{f}}}]\}\end{split}$ (14)
Remark 5.
A fixed time horizon is reasonable because we seek to compute a trajectory near danger; the robot is already near its goal, but reaching the goal is difficult. The BRAS is a large set far from obstacles that can be used as a target set for other methods to reach safely.
Remark 6.
We restrict the goal and obstacles to the workspace only to simplify exposition; we demonstrate more general goals and obstacles, such as in the robot’s orientation, in Sections V and VI.
III-D Example System: TurtleBot
We use the TurtleBot unicycle model as an illustrative example from the safe planning and control literature [5, 9]. Our planning model is a Dubins car:
$\dot{\mathbf{p}}=\frac{d}{dt}\begin{bmatrix}p_{x}\\ p_{y}\\ \theta\end{bmatrix}=\begin{bmatrix}v_{\mathrm{\textnormal{des}}}\cos\theta\\ v_{\mathrm{\textnormal{des}}}\sin\theta\\ \omega_{\mathrm{\textnormal{des}}}\end{bmatrix},$ (21)
with workspace (position) states $\mathbf{w}=[p_{x},p_{y}]^{\intercal}$, other (heading) state $\mathbf{p}_{\mathrm{\textnormal{other}}}=\theta$, and trajectory parameters $\mathbf{k}=[v_{\mathrm{\textnormal{des}}},\omega_{\mathrm{\textnormal{des}}}]^{\intercal}$. Note that a Dubins car model (21) is ETI in the workspace (Assum. 2) since its dynamics in workspace dimension is translation-invariant (i.e. $\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\right\|_{2}=0$.)
We use the following tracking model:
$\dot{\mathbf{z}}=\frac{d}{dt}\begin{bmatrix}p_{x}\\ p_{y}\\ \theta\\ v\end{bmatrix}=\begin{bmatrix}v\cos\theta\\ v\sin\theta\\ u_{\omega}\\ u_{v}\end{bmatrix},$ (30)
where $\mathbf{z}_{\mathrm{\textnormal{other}}}=[\theta,v]^{\intercal}$ are heading and linear velocity, and $\mathbf{u}_{\mathrm{\textnormal{fb}}}=[u_{\omega},u_{v}]^{\intercal}$ is implemented as an iterative LQR tracking controller [31] that commands angular velocity and linear acceleration.
We define a goal set as a unit box, $\overline{P_{\mathrm{\textnormal{G}}}}=\{\mathbf{w}\mid p_{x},p_{y}\in[-1,1]\}$, and an obstacle near the goal, $\overline{\mathcal{O}}=\{\mathbf{w}\mid p_{x}\in[-2.5,-1.5],p_{y}\in[-1,0]\}$ (see Figures 2 and 3).
IV Proposed Method
In this section, we detail our approach to solving the BRAS problem. First, we convert the planning model to a PWA system to enable our Piecewise Affine Reach-Avoid Computation (PARC). We then adapt the method in [19] to underapproximate the reach set of the PWA planning model without tracking error, to simplify exposition. Next, we propose a novel approach to overapproximate the continuous time avoid set of the PWA system without tracking error (Key result: Theorem 11). To enable incorporating tracking error between the planned and realized trajectories, we adapt the methods in [10, 9] to estimate time-varying tracking error. Finally, we propose PARC to incorporate tracking error and compute the BRAS (Key result: Theorem 17), enabling provably-safe goal-reaching.
IV-A Reformulation to Enable PARC
IV-A1 Augmented State, Goal, and Obstacles
Before proceeding, to enable PARC, we define an augmented planning state $\mathbf{x}(t)\in X=W\times K\times P_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{n}_{w}+{n}_{k}+{{n}_{p}}_{\mathrm{\textnormal{other}}}}$ at time $t$ as:
$\mathbf{x}(t)=\begin{bmatrix}\mathbf{w}(t)\\ \mathbf{k}\\ \mathbf{p}_{\mathrm{\textnormal{other}}}(t)\end{bmatrix}.$ (34)
We perform this reordering and stacking because, as we see later in Lem. 10, the $\mathbf{w}$ and $\mathbf{k}$ are ETI states, so it indexing them as the first states of $\mathbf{x}$ simplifies avoid set computation. We similarly augment the goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ and the obstacles $\overline{\mathcal{O}}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\overline{\mathcal{O}}_{i}$, which we assume are union of H-polytopes:
$\begin{split}P_{\mathrm{\textnormal{G}}}&=\overline{P_{\mathrm{\textnormal{G}}}}\times K\times P_{\mathrm{\textnormal{other}}}\\ \mathcal{O}_{i}&=\overline{\mathcal{O}}_{i}\times K\times P_{\mathrm{\textnormal{other}}}.\end{split}$ (35)
IV-A2 PWA Planning Model
We approximate the planning model $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ as a discrete-time, time-variant PWA to enable computation. First, we define linearization points $\mathbf{x}^{*}_{t,i}\in X_{t}$, $i=1,\cdots,{n}^{*}_{t}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which can be generated by gridding or uniform sampling. We define the PWA regions as the Voronoi partition of the linearization points [32] (see Appendix -C). Each $i$th region is an H-polytope $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$. Then, in each region, at each time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define the dynamics $\mathbf{C}_{i,t}$ and $\mathbf{d}_{i,t}$ by Taylor expanding about the linearization points:
$\mathbf{x}(t+\Delta t)=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})+\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{\mathbf{x}={\mathbf{x}^{*}_{i,t}}}(\mathbf{x}(t)-\mathbf{x}^{*}_{i,t}),$ (36a)
$\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))=\begin{bmatrix}(\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{1:{n}_{w}}\\ \mathbf{k}\\ (\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{{n}_{w}+1:{n}_{p}}\\ \end{bmatrix}$ (36e)
$\mathbf{C}_{i,t}=\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}},\quad\mathrm{\textnormal{and}}$ (36f)
$\mathbf{d}_{i,t}=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})-\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}}\mathbf{x}^{*}_{i,t}.$ (36g)
Depending on the planning model, there may be some PWA regions with identical dynamics. If the union of the regions is convex, we combine them into a single PWA region (note, the linearization points are discarded).
IV-B Backward Reach Set without Tracking Error
Without accounting for tracking error, the reach set of a planning model is the set of initial conditions $\mathbf{x}_{0}$ of trajectories that are guaranteed to reach the goal region $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$. Instead of naively computing the full reach set, which scales exponentially in memory, computation time, and a number of set representations with the number of timesteps, PARC is planning aware: given an expert plan, we compute only a local, underapproximated BRS efficiently for the neighborhood that shares the same mode sequence as the expert plan. Fixing the mode sequence reduces the PWA system to a time-variant affine system, so the BRS is a single H-polytope using the inverse affine map of the goal (see (8)), with computation time scaling linearly with the number of timesteps and quadratically with dimensions from matrix multiplication.
Assumption 7 (Expert Plan w/o Tracking Error).
We have access to at least one goal-reaching initial condition $\mathbf{x}_{0}$ for which the PWA system reaches the goal without necessarily avoiding obstacles:
$\mathbf{\psi}\!\left(\mathbf{x}_{0}\right)\in P_{\mathrm{\textnormal{G}}}.$ (37)
Obtaining expert plans is system-dependent, but is not difficult, because we do not require the expert plan to avoid obstacles. Thus, one option is to solve a two-point boundary value nonlinear MPC problem, for which solution guarantees exist [33]. In practice, it is often sufficient to uniformly sample trajectory parameters $\mathbf{k}$ and roll out the PWA planning model, since the robot starts out near the goal. This assumption is even easier to fulfill if the planning model is time-variant affine (which we use for a general 3D quadrotor and drift vehicle planning in Section V, VI) since plans have only one mode.
If Assumption 7 is satisfied, we can obtain $\mathbf{x}_{0}$’s mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$ using (3). That is,
$s_{t}$ $=\min\{i\ |\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\},$ (38)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$. Then, we can compute the exact $t_{\mathrm{\textnormal{f}}}$-time BRS using the method in [19]:
Proposition 8 (Reach Set w/o Tracking Error).
Consider a mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Then the exact $t_{\mathrm{\textnormal{f}}}$-time BRS $\overline{\Omega}_{0}$ of the goal for the affine dynamics associated with $S$ is
$\begin{split}\overline{\Omega}_{t_{\mathrm{\textnormal{f}}}}&=P_{\mathrm{\textnormal{G}}},\\ \overline{\Omega}_{t}&=\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)\end{split}$ (39)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
Proof.
This follows directly from Lemma 1. ∎
This means the reference trajectories generated from any augmented planning state within $\overline{\Omega}_{0}$ are guaranteed to reach $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$ with mode sequence $S$. This process, along with an expert plan, is illustrated in Fig. 2.
Figure 2: (Right) Visualization of the reach set computation in Section IV-B for the TurtleBot example (Section III-D), projected onto the planning state space $P$, when $\mathbf{p}_{0}=[-4,0,\pi/5]^{\intercal}$, $t_{\mathrm{\textnormal{f}}}=4(s)$. The goal set is the green polytope extended to all $\theta\in[-\pi,\pi]$. An expert plan, computed by solving a two-point boundary value problem, passes through two modes (blue and magenta line). We compute the corresponding $t_{\mathrm{\textnormal{f}}}$-time BRS $\overline{\Omega}_{0}$ (blue polytope with solid outline). Intermediate BRSs $\overline{\Omega}_{t}$ are the polytopes without outlines, colored by mode. (Left) Illustration of the trajectory parameter space; the grey region represents $K$, the green region represents the projection of $\overline{\Omega}_{0}$ on $K$. After reach-set computation, the state $\mathbf{x}_{\mathrm{\textnormal{sample}}}=[\mathbf{p}_{\mathrm{\textnormal{sample}}}^{\intercal},\mathbf{k}_{\mathrm{\textnormal{sample}}}^{\intercal}]^{\intercal}$ is sampled from $\overline{\Omega}_{0}$: black dot (right) denotes the $\mathbf{p}_{\mathrm{\textnormal{sample}}}$ and black star (left) denotes the $\mathbf{k}_{\mathrm{\textnormal{sample}}}$. The resulting trajectory is denoted as the black line which reaches the goal.
IV-C PARC: Avoid Set without Tracking Error
Without accounting for tracking error, the avoid set is the set of pairs of initial planning model state and trajectory parameter such that any plans generated from outside the avoid set are guaranteed to not collide with the obstacles at all continuous time before $t_{\mathrm{\textnormal{f}}}$. Recall that, per (7), we model the continuous-time behavior as straight line segments between the discrete states of the PWA system.
Intuitively, the avoid set identifies all trajectories where the straight lines connecting the PWA planning model’s discrete states collide with any obstacles:
Proposition 9 (Unsafe Plans w/o Tracking Error).
Consider the obstacle $\mathcal{O}_{i}$, some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Suppose we compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ and the $(t_{\mathrm{\textnormal{f}}}-t-\Delta t)$-time BRS $\overline{\Omega}_{t+\Delta t}$ as in (39). If any trajectories between $\overline{\Omega}_{t}$ and $\overline{\Omega}_{t+\Delta t}$ collide with $\mathcal{O}_{i}$, that is, if $\exists\ \mathbf{x}(t)\in\overline{\Omega}_{t},\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\in\overline{\Omega}_{t+\Delta t}$ such that
[…]
then
$\mathcal{O}_{i}\cap\mathrm{\textnormal{conv}}\!\left(\overline{\Omega}_{t},\overline{\Omega}_{t+\Delta t}\right)\neq\emptyset.$ (41)
Proof.
By the definition of convex-hull, we have
[…]
$\mathbf{x}(t),\mathbf{x}(t+\Delta t)\in\overline{\Omega}_{t}\bigcup\overline{\Omega}_{t+\Delta t}\},$ (42a)
[…]
$\mathbf{x}(t)\in\overline{\Omega}_{t},\mathbf{x}(t+\Delta t)\in\overline{\Omega}_{t+\Delta t}\}.$ (42b)
Thus (40) necessarily implies (41). ∎
Next, before we can overapproximate the avoid set between two-time steps, we must adapt our notion of extended translation invariance to the PWA planning model:
Lemma 10 (PWA ETI).
Suppose $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in the workspace as in Assum. 2. Also suppose we convert it to a PWA system as in Section IV-A2 and (IV-A2). Then in mode $i$ at time $t$, each state $\mathbf{x}$ of the PWA system obeys
$\left\|\hat{\mathbf{C}}_{i,t}^{W}{(\hat{\mathbf{C}}_{i,t}\mathbf{x}+\mathbf{d}_{i,t})_{1:{n}_{w}}}\right\|_{\mathrm{\textnormal{2}}}=0\quad\forall\mathbf{x}\in{\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})}$ (43)
where $\hat{\mathbf{C}}_{i,t}=\mathbf{C}_{i,t}-\mathbf{I}_{{n}_{x}}$ and $\hat{\mathbf{C}}^{W}_{i,t}$ denotes the ${n}_{w}\times{n}_{w}$ upper-left submatrix of $\hat{\mathbf{C}}_{i,t}$, namely $\hat{\mathbf{C}}_{i,t}^{W}\triangleq(\hat{\mathbf{C}}_{i,t})_{1:{n}_{w},1:{n}_{w}}$
Proof.
This follows directly from Assum. 2 and (IV-A2). ∎
The definition of PWA-ETI (43) in other subspace is similarly defined by substituting $W$ to the subspace of interest. Note that $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in $W\times K$ since the trajectory parameter $\mathbf{k}\in K$ is constant.
Now, for avoid-set computation, it is helpful to relabel the ETI states of $\mathbf{x}(t)$, identified per Lemma 10:
$\mathbf{x}(t)$ $=\begin{bmatrix}\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\\ \mathbf{x}_{\mathrm{\textnormal{other}}}(t)\end{bmatrix},$ (46)
where $\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\in X_{\mathrm{\textnormal{ETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}$ are the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ states in $\mathbf{x}(t)$ that are ETI-states, and $\mathbf{x}_{\mathrm{\textnormal{other}}}(t)\in X_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{other}}}}$ are the remaining states in $\mathbf{x}(t)$. Thus ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}\geq{n}_{w}+{n}_{k}$. Note that all sets and matrices including $\mathcal{O}$, and $P_{\mathrm{\textnormal{G}}}$ are relabeled accordingly. This reordering is only to simplify the exposition.
We are now ready to state our first key contribution: overapproximating the avoid set without tracking error. The following theorem overapproximates the continuous-time BRS of an obstacle and intersects it with the discrete-time reach-set of the PWA planning model to identify unsafe plans.
Theorem 11 (Avoid set w/o Tracking Error).
Consider the obstacle $\mathcal{O}_{i}$, some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and some mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Suppose we compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ as in (39). Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\mathcal{O}_{i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{other}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$, and define the intermediate avoid set $\overline{\Lambda}_{i,t,t}\subset\overline{\Omega}_{t}$ as:
$\overline{\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{other}}},\mathcal{O}_{i}\right)\cap\overline{\Omega}_{t}$ (47)
Then, for any $\mathbf{x}(t)\in\overline{\Omega}_{t}\setminus\overline{\Lambda}_{i,t,t}$, $\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}$,
[…]
Proof.
See Appendix -D. ∎
In summary, Proposition 9 establishes a necessary condition for the collision to occur between two discrete timesteps, while Theorem 11 overapproximates the avoid sets contained in each intermediate BRS.
We now propose a straightforward scheme to construct the avoid set using Proposition 9 and Theorem 11: reference trajectories generated from outside the avoid set are guaranteed to avoid the obstacle. First, for each timestep $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and for each obstacle $i=1,\cdots,{n}_{\mathcal{O}}$, we check if (41) is true; this reduces conservativeness because we identify obstacles for which we do not have to compute the avoid set $\overline{\Lambda}_{i,t,t}$ with Theorem 11, at the expense of additional convex hull computations. That is, one can speed up computation but potentially make the avoid set more conservative by skipping checking (41). Then, we apply the following corollary:
Corollary 12 (Avoid Set Computation).
We compute the avoid set $\overline{\Lambda}_{i,t,0}$ by computing the $t$-time BRS of $\overline{\Lambda}_{i,t,t}$. Mathematically,
$\overline{\Lambda}_{i,t,j}$ $=\mathcal{B}\!\left(\overline{\Lambda}_{i,t,j+\Delta t},\mathbf{C}_{s_{j},j},\mathbf{d}_{s_{j},j}\right)\ \forall j=0,\Delta t,\cdots,t-\Delta t.$ (49)
Then, the final, overapproximated avoid set $\overline{\Lambda}$ can be stored as a union of H-polytopes as:
[…]
Proof.
This follows as a consequence of Theorem 11 and since the BRS operation, $\mathcal{B}$ is exact for an affine system defined for a mode sequence (Lemma 1). ∎
To sample from within the reach set but not from the avoid set, one can either rejection-sample, sample via random-walk algorithms [34, 35], or compute the set difference between the reach set and the avoid set [36], and then sample via solving an LP. The reach set and the avoid set are shown in Fig. 3; notice that the avoid set is a nonconvex union of polytopes.
Figure 3: Visualization of the reach set (blue) and the avoid set (red) from Section IV-C for the TurtleBot example (Section III-D and Fig. 2), with an obstacle (light red), extended to all $\theta\in[-\pi,\pi]$. The expert plan used to compute the reach set is shown as a dashed black line; notice that it intersects the obstacle. However, PARC is still able to compute the reach set (blue) and thereby find plans that safely reach the goal set (green).
IV-D Estimating Tracking Error
Now that we have a basis for computing the BRAS, we will model tracking error such that, when incorporated into the PARC framework, we can produce plans that, when followed by the tracking model, guaranteed to reach the goal region while avoiding all obstacles.
Our method is based on [10, 37]. We sample $(\mathbf{z}_{0,j},\mathbf{k}_{j})\in\tilde{Z}\times\tilde{K}\subset Z\times K$ for $j=1,\cdots,{n}_{\mathrm{\textnormal{sample}}}$ uniformly, where $\tilde{Z}\times\tilde{K}$ is a hyperrectangle created by partitioning $Z\times K$; we do not index this partition to ease notation. We then have the robot start at $\mathbf{z}_{0,j}$ to track the plans parameterized by initial planning state $\mathbf{p}_{0,j}\triangleq(\mathbf{z}_{0,j})_{1:{n}_{p}}$ and trajectory parameter $\mathbf{k}_{j}$, and compute the error between the planned and realized trajectories in the workspace states (recall that we assume the goal and obstacles are subsets of workspace).
We consider two types of errors: the maximum final error $e_{t_{\mathrm{\textnormal{f}}},i}\in\mathbb{R}$ for goal-reaching, and maximum interval error $\tilde{e}_{t,i}\in\mathbb{R}$ for obstacle avoidance, defined for the discrete-time at time $t$ and the $i^{\mathrm{\textnormal{th}}}$ dimension of $\mathbf{z}$ and $\mathbf{p}$. PARC require two types of errors because it provides continuous-time guarantees for obstacle avoidance, but gives guarantees at the final timestep only for goal-reaching.
For each workspace coordinate $i=1,\cdots,{n}_{w}$, we define the maximum final error $e_{t_{\mathrm{\textnormal{f}}},i}$ as:
$e_{t_{\mathrm{\textnormal{f}}},i}=\max_{j}\left|\mathbf{p}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{p}_{0,j},\mathbf{k}_{j})-\mathbf{z}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0,j},\mathbf{k}_{j})\right|,$ (51)
where $\mathbf{p}_{i}(\cdot)$ and $\mathbf{z}_{i}(\cdot)$ are the $i^{\mathrm{\textnormal{th}}}$ element of $\mathbf{p}(\cdot)$ and $\mathbf{z}(\cdot)$. In plain words, $e_{t,i}$ represents the maximum error observed at the final time $t_{\mathrm{\textnormal{f}}}$ for the $i^{\mathrm{\textnormal{th}}}$ workspace dimension.
For $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, $i=1,\cdots,{n}_{w}$, we define the maximum interval error $\tilde{e}_{t,i}$ as:
$\tilde{e}_{t,i}=\max_{j,\ t^{\prime}\in[t,t+\Delta t]}\left|\mathbf{p}_{i}(t^{\prime};\mathbf{p}_{0,j},\mathbf{k}_{j})-\mathbf{z}_{i}(t^{\prime};\mathbf{z}_{0,j},\mathbf{k}_{j})\right|$ (52)
which is the maximum error observed between each discrete timestep for the $i^{\mathrm{\textnormal{th}}}$ workspace dimension.
We assume this approach provides an upper bound for feasible tracking error, which can be proven for simple planning models [10, 37].
Assumption 13 (Maximum Sampled Error Contains Maximum Error).
We assume that, for a large enough ${n}_{\mathrm{\textnormal{sample}}}$,
$e_{t_{\mathrm{\textnormal{f}}},i}$ $\geq\max_{\mathbf{z}_{0}\in\tilde{Z},\,\mathbf{k}\in\tilde{K}}\left|\mathbf{p}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{p}_{0},\mathbf{k})-\mathbf{z}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0},\mathbf{k})\right|,$ (53a)
$\tilde{e}_{t,i}$ $\geq\max_{\mathbf{z}_{0}\in\tilde{Z},\,\mathbf{k}\in\tilde{K},\,t^{\prime}\in[t,t+\Delta t]}\left|\mathbf{p}_{i}(t^{\prime};\mathbf{p}_{0},\mathbf{k})-\mathbf{z}_{i}(t^{\prime};\mathbf{z}_{0},\mathbf{k})\right|,$ (53b)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, where $\mathbf{p}_{0}=(\mathbf{z}_{0})_{1:{n}_{p}}$.
This assumption is clearly a potential weakness of the method, but we find that it holds in practice because the sampling process produces adversarial $(\mathbf{z}_{0},\mathbf{k})$ pairs. Furthermore, we can refine the tracking error estimate by increasing the fineness of the partition of $Z\times K$ into subsets $\tilde{Z}\times\tilde{K}$ [10].
To incorporate these errors into the geometric computation of PARC, we express them as the maximum final error set $E_{t_{\mathrm{\textnormal{f}}}}\subset\mathbb{R}^{{n}_{x}}$ and the maximum interval error set $\tilde{E}_{t}\subset\mathbb{R}^{{n}_{x}}$:
$E_{t_{\mathrm{\textnormal{f}}}}$ $=[-e_{t_{\mathrm{\textnormal{f}}},1},e_{t_{\mathrm{\textnormal{f}}},1}]\times\cdots\times[-e_{t_{\mathrm{\textnormal{f}}},{n}_{w}},e_{t_{\mathrm{\textnormal{f}}},{n}_{w}}]\times\underbrace{\{0\}\times\cdots\times\{0\}}_{({{n}_{p}}_{\mathrm{\textnormal{other}}}+{n}_{k})\mathrm{\textnormal{ times}}},$ (54)
$\tilde{E}_{t}$ $=[-\tilde{e}_{t,1},\tilde{e}_{t,1}]\times\cdots\times[-\tilde{e}_{t,{n}_{w}},\tilde{e}_{t,{n}_{w}}]\times\underbrace{\{0\}\times\cdots\times\{0\}}_{({{n}_{p}}_{\mathrm{\textnormal{other}}}+{n}_{k})\mathrm{\textnormal{ times}}}.$ (55)
It is straightforward to express these intervals as H-polytopes [22]. The Cartesian product with empty sets is necessary to augment the states for Minkowski sum and Pontryagin difference operations with the obstacles and the goal.
An example of computed tracking error for a near-hover quadrotor dynamical model is shown in Fig. 8.
IV-E Incorporating Tracking Error into PARC
To conclude our proposed method, we incorporate tracking error into PARC’s BRAS computation, thereby transferring guarantees for the planning model to the tracking model. In short, we get reach-avoid guarantees that incorporate tracking error by simply Minkowski summing the error with obstacles and Pontryagin differencing from the goal set.
IV-E1 PARC: Reach Set with Tracking Error
From Assumption 13, we assume the sampled tracking error is valid only over the compact hyperrectangle $\tilde{Z}\times\tilde{K}$. Thus, we must define a valid input region $\tilde{X}\subset X$ for which the computed reach set will have guarantees over:
$\tilde{X}=\mathrm{\textnormal{proj}}_{1:{n}_{w}}\!\left(\tilde{Z}\right)\times\tilde{K}\times\mathrm{\textnormal{proj}}_{({n}_{w}+1):{n}_{p}}\!\left(\tilde{Z}\right).$ (56)
The projection and Cartesian product are necessary to reorder $\tilde{Z}$ and $\tilde{K}$ into that of (34).
To incorporate tracking error, we first tighten our assumption on the expert plan to find a mode sequence:
Assumption 14 (Expert Plan with Tracking Error).
We assume that a goal-reaching initial condition $\mathbf{x}_{0}$ that defines an expert plan is known and given for the PWA planning model in Section IV-A. Particularly, $\mathbf{x}_{0}$ is any point that satisfies:
$\mathbf{\psi}\!\left(\mathbf{x}_{0}\right)\in P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}}.$ (57)
Intuitively, $E_{t_{\mathrm{\textnormal{f}}},i}$ captures the maximum tracking error at time $t_{\mathrm{\textnormal{f}}}$: if the plans can arrive at the smaller goal region $P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}}$, then the robot should be guaranteed to arrive at the original, bigger goal region $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$.
With a given expert plan, we can obtain $\mathbf{x}_{0}$’s mode sequence using (38), then compute the underapproximated reach set $\Omega_{0}$ of the goal similar to Proposition 8.
Proposition 15 (Reach Set with Tracking Error).
Consider a mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$.. Then we can underapproximate the reach set $\Omega_{0}$ corresponding to this mode sequence by:
$\Omega_{t_{\mathrm{\textnormal{f}}}}=$ $P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}},$ (58a)
$\Omega_{t}=$ $\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$
$\forall t=\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t,$ (58b)
$\Omega_{0}=$ $\mathcal{B}\!\left(\overline{\Omega}_{\Delta t},\mathbf{C}_{s_{\Delta t},\Delta t},\mathbf{d}_{s_{\Delta t},\Delta t}\right)\cap\tilde{X},$ (58c)
Proof.
This follows from Lemma 1, Assumption 13, and Proposition 8. ∎
This extends Proposition 8 by shrinking the goal set and intersection with the tracking error data coverage domain to include tracking error. With $\Omega_{0}$ computed, if the robot tracks any plan generated from an initial condition within $\Omega_{0}$, it is guaranteed to reach the goal at time $t_{\mathrm{\textnormal{f}}}$, but may not be safe; next we guarantee safety by computing the avoid set.
IV-E2 PARC: Avoid Set with Tracking Error
Consider the obstacle $\mathcal{O}_{i}$ and some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$. We define the $i^{\mathrm{\textnormal{th}}}$ buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ between time $t$ and $t+\Delta t$ as:
$\tilde{\mathcal{O}}_{t,i}$ $=\mathcal{O}_{i}\oplus\tilde{E}_{t}.$ (59)
Assuming $\tilde{E}_{t}$ captures the maximum tracking error from $t$ to $t+\Delta t$, by buffering $\mathcal{O}_{i}$ with $\tilde{E}_{t}$, if the trajectory does not collide with $\tilde{\mathcal{O}}_{t,i}$, then the robot following the trajectory should never collide with $\mathcal{O}_{t,i}$. Thus, incorporating tracking errors into the PARC formulation for avoid set follows closely to that of Section IV-C, but with $\tilde{\mathcal{O}}_{t,i}$ instead of $\mathcal{O}_{t,i}$ as an obstacle.
Proposition 16 (Unsafe Plans w/ Tracking Error).
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ and the $(t_{\mathrm{\textnormal{f}}}-t-\Delta t)$-time BRS ${\Omega}_{t+\Delta t}$ using (15). If any planning trajectories between ${\Omega}_{t}$ and ${\Omega}_{t+\Delta t}$ collide with $\tilde{\mathcal{O}}_{t,i}$, that is, if $\exists\ \mathbf{x}(t)\in{\Omega}_{t},\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\in{\Omega}_{t+\Delta t}$ such that
[…]
then
$\tilde{\mathcal{O}}_{t,i}\cap\mathrm{\textnormal{conv}}\!\left({\Omega}_{t},{\Omega}_{t+\Delta t}\right)\neq\emptyset.$ (61)
Proof.
Follows from Propositions 15 and Proposition 9. ∎
Finally, we can state our second key contribution: overapproximating the avoid set with tracking error (i.e., computing the BRAS).
Theorem 17 (Avoid set w/ Tracking Error).
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and from (15), the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ of the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\tilde{\mathcal{O}}_{t,i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{other}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$. Define the intermediate avoid set ${\Lambda}_{i,t,t}\subset{\Omega}_{t}$ as:
${\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{other}}},\tilde{\mathcal{O}}_{t,i}\right)\cap{\Omega}_{t}.$ (62)
Then, for any $\mathbf{x}(t)\in{\Omega}_{t}\setminus{\Lambda}_{i,t,t}$, $\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}$,
[…]
Proof.
Follows from Proposition 16 and Theorem 11. ∎
We can adapt Corollary 12 to compute the avoid set ${\Lambda}_{i,t,0}$ by computing the $t$-time BRS of ${\Lambda}_{i,t,t}$ as
${\Lambda}_{i,t,j}$ $=\mathcal{B}\!\left({\Lambda}_{i,t,j+1},\mathbf{C}_{s_{j},j},\mathbf{d}_{s_{j},j}\right)\ \forall j=0,\Delta t,\cdots,t-\Delta t.$ (64)
Then, the final overapproximated avoid set ${\Lambda}$ can be stored as a union of H-polytopes as:
[…]
With the avoid set and the reach set from Section IV-E1 computed, if the robot tracks any plans starting in $\Omega_{0}\setminus\Lambda$ it is guaranteed to reach the goal region while avoiding all obstacles. Thus, the set $\Omega_{0}\setminus\Lambda$ is the desired object $\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})$. Next, we demonstrate PARC’s BRAS computation on a variety of numerical examples.
V Experiments
We now assess the utility of our proposed PARC method for generating safe trajectory plans near danger. We seek to understand the impact of planning model design (Section V-A), cooperative vs. adversarial planner-trackers (Section V-B), and tight approximation of planning model reachable sets (Section V-C). We also assess the utility of learning-based reach-avoid in Sections V-A and V-B.
Key result: PARC outperforms other methods in goal reaching, safety, and computation time, but requires careful choice of planning model.
To ensure fair assessment, we apply PARC to robots from the safe motion planning and control literature: a near-hover quadrotor in 2-D [15] and 3-D [5], and a more general quadrotor model in 3-D [10]. Implementation details are reported in Appendix -E. All experiments, and training of learning-based methods, are run on a desktop computer with a 24 core i9 CPU, 32 GB RAM, and an Nvidia RTX 4090 GPU.
V-A Impact of Planning Model Design
This experiment seeks to understand the impact of choosing a planning model for a given tracking model. We deploy PARC on a tracking model and environment that have been successfully solved by a Neural CLBF [15] approach.
To examine the importance of the planning model, we employ the same dynamics for both the planning and tracking model, namely 2-D near-hover quadrotor dynamics:
$\dot{\mathbf{p}}=\dot{\mathbf{z}}=\frac{d}{dt}\begin{bmatrix}p_{x}\\ p_{z}\\ \theta\\ v_{x}\\ v_{z}\\ \omega\end{bmatrix}=\left[\begin{array}[]{l}v_{x}\\ v_{z}\\ \omega\\ \tfrac{1}{m}\sin(\theta)({F_{\mathrm{\textnormal{l}}}}+{F_{\mathrm{\textnormal{r}}}})\\ \tfrac{1}{m}\cos(\theta)({F_{\mathrm{\textnormal{l}}}}+{F_{\mathrm{\textnormal{r}}}})-g\\ \tfrac{r}{I}({F_{\mathrm{\textnormal{l}}}}-{F_{\mathrm{\textnormal{r}}}})\end{array}\right]$ (78)
The workspace states are $\mathbf{w}=\begin{bmatrix}p_{x}&p_{y}\end{bmatrix}^{\intercal}$, and the inputs are left- and right-side propeller forces
$\mathbf{u}=\mathbf{k}$ $=\begin{bmatrix}{F_{\mathrm{\textnormal{l}}}}&{F_{\mathrm{\textnormal{r}}}}\end{bmatrix}^{\intercal},$
which are held constant for $[0,t_{\mathrm{\textnormal{f}}}]$ to generate plans. The feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ produces time-varying left- and right-side propeller forces using an iLQR controller [31].
[…]
V-A1 Experiment Setup
We use the obstacle and goal environment from [15, Sec. 6.3], wherein the quadrotor must navigate over an obstacle and through a small gap to a goal region (see the left side of Fig. 4 and Fig. 5).
We use PARC to compute a BRAS for this scenario with an expert plan, accounting and without accounting for tracking error. The planning model piece-wise affinize (78) with 80 PWA regions, timestep $\Delta t=0.1$unitsec, and final time $t_{\mathrm{\textnormal{f}}}=2$unitsec. We assess the computation time, as well as the size of the BRAS to investigate the impact of planning model design.
V-A2 Hypothesis
We expect PARC to be able to compute plans to the goal reliably, with no safety violations during $[0,t_{\mathrm{\textnormal{f}}}]$. However, due to the naïve choice of planning model, we expect the avoid set to be overly conservative, especially since (78) has three non-ETI states ($v_{x},v_{z},\omega$) in its PWA formulation.
V-A3 Results
[…]
Figure 4: Results of PARC on a naïve choice of planning model, without accounting for tracking error. Green indicates the goal region, red indicates the obstacles with the volume of the drone accounted for, and blue shows the reach set set-differenced with the avoid set without tracking error. Safe trajectories were successfully generated from the left of the leftmost obstacle.
Figure 5: Results of PARC on a naïve choice of planning model after accounting for tracking error. Green indicates the goal region, red indicates the obstacles with the volume of the drone accounted for, and blue shows the computed BRAS. There were no longer any safe trajectories generating from the left of the leftmost obstacle, which the Neural CLBF was able to accomplish.
V-A4 Discussion
The decrease in BRAS volume was expected due to the many non-ETI states, leading to overconservativeness in avoid set computation. Moreover, as the problem framework requires the trajectory parameters ${F_{\mathrm{\textnormal{l}}}},{F_{\mathrm{\textnormal{r}}}}$ to be constant, the quadrotor shall continue accelerating after reaching the goal region, leading to instability should ${F_{\mathrm{\textnormal{l}}}},{F_{\mathrm{\textnormal{r}}}}$ be tracked.
In Appendix -G and the next few examples, we show how a careful design of the planning model would lead to much better performance on more complicated scenarios, as well as enable stabilizing behavior such as braking without explicitly modeling velocity as a goal state.
V-B Impact of Planner-Tracker Cooperation
We now study cooperative vs. adversarial planning and tracking models. We also further study the Neural CLBF [15] learning-based reach-avoid approach. For this experiment, we use a 3D near-hover quadrotor introduced for FaSTrack [5], a reachability-based planner tracker that uses adversarial planning and tracking models.
V-B1 Experiment Setup
The tracking model and PARC’s planning model are listed in Appendix -E. The quadrotor must navigate a narrow gap (width 1.8 m, compared to drone width 0.5 m) to reach its goal as shown in Fig. 7. We simulate 8,100 trials from different initial conditions. We measure the number of trials that reach the goal, and the number that crashes. We also consider the total computation time of each method.
V-B2 Hypotheses
We anticipate that our PARC planner always reaches the goal safely. We also anticipate faster computation time of our reachability analysis versus the other two approaches, but note that FaSTrack does not need to recompute the reachable set for every obstacle. Since FaSTrack provides formal proof of safety, we anticipate it has no crashes. However, since it treats the planner and tracker as adversarial, we expect it to struggle to reach the goal. On the other hand, since CLBFs specifically balance safety and liveness, we expect it to reach the goal more often than FaSTrack but at the expense of crashing due to the high chance of learning an incorrect representation of safety when forced to operate near obstacles.
V-B3 Results
Fig. 7 illustrates results for all methods with an initial position of $[4,-1,3]^{\intercal}$. FaSTrack (orange) fails to find a feasible plan that reaches the goal, the trajectory rolled out by Neural CLBF (magenta) violates safety, while the trajectory rolled out by PARC (blue) finds a plan (black) that reaches the goal safely despite tracking error.
Fig. 6 illustrates the trajectories simulated on the grid of 8,100 initial positions using each method. Neural CLBF finds 576 successful trajectories (7.1 %), while safety is violated for the rest. FaSTrack finds 0 successful trajectories but never violates safety. However, PARC finds the plan to reach the goal without collision for 2,592 initial states (32 %).
[…]
Figure 6: Simulated trajectories of 10D quadrotor for 8,100 initial states using three methods: PARC, FaSTrack [5], Neural CLBF [15]. The green cuboid is the goal set and the pink cuboids are obstacles.
Figure 7: Realized trajectory of the 10D quadrotor in the environment for 10 seconds using three methods: ours (blue), FaSTrack [5] (orange), Neural CLBF [15] (magenta). The black dashed line denotes the plan computed using PARC. The goal is represented as a green cube while the obstacle is magenta cuboids. The boundary of environments exist although not illustrated in the figure.
Figure 8: Comparison of tracking error function between Fastrack and PARC for planning horizon $t_{\mathrm{\textnormal{f}}}=10$unitsec. FaSTrack computes a fixed, worst-case tracking error bound for all time because it assumes an adversarial planner, whereas PARC takes advantage of the smaller, time-varying tracking error that results from a planner and controller working together. The pale pink trajectories are samples used to compute the error as in Section IV-D.
V-B4 Discussion
The performance of FaSTrack is notably compromised by the large, conservatively computed TEB, specified as a $2.47\times 2.47\times 0.1$ cuboid. This often results in the inability to compute feasible plans in spaces tighter than the TEB. However, as seen in the Fig. 8, sampled tracking error under parameterized plan, the tracking error is much smaller than TEB, enabling PARC to find a plan in challenging scenarios.
Although Neural CLBF succeeds in capturing most of the safe, unsafe, goal regions, the synthesized control policy steers some of the initial states to violate safety. Thus, the learning-based method sacrifices guarantees.
PARC has a strong safety guarantee, so every plan PARC found did not collide with the obstacles. We note that a better choice of planning model could increase PARC’s success rate; the simplicity of the single-integrator planning model, employed mainly for fair comparison with FaSTrack, reduced the flexibility of possible plans. Indeed we leverage a better choice of trajectory model next.
V-C Impact of Tight Approximation for Planning Reachability
Finally, we seek to understand the importance of computing a tighter approximation of reachable set for the planning model. To do this, we compare against RTD implemented on a general 3-D quadrotor [10] using zonotope reachability [23]. This allows us to use the same tracking error model but employs a different approach to planning model reachability analysis.
V-C1 Experiment Setup
Our experimental setup is similar to that of Section V-B but with a narrower gap.
We use the same planning model as quadrotor RTD [10], which generates time-varying polynomial positions in 3-D [38] with $\Delta t=0.02$unitsec with $t_{\mathrm{\textnormal{f}}}=3$unitsec; this formulation is affine time-varying in the trajectory parameters with no non-ETI states, so reachability can be tightly approximated by PARC. Both PARC and RTD use the same tracking error that is precomputed in [10].
[…]
We declare success if the quadrotor reaches the goal region within 30 seconds without colliding with any obstacles.
V-C2 Hypothesis
We anticipate that PARC will always safely find a way to reach the goal since we know the narrow gap is large enough to do so, and PARC’s primary source of conservativeness is its tracking error estimation (which is identical to that of RTD). We anticipate that RTD will reach the goal much less often because it suffers an additional course of conservativeness from using an approximate method to compute the reachable set of its planning model.
V-C3 Results
[…]
V-C4 Discussion
As expected, both RTD and PARC guaranteed safety. However, only PARC was able to traverse through the narrow gap. Thus, this example clearly illustrates the utility of PARC’s tightly approximating reachable set computation in comparison to RTD’s approximation. Although PARC is significantly slower in planning when compared to RTD, once a BRAS is computed, the robot obtains a continuum of safe plans.
[…]
V-D Overall Discussion
Through the above experiments, we have shown how our proposed PARC approach confers advantages in computing and using Backward Reach-Avoid Set (BRAS). By careful use of PWA systems, plus bounded tracking error, we are able to generate safe motion plans in cases where other methods are either too conservative or fail to compute entirely.
However, PARC still has room for improvement. Most obviously, the safe plan computation time is on the order of seconds for just a few obstacles; this is too slow for real-time operation. Furthermore, the planning model requires careful hand-crafting as shown in our experiments.
VI Drift Vehicle Demonstration
We now demonstrate the utility of PARC in safely planning extreme near-danger vehicle drifting maneuvers, where the robot loses controllability.
Key result: PARC enables computation of provably safe drift parking trajectories for the first time, extending the state of the art.
VI-A Background
VI-A1 Drifting
Drifting is when a vehicle turns with a high sideslip angle, causing tire saturation and losing direct control of its heading [39]. Stable drifting can be achieved by finding “drift equilibria” with constant yaw rate and sideslip angle of the highly nonlinear dynamics [39, 40, 41, 42]. Drift parking, however, remains challenging due to the highly coupled inputs [43] and reduced controllability as the vehicle specifically must not maintain a drift equilibrium [44].
VI-A2 Reach-Avoid Method Comparison
To the best of our knowledge, no other method can guarantee safety and liveness for drift parking.
Our drifting tracking model dynamics (see (122) below) are not control affine, so directly applying a CBF or CLBF control approach is not tractable [45, 46]; future work may apply results for non-control-affine systems [47, 48]. Note, [15] poses a control affine vehicle dynamics model but does not include tire saturation, so cannot drift. Furthermore, our planning and tracking models are too high-dimensional for standard HJB value function computation [5, 49]; future work will compare against learning-based HJB [13]. The most similar work is RTD [9, 50], which has not been applied to these extreme dynamics. As a future implementation, RTD’s reachable set can use PARC’s formulation to enable a direct drifting comparison; then, the only difference with PARC is that RTD implicitly represents the avoid set. Due to this similarity, we focus on demonstrating PARC’s capabilities as opposed to a head-to-head comparison.
VI-B Demo Details
Extensive details are presented in Appendix -H. Here, we briefly present setup, simulation results, and a discussion.
VI-B1 Planning and Tracking Models
As we saw in Section V-A, carefully planning model design is critical to PARC’s reach-avoid. In this case, we fit an affine time-varying model to drifting data collected by rolling out open-loop drifting maneuvers. This not only guarantees a reasonable planning model (i.e., it successfully parameterizes drifting maneuvers) but also demonstrates the flexibility of using PWA models for PARC since they can readily be learned from expert demonstrations.
We use a tracking model based on [51, 40, 42, 39], detailed in Appendix -H. Our tracking controller is based on [51]: we first use nonlinear MPC to bring the car into a drifting state, then switch to open-loop control to complete the parking action (since the vehicle loses controllability when sliding sideways to a stop).
VI-B2 Environment and Demo Setup
The objective of this demo is to drift safely into a parallel parked state between two closely parked cars, as shown in Fig. 1 and Fig. 10. Our environment is inspired by a real-world demo video [52].
We assess PARC’s ability to compute the BRAS for this maneuver. We further evaluate safety trajectory rollouts using initial conditions and trajectory parameters sampled from the BRAS to check for collision.
VI-B3 Results
Figure 10: Example planned drifting trajectories sampled from the Backward Reach-Avoid Set using PARC. The two red cars are obstacles. The goal set for the center of mass is in green. The trajectory with a very low tracking error is shown in (a), whereas a high tracking error is shown in (b). Critically neither collide with the obstacles, and both are represented in the single BRAS.
[…]
VI-B4 Discussion
This demo showcased (i) PARC’s capability to generate safe and goal-reaching trajectories with challenging near-danger tasks such as reach-avoid drift parking, and (ii) the flexibility of PARC in terms of handling a suitable planning model from data. We notice that the BRAS is relatively small (see Fig. 0(b)) in the state space, indicating that drift parking is indeed a difficult maneuver. This emphasizes the importance of PARC minimizing conservativeness in the BRAS computation. However, it also suggests that this maneuver is brittle to large changes in initial conditions, and PARC will struggle to find drifting maneuvers for different goal sets or obstacle configurations. In future work, we plan to test more robust drifting controllers to produce a larger BRAS. With that in mind, our ultimate goal is physical hardware experiments on a platform such as the F1:10 car in Fig. 12.
VII Conclusion
This paper has presented an approach to use Piecewise Affine Reach-avoid Computation (PARC) for safe goal-reaching robot trajectory planning near danger. PARC enables goal-reaching through narrow gaps and in tight obstacle configurations, even with extreme dynamics. The method outperforms a variety of state-of-the-art reach-avoid methods and establishes a new benchmark for safe, extreme vehicle motion planning. Future work will improve PARC’s computation time,
Limitations. PARC has several key limitations. First and foremost, it requires an assumption that the data coverage of sampled tracking error is large enough. While this assumption holds in practice (over hundreds of collision-free trials planned in this paper), it means that further mathematical work is necessary for true safety guarantees. Second, although PARC is faster than compared methods, it is still too slow for real-time; the present implementation has not been parallelized and may benefit from alternative representations to H-polytopes. Third, PARC has only been compared on hand-crafted examples of goal reaching near danger, and only in simulation. It is critical to validate the method across randomly-generated, realistic environments, on real hardware. To this end, the authors are preparing a small robotic car to perform hardware experiments in the coming months.