### v3

# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

Abstract
Autonomous mobile robots must maintain safety, but should not sacrifice performance, leading to the classical reach-avoid problem. This paper seeks to compute trajectory plans for which a robot is guaranteed to reach a goal and to avoid obstacles in the specific near danger case, also known as a narrow gap, where the agent starts near the goal, but must navigate through tight obstacles that block its path. The proposed method builds off of a common approach of using a simplified planning model to generate plans, which are then tracked using a high-fidelity tracking model and controller. Existing safe planning approaches use reachability analysis to overapproximate the error between these models, but this introduces additional numerical approximation error and thereby conservativeness that prevents goal reaching. The present work instead proposes a Piecewise Affine Reach-avoid Computation (PARC) method to tightly approximate the reachable set of the planning model. PARC significantly reduces conservativeness through a careful choice of the planning model and set representation, along with an effective approach to handling time-varying tracking errors. The utility of this method is demonstrated through extensive numerical experiments in which PARC outperforms state-of-the-art reach avoid methods in near-danger goal reaching. Furthermore, in a simulated demonstration, PARC enables generation of provably-safe extreme vehicle dynamics drift parking maneuvers. A preliminary hardware demo on a TurtleBot3 also validates the method.
I Introduction
The increasing deployment of autonomous robots in daily tasks necessitates addressing complex operational challenges. Many of these tasks can be conceptualized as reach-avoid problems, wherein a robot must steer its states into a set of goal states without violating safety-critical constraints [fisac2015reach]. Achieving both liveness – the assurance of goal fulfillment [alpern1985defining] – and adherence to safety constraints is pivotal in these scenarios. In this paper, we are particularly interested in collision-free goal reaching near danger, meaning that the agent starts near the goal, but must navigate through tight obstacles—typically within the width of a robot’s body—that block its path. As we will show, this problem setup is very difficult for existing robust planning and control approaches to provide reach-avoid guarantees on general dynamical systems due to over-conservativeness. We note that the specific “near danger” condition where these planners and controllers fail is task-dependent. Some examples include flying a quadrotor drone through a narrow gap, or parallel parking an autonomous car.
I-A Reach-Avoid Approaches and Challenges
The reach-avoid problem is typically solved with one of two approaches, both based on reachability analysis or computing reachable sets of a system’s trajectories. Note that an extended literature review is presented in Appendix -A.
Figure 1: Our overall goal is to find safe goal-reaching motion plans near danger. In (a), we show the components of our proposed Piecewise Affine Reach-avoid Computation (PARC) method, with the paper section number on the top left, and the corresponding symbol on the bottom. This method enables a new state-of-the-art for safe reach-avoid problems, as evidenced by (b), which shows examples of extreme vehicle maneuvers. These planned drift parking trajectories are sampled from the Backward Reach-Avoid Set (BRAS, green triangle on the top left) using PARC. Our vehicle is in a blue time lapse. The two red cars are obstacles. The goal set for the center of mass is the green rectangle. The small size of the BRAS indicates the challenge of finding safe trajectories.
The first approach is control intervention, wherein a motion planner proposes actions, while a safety filter [wabersich2023data, hsu2023safety] prevents the robot from executing actions for which its reachable set would intersect obstacles. This approach often requires compensating for the worst-case behavior of the motion planner [chen2021fastrack, garg2021multi], which we find leads to conservativeness (i.e., sacrificing liveness) in our experiments (Section V). The second approach is a planner-tracker framework, wherein a motion planner proposes plans that are safe by leveraging the knowledge of a low-level controller’s tracking ability. This controller can be synthesized [majumdar2017funnel, tedrake2010lqr] or hand-engineered [kousik2020bridging, kousik2019safe]; the key difference is that the planner need not be treated as adversarial, which reduces conservativeness. However, this approach typically limits the space of possible plans due to the need to compute the robot’s tracking ability. Furthermore, one must compensate for tracking error between the simplified planning model and the high-fidelity tracking model (i.e., the robot cannot perfectly track plans).
Both control intervention and planner-tracker approaches suffer from challenges in numerical representation due to the need to underapproximate the set of states that reach the target set and overapproximate those that avoids obstacles when guaranteeing safety and liveness in implementation. Implementations such as sums-of-squares polynomials [majumdar2017funnel, kousik2020bridging, landry2018reach], grid-based PDE solutions [chen2021fastrack], or polytope methods [kousik2019safe, althoff2021set] all introduce additional conservativeness purely due to compensating for numerical representation error. This challenge is exacerbated for systems with high-dimensional, nonlinear dynamics or observations; machine learning can address some of these concerns [bansal2021deepreach, michaux2023reachability, dawson2022safe, xiao2023barriernet, selim2022safe], at the expense of losing guarantees.
I-B Proposed Method Summary and Context
In this work, we leverage two key insights to perform goal-reaching near danger. First, we follow the literature [kousik2020bridging, kousik2019safe] in using a planner-tracker framework to reduce conservativeness by treating the planner and tracking controller as cooperative, as opposed to adversarial. Second, by representing our motion planner as a piecewise affine (PWA) system [christophersen2007optimal], we tightly approximate its reachable set, with minimal conservativeness from numerical approximation error. We demonstrate how this unlocks an efficient and less conservative approach to a variety of reach-avoid problems from the robotics literature, and provides the solution to a reachable set computation for drift parking vehicle dynamics. An overview and example of our approach is shown in Fig. 1.
Our method fits in the context of both control and planning. We build off of a long history of discrete-time PWA system analysis in optimal control (e.g., [thomas2006robust, kerrigan2002optimal, rakovic2005robust]), which we extend to the continuous-time reach-avoid problem for robot motion planning. Furthermore, our approach is complementary to sampling-based motion planners such as RRT [lavalle1998rapidly]; our approach could be applied in the future to compute a safety funnel around a sampled motion plan, or we could leverage sampling to seed our reachability analysis.
I-C Contributions
  1. 1.
    We propose an efficient, parallelizable method for Piecewise Affine Reach-avoid Computation (PARC) to underapproximate the Backwards Reach-Avoid Set (BRAS) of a planning model represented as a low-dimensional discrete-time, time-variant PWA system. Our method extends to discrete-time, time-variant affine systems as a particular case.
  2. 2.
    We propose a scheme to incorporate tracking error into PARC to compute a subset of the reach set and overapproximate the avoid set for a trajectory planning model. This allows us to represent the BRAS that compactly represents all trajectory plans that reach the goal and avoid obstacles when tracked by the robot, represented by a high-fidelity dynamical model.
  3. 3.
    We stress-test PARC on examples drawn from the safe control and planning literature: a 4-D TurtleBot, a 6-D planar quadrotor, a 10-D near-hover quadrotor, and a 13-D general quadrotor. We show that PARC outperforms state-of-the-art methods such as FaSTrack [chen2021fastrack], Neural CLBF [dawson2022safe], RTD [kousik2020bridging, kousik2019safe], and KDF [verginis2022kdf] in near danger reach-avoid tasks due to reduced conservativeness from the combination of planning model and set representation choice and better implementation of tracking error. We also demonstrate PARC on a simulated 6-D vehicle model to plan safe, extreme drift parallel parking maneuvers, and on a 4-D physical TurtleBot as a preliminary hardware experiment.
Paper Organization: Next, we introduce mathematical objects used throughout our formulation (Section II). We then present our problem formulation (Section III) and proposed approach (Section IV). We assess PARC with numerical experiments (Section V), then demonstrate a drift parking extreme vehicle dynamics example in simulation, extending beyond the state of the art in reachability-based safe motion-planning and control methods (Section VI). We also showcase PARC’s ability to extend to real robots with a preliminary hardware experiment (Section VII). Finally, we discuss takeaways and limitations in Section VIII. Note, we have extensive appendices, and call out key results in the text.
II Preliminaries
We now introduce our notation conventions and define H-polytopes and PWA functions, which we use to represent sets and dynamics of planning model throughout the paper. We then present a one-step Backward Reachable Set (BRS) computation that is the core of our proposed PARC method.
II-A Notation
The real numbers are $\mathbb{R}$ and the natural numbers are $\mathbb{N}$. Scalars are lowercase italic and sets are in uppercase italic (e.g., $a\in A\subset\mathbb{R}^{n}$). Vectors are lowercase bold, and matrices are uppercase bold (e.g., $\mathbf{y}=\mathbf{A}\mathbf{x}$). An $n\times m$ matrix of zeros is $\mathbf{0}_{n\times m}$, and an $n\times n$ identity matrix is $\mathbf{I}_{n}$. The $p^{\mathrm{\textnormal{th}}}$ to $q^{\mathrm{\textnormal{th}}}$ dimensions of vector $\mathbf{x}$ and the submatrix of $\mathbf{A}$ defined by row dimensions $p_{1}$:$q_{1}$ and column dimensions $p_{2}$:$q_{2}$ are denoted as $(\mathbf{x})_{p:q}$ and $(\mathbf{A})_{p_{1}:q_{1},p_{2}:q_{2}}$, respectively.
II-B H-Polytopes
In this work, we represent all sets as H-polytopes or unions of H-polytopes, which have extensive algorithm and toolbox support [herceg2013multi, althoff2015introduction]. An ${n}$-dimensional H-polytope $\mathcal{P}(\mathbf{A},\mathbf{b})\subset\mathbb{R}^{{n}}$ is a convex set parameterized by ${n}_{\mathrm{\textnormal{h}}}$ linear constraints $\mathbf{A}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}}$ and $\mathbf{b}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}}$:
$\mathcal{P}(\mathbf{A},\mathbf{b})$ =$=\left\{\mathbf{x}\ |\ \mathbf{A}\mathbf{x}\leq\mathbf{b}\right\}.$ (1)
Unless otherwise stated, we assume all H-polytopes are compact.
Consider a pair of H-polytopes $P_{1}=\mathcal{P}(\mathbf{A}_{1},\mathbf{b}_{1})$, $P_{2}=\mathcal{P}(\mathbf{A}_{2},\mathbf{b}_{2})$. We make use of intersections $P_{1}\cap P_{2}$, Minkowski sums ($P_{1}\oplus P_{2}=\left\{\mathbf{x}+\mathbf{y}\ |\ \mathbf{x}\in P_{1},\mathbf{y}\in P_{2}\right\}$), Pontryagin differences ($P_{1}\ominus P_{2}=\left\{\mathbf{x}\in P_{1}\ |\ \mathbf{x}+\mathbf{y}\in P_{1}\ \forall\ \mathbf{y}\in P_{2}\right\}$), set differences $P_{1}\setminus P_{2}$, Cartesian products $P_{1}\times P_{2}$, and convex hulls $\mathrm{\textnormal{conv}}\!\left(P_{1},P_{2}\right)$. We also project an $n$-dimensional polytope $P$ into dimensions $p:q$ as $\mathrm{\textnormal{proj}}_{{p}:{q}}\!\left(P\right)$, and slice it in the dimensions $p:q$ with respect to some constant $\mathbf{x}_{0}\in\mathbb{R}^{q-p+1}$ as $\mathrm{\textnormal{slice}}_{{p}:{q}}\!\left(P,\mathbf{x}_{0}\right)$. These operations are detailed in Appendix -B.
II-C PWA Systems
We consider discrete-time, time-variant PWA functions to enable our reach-avoid analysis. We denote a PWA function as $\psi:X_{0}\to X_{t_{\mathrm{\textnormal{f}}}}\subseteq\mathbb{R}^{{n}}$ with output $\mathbf{x}(t_{\mathrm{\textnormal{f}}})=\mathbf{\psi}\!\left(\mathbf{x}(0)\right)$ given an input $\mathbf{x}(0)=\mathbf{x}_{0}\in X_{0}\subset\mathbb{R}^{{n}}$. The system has state $\mathbf{x}(t)\in X_{t}\subset\mathbb{R}^{{n}}$ at time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which is the solution of difference equation of dynamics:
$\mathbf{x}(t+\Delta t)=\mathbf{C}_{\ell,t}\mathbf{x}(t)+\mathbf{d}_{\ell,t},$ (2)
where $\mathbf{C}_{\ell,t}\in\mathbb{R}^{{n}\times{n}}$ and $\mathbf{d}_{\ell,t}\in\mathbb{R}^{{n}}$ define the dynamics of the $\ell^{\mathrm{\textnormal{th}}}$ PWA region. The index $\ell$ is given by
$\ell$ $=\min\{i\ |\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\},$ (3)
where $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$ is the $i^{\mathrm{\textnormal{th}}}$ of the ${n}_{\mathrm{\textnormal{PWA}},t}$ PWA regions at time $t$. We say that $\mathbf{x}(t)$ has the mode $s_{t}=\ell$ at time $t$.
In addition, if the mode of $\mathbf{x}(0),\mathbf{x}(\Delta t),\cdots,\mathbf{x}(t_{\mathrm{\textnormal{f}}}-\Delta t)$ is $s_{0},s_{\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t}$ for some input $\mathbf{x}(0)\in X_{0}$, we denote the mode sequence for $\mathbf{x}(0)$ as:
$S=(s_{0},s_{\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t}).$ (4)
For all $\mathbf{x}(0)$ that share the same mode sequence, the dynamics reduce from time-variant PWA to time-variant affine up to time $t_{\mathrm{\textnormal{f}}}$. Mathematically, the discrete-time, time-variant affine dynamics with respect to the mode sequence (4) is:
$\begin{split}&\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\\ &\forall\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{s_{t},t},\mathbf{b}_{s_{t},t}),\end{split}$ (5)
where $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
The reduction of a PWA system to a specific affine dynamics is the core idea that enables the efficient underapproximating BRS computation of PARC. Note, taking advantage of fixed mode sequences is well-established in robotics and controls (e.g., [thomas2006robust, desimini2020robust, kerrigan2002optimal, wensing2023optimization]), but we believe we are the first to show its utility in safe reach-avoid computation of robot planning models.
We require that the interior of each PWA region does not intersect with other regions, such that any two regions are either disjoint or intersect only at their boundaries:
$\left\{\mathbf{x}\ |\ \mathbf{A}_{i}(t)\mathbf{x}<\mathbf{b}_{i}(t)\right\}\cap\mathcal{P}(\mathbf{A}_{j}(t),\mathbf{b}_{j}(t))=\emptyset,$ (6)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and $i,j\in\{1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\}$, $i\neq j$. Thus, the $\min$ in (3) dictates the dynamics’ behavior at the intersecting subsets of boundaries.
Note that, for $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define:
$X_{t}=\bigcup_{i=1}^{{n}_{\mathrm{\textnormal{PWA}},t}}\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),$ (7)
such that the input to $\psi$ is restricted to $\{\mathbf{x}_{0}\ |\ \mathbf{x}_{t}\in X_{t},t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$. Thus, a PWA function $\psi$ can be fully defined and parameterized by some given $\mathbf{C}_{i,t}$, $\mathbf{d}_{i,t}$, $\mathbf{A}_{i,t}$, and $\mathbf{b}_{i,t}$ for each $i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}$ and $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
As a particular case of this problem formulation, when ${n}_{\mathrm{\textnormal{PWA}},t}=1$ for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, one obtains the discrete-time, time-variant affine system (LABEL:eq:time_variant_affine) from the PWA system (2) and (3) automatically, as $S=(1,\cdots,1)$ always.
Finally, to ensure safety for realistic systems, we must extend our system definition from discrete to continuous time. We do so with linear interpolation: at each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, $\mathbf{x}(t^{\prime})$ for $t<t^{\prime}<t+\Delta t$ is defined by
$\mathbf{x}(t^{\prime})=\mathbf{x}(t)+(\mathbf{x}(t+\Delta t)-\mathbf{x}(t))(\tfrac{t^{\prime}-t}{\Delta t}).$ (8)
As will be shown in Section IV-C, the use of linear interpolation to bridge between discrete and continuous time is necessary to enable avoid set computation with H-polytopes. Similar reachability techniques to ours have been shown to work with Bernstein polynomials [michaux2023can], but only for forward reachability, to the best of our knowledge. Nonetheless, it may still be possible for PARC to handle alternative interpolation methods such as cubic splines or Bézier curves by using other set representations such as polynomial zonotopes [kochdumper2020sparse]. In this paper, we opted for the most straightforward approach of linear interpolation and leave the investigation with other interpolation methods as future work.
II-D One Step BRS via Inverse Affine Map
Finally, we take advantage of the H-polytope analytic solution for inverse affine maps to enable BRS computation using the following well-known lemma (c.f., [thomas2006robust, vincent2021reachable, herceg2013multi]).
Lemma 1.
Consider an affine system $\mathbf{x}(t+\Delta t)=\mathbf{C}\mathbf{x}(t)+\mathbf{d}$. Define the one-step BRS as the map $\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)=\left\{\mathbf{x}|\ \mathbf{A}\mathbf{x}_{next}\leq\mathbf{b},\ \mathbf{x}_{next}=\mathbf{C}\mathbf{x}+\mathbf{d}\right\}$. This one-step BRS is in fact an H-polytope:
$\begin{split}\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)=\mathcal{P}(\mathbf{A}\mathbf{C},\mathbf{b}-\mathbf{A}\mathbf{d}).\end{split}$ (9)
Proof.
$\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)$ =$=\left\{\mathbf{x}|\ \mathbf{A}(\mathbf{C}\mathbf{x}+\mathbf{d})\leq\mathbf{b}\right\},$ (10)
$=\mathcal{P}(\mathbf{A}\mathbf{C},\mathbf{b}-\mathbf{A}\mathbf{d}).$ (11)
We take advantage of the fact that repeated applications of Lemma 1 to an H-polytope still produce an H-polytope. Notice that (9) does not require invertibility of $\mathbf{C}$, conferring an advantage over other set polytope representations such as constrained zonotopes [scott2016constrained]. This makes H-polytopes an excellent choice for representing the BRS.
III Problem Statement
Our overall goal is to design a reach-avoid trajectory using a simple planning model, which enables computational efficiency while also accounting for the tracking error when the plan is tracked by a robot, represented by a tracking model. We now describe each of these models and the problem we seek to solve.
III-A Planning Model
A planning model provides a low-fidelity but easy-to-compute model of the robot’s dynamics. In this work, we use parameterized reference trajectories, which can allow quick computation and strong safety guarantees [kousik2020bridging, kousik2019safe, hess2014formal].
Before defining the model, we define a workspace state $\mathbf{w}\in W\subset\mathbb{R}^{{n}_{w}}$ (i.e., $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$) that represents the position of the robot. The planning state which we denote as $\mathbf{p}\in P\subset\mathbb{R}^{{n}_{p}}$ is chosen to include the workspace state, such that $\mathbf{p}=[\mathbf{w}^{\intercal},\mathbf{p}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$ where $\mathbf{p}_{\mathrm{\textnormal{other}}}\in P_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{{n}_{p}}_{\mathrm{\textnormal{other}}}}$ represents states such as heading, velocity, etc.
We then define the planning model as
$\dot{\mathbf{p}}(t)=\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}),\quad\mathbf{p}(0)=\mathbf{p}_{0}$ (12)
where $\mathbf{p}_{0}$in $P\subset\mathbb{R}^{{n}_{p}}$ defines the initial planning state, $\mathbf{k}\in K\subset\mathbb{R}^{{n}_{k}}$ represents a trajectory parameter, and $t\in[0,t_{\mathrm{\textnormal{f}}}]$ is the time horizon. A solution $\mathbf{p}:[0,t_{\mathrm{\textnormal{f}}}]\rightarrow P$ is referred to as a plan, or equivalently, reference trajectory. For a given initial planning state and trajectory parameter, the plan at time $t$ is $\mathbf{p}(t;\mathbf{p}_{0},\mathbf{k})$. We restrict the class of planning models to enable computing reachable sets for parameterized planning models:
Assumption 2 (Extended Translation Invariance (ETI) in Workspace).
Changing the workspace state either does not affect the planning dynamics $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ in the workspace or else the dynamics are constant:
$\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\mathbf{f}_{\mathbf{w}}\right\|_{2}=0\quad\mathrm{\textnormal{and}}\quad\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{p}}\right\|_{\mathrm{\textnormal{F}}}=0,$ (13)
where $\mathbf{f}_{\mathbf{w}}=({\mathbf{f}_{\mathrm{\textnormal{plan}}}})_{1:{n}_{w}}$ and $\left\|\cdot\right\|_{\mathrm{\textnormal{F}}}$ indicates the Frobenius norm.
This is not overly restrictive; a variety of ETI planning models are shown in Section V. The need for the dynamics to be constant in some states is because we treat $\mathbf{k}$ as states with 00 dynamics to formulate our reachability method.
III-B Tracking Model
The tracking model is a high-fidelity model of robot dynamics, denoted
$\dot{\mathbf{z}}(t)=\mathbf{f}_{\mathrm{\textnormal{track}}}({\mathbf{z}(t)},{\mathbf{u}_{\mathrm{\textnormal{fb}}}(t)}),~{}~{}~{}\mathbf{z}(0)=\mathbf{z}_{0},$ (14)
where $\mathbf{z}_{0}\in Z\subset\mathbb{R}^{{n}_{z}}$ is the initial state. The tracking state $\mathbf{z}$ includes the planning state $\mathbf{p}$, i.e., $\mathbf{z}=[\mathbf{p}^{\intercal},\mathbf{z}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$. Given a plan $\mathbf{p}$ parameterized by $\mathbf{k}$, we use a feedback tracking controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}(t,\mathbf{z}(t);\mathbf{k})\in U$ (e.g., PID or MPC) to control the tracking model. We assume $Z$ and $U$ are compact and the dynamics are Lipschitz such that trajectories exist [grant2014theory]. For a particular $\mathbf{z}_{0}$, we denote the resulting, unique realized trajectory (i.e., solution to (14)) as $\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})$ at time $t$.
Remark 3.
Designing the feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ is not the focus of this work. Conveniently, specifying a parameterized (i.e., restricted) space of plans $\mathbf{p}$ simplifies controller design, as we find experimentally in Section V-D. Our overall goal is to compensate for the fact that the controller is imperfect.
III-C Reach-Avoid Problem
We now define our reach-avoid problem: find all plans for which the tracking model safely reaches a goal set at time $t_{\mathrm{\textnormal{f}}}$. We denote the goal $\overline{P_{\mathrm{\textnormal{G}}}}\subset W$ and obstacles $\overline{\mathcal{O}}\subset W$; the overline indicates these are restricted to the workspace, which we modify later to enable the proposed method.
Note, we only consider static obstacles, because our goal is to plan a robot’s motion near danger. The danger resulting from moving or adversarial obstacles depends upon accurately predicting obstacle motion, which introduces an independent variable unrelated to our purpose. We leave extension to dynamic scenes for future work, but we anticipate we can adapt reachability-based methods such as [vaskov2019towards, vaskov2019not].
Problem 4 (Backward Reach-Avoid Set (BRAS)).
Given the the planning model (12), tracking model (14), goal set $\overline{P_{\mathrm{\textnormal{G}}}}$, obstacles $\overline{\mathcal{O}}$, and a fixed time $t_{\mathrm{\textnormal{f}}}$, find the set of planning models states and trajectory parameters for which the robot reaches $\overline{P_{\mathrm{\textnormal{G}}}}$ exactly at time $t_{\mathrm{\textnormal{f}}}$ while avoiding collisions with obstacles $\overline{\mathcal{O}}$ :
$\begin{split}\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})=\{&(\mathbf{p}_{0},\mathbf{k})\in P\times K\ \mid\forall~{}\mathbf{z}_{0}~{}\mathrm{\textnormal{s.t.}}\\ &(\mathbf{z}_{0})_{1:{n}_{w}}=\mathbf{p}_{0},\\ &(\mathbf{z}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\in\overline{P_{\mathrm{\textnormal{G}}}},~{}\mathrm{\textnormal{and}}\\ &(\mathbf{z}(\tau;\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\notin\overline{\mathcal{O}}\quad\forall\tau\in[0,t_{\mathrm{\textnormal{f}}}]\}\end{split}$ (15)
Several remarks are in order to discuss our fixed time horizon, choice of goals and obstacles, and choice of tracking error model.
Remark 5 (Fixed Final Time).
A fixed time horizon is reasonable because we seek to compute a trajectory near danger; the robot is already near its goal, but reaching the goal is difficult. The BRAS is ideally a large set of initial conditions far from obstacles that can be used as a target set for other methods to reach safely.
Remark 6 (Goal and Obstacles).
We restrict the goal and obstacles to the workspace only to simplify exposition; we demonstrate more general goals and obstacles, such as in the robot’s orientation, in Sections V and VI.
Remark 7 (Disturbances to the System).
To simplify exposition, we treat tracking error as a lumped uncertainty representation between planning and control, encompassing multiple error sources – disturbances, state estimation, object detection, and trajectory prediction. Separately analyzing these sources would introduce confounding variables, complicating the interpretation of our results. Instead, we consolidate them into a single metric of tracking error. For example, one can design a robust tracking controller that yields a tracking error bound, robust to external disturbances and model uncertainties [chen2021fastrack, bechlioulis2014low]. Our planning level only considers the result of those errors as a tracking error bound without separately accounting for each. Alternatively, if the bounds of inertial uncertainties [michaux2023can], dynamic disturbance [selim2022safe, alanwar2023data], and perception uncertainty [kousik2020bridging] are known (which are strong assumptions), then planner-tracker frameworks similar to ours have been shown to maintain reach-avoid guarantees in the past by “buffering” the bounds into the tracking error [kousik2020bridging, vaskov2019towards, selim2022safe, chen2021fastrack]. In Section VII, we show the preliminary results of a hardware experiment that successfully includes localization drift from the inertial measurement unit (IMU) and the encoders.
III-D Example System: TurtleBot
We use the TurtleBot unicycle model as an illustrative example from the safe planning and control literature [chen2021fastrack, kousik2020bridging]. Our planning model is a Dubins car:
$\dot{\mathbf{p}}=\frac{d}{dt}\begin{bmatrix}p_{x}\\ p_{y}\\ \theta\end{bmatrix}=\begin{bmatrix}v_{\mathrm{\textnormal{des}}}\cos\theta\\ v_{\mathrm{\textnormal{des}}}\sin\theta\\ \omega_{\mathrm{\textnormal{des}}}\end{bmatrix},$ (16)
with workspace (position) states $\mathbf{w}=[p_{x},p_{y}]^{\intercal}$, other (heading) state $\mathbf{p}_{\mathrm{\textnormal{other}}}=\theta$, and trajectory parameters $\mathbf{k}=[v_{\mathrm{\textnormal{des}}},\omega_{\mathrm{\textnormal{des}}}]^{\intercal}$. Note that a Dubins car model (16) is ETI in the workspace (Assum. 2) since its dynamics in workspace dimension is translation-invariant (i.e. $\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\right\|_{2}=0$.)
We use the following tracking model:
$\dot{\mathbf{z}}=\frac{d}{dt}\begin{bmatrix}p_{x}\\ p_{y}\\ \theta\\ v\end{bmatrix}=\begin{bmatrix}v\cos\theta\\ v\sin\theta\\ u_{\omega}\\ u_{v}\end{bmatrix},$ (17)
where $\mathbf{z}_{\mathrm{\textnormal{other}}}=[\theta,v]^{\intercal}$ are heading and linear velocity, and $\mathbf{u}_{\mathrm{\textnormal{fb}}}=[u_{\omega},u_{v}]^{\intercal}$ is implemented as an iterative LQR tracking controller [kong2021ilqr] that commands angular velocity and linear acceleration.
We define a goal set as a unit box, $\overline{P_{\mathrm{\textnormal{G}}}}=\{\mathbf{w}\mid p_{x},p_{y}\in[-1,1]\}$, and an obstacle near the goal, $\overline{\mathcal{O}}=\{\mathbf{w}\mid p_{x}\in[-1.75,-1.25],p_{y}\in[-0.25,0.25]\}$ (see Figures 2 and 3).
IV Proposed Method
In this section, we detail our approach to solving the BRAS problem. First, we convert the planning model to a PWA system to enable our Piecewise Affine Reach-Avoid Computation (PARC). We then adapt the method in [thomas2006robust] to compute a subset of the reach set of the PWA planning model without tracking error, to simplify exposition. Next, we propose a novel approach to overapproximate the continuous time avoid set of the PWA system without tracking error (Key result: Theorem 13). To enable incorporating tracking error between the planned and realized trajectories, we adapt the methods in [kousik2019safe, kousik2020bridging] to estimate time-varying tracking error. Finally, we propose PARC to incorporate tracking error and compute the BRAS (Key result: Theorem 20), enabling provably-safe goal-reaching.
IV-A Reformulation to Enable PARC
IV-A1 Augmented State, Goal, and Obstacles
Before proceeding, to enable PARC, we define an augmented planning state $\mathbf{x}(t)\in X=W\times K\times P_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{n}_{w}+{n}_{k}+{{n}_{p}}_{\mathrm{\textnormal{other}}}}$ at time $t$ as:
$\mathbf{x}(t)=\begin{bmatrix}\mathbf{w}(t)\\ \mathbf{k}\\ \mathbf{p}_{\mathrm{\textnormal{other}}}(t)\end{bmatrix}.$ (18)
We perform this reordering and stacking because, as we will see later in Lem. 12, the $\mathbf{w}$ and $\mathbf{k}$ are ETI states, so indexing them as the first states of $\mathbf{x}$ will simplify avoid set computation. We similarly augment the goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ and the obstacles $\overline{\mathcal{O}}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\overline{\mathcal{O}}_{i}$, which we assume are union of H-polytopes:
$\begin{split}P_{\mathrm{\textnormal{G}}}&=\overline{P_{\mathrm{\textnormal{G}}}}\times K\times P_{\mathrm{\textnormal{other}}}\\ \mathcal{O}_{i}&=\overline{\mathcal{O}}_{i}\times K\times P_{\mathrm{\textnormal{other}}}.\end{split}$ (19)
IV-A2 PWA Planning Model
We approximate the planning model $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ as a discrete-time, time-variant PWA to enable computation. First, we define linearization points $\mathbf{x}^{*}_{t,i}\in X_{t}$, $i=1,\cdots,{n}^{*}_{t}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which can be generated by gridding or uniform sampling. We define the PWA regions as the Voronoi partition of the linearization points [casselman2009new] (see Appendix -C). Each $i$th region is an H-polytope $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$. Then, in each region, at each time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define the dynamics $\mathbf{C}_{i,t}$ and $\mathbf{d}_{i,t}$ by Taylor expanding about the linearization points:
$\mathbf{x}(t+\Delta t)=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})+\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{\mathbf{x}={\mathbf{x}^{*}_{i,t}}}(\mathbf{x}(t)-\mathbf{x}^{*}_{i,t}),$ (20a)
$\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))=\begin{bmatrix}(\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{1:{n}_{w}}\\ \mathbf{k}\\ (\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{{n}_{w}+1:{n}_{p}}\\ \end{bmatrix}$ (20b)
$\mathbf{C}_{i,t}=\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}},\quad\mathrm{\textnormal{and}}$ (20c)
$\mathbf{d}_{i,t}=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})-\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}}\mathbf{x}^{*}_{i,t}.$ (20d)
Depending on the planning model, there may be some PWA regions with identical dynamics. If the union of the regions is convex, we combine them into a single PWA region (note, the linearization points are discarded).
From this point onwards, we will be using the PWA system defined by (20) as the planning model instead of (12). That is, the tracking error, BRAS, and reach-avoid guarantees will be computed with respect to the tracking model (14) and the planning model (20). Thus, PARC is concerned with the tracking error between (14) and (20), but not the error induced by affinizing (12) to (20).
Remark 8.
Choosing a larger $\Delta t$ would induce higher mismatch between the tracking model and the planning model, thus increasing the conservativeness of the method. In scenarios with highly dynamic systems or cluttered environments, selecting a smaller $\Delta t$ reduces conservatism, as demonstrated in Fig. 4. As will be shown in the rest of this section, PARC’s reach-avoid guarantees will still be maintained regardless of the choice of $\Delta t$, and its computation time is inversely proportional to $\Delta t$.
IV-B PARC: Reach Set without Tracking Error
Without accounting for tracking error, the reach set of a planning model is the set of initial conditions $\mathbf{x}_{0}$ of trajectories that are guaranteed to reach the goal region $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$. Instead of naively computing the full reach set, which scales exponentially in memory, computation time, and a number of set representations with the number of timesteps, PARC is planning aware: given an expert plan, we compute only a local, underapproximated BRS efficiently for the neighborhood that shares the same mode sequence as the expert plan. Fixing the mode sequence reduces the PWA system to a time-variant affine system, so the BRS is a single H-polytope using the inverse affine map of the goal (see (9)), with computation time scaling linearly with the number of timesteps and quadratically with dimensions from matrix multiplication.
Assumption 9 (Expert Plan w/o Tracking Error).
We have access to at least one goal-reaching initial condition $\mathbf{x}_{0}$ for which the PWA system reaches the goal without necessarily avoiding obstacles:
$\mathbf{\psi}\!\left(\mathbf{x}_{0}\right)\in P_{\mathrm{\textnormal{G}}}.$ (21)
Obtaining expert plans is system-dependent, but is not difficult, because we do not require the expert plan to avoid obstacles. Thus, one option is to solve a two-point boundary value nonlinear MPC problem, for which solution guarantees exist [bonalli2019gusto]. In practice, it is often sufficient to uniformly sample trajectory parameters $\mathbf{k}$ and roll out the PWA planning model, since the robot starts out near the goal. This assumption is even easier to fulfill if the planning model is time-variant affine (which we use for a general 3D quadrotor and drift vehicle planning in Section V, VI) since plans have only one mode.
If Assumption 9 is satisfied, we can obtain $\mathbf{x}_{0}$’s mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$ using (3). That is,
$s_{t}$ $=\min\{i\ |\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\},$ (22)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$. Then, we can compute the exact $t_{\mathrm{\textnormal{f}}}$-time BRS using the method in [thomas2006robust]:
Proposition 10 (Reach Set w/o Tracking Error).
Consider a mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Then the exact $t_{\mathrm{\textnormal{f}}}$-time BRS $\overline{\Omega}_{0}$ of the goal for the affine dynamics associated with $S$ is
$\begin{split}\overline{\Omega}_{t_{\mathrm{\textnormal{f}}}}&=P_{\mathrm{\textnormal{G}}},\\ \overline{\Omega}_{t}&=\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)\end{split}$ (23)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
Proof.
This follows directly from Lemma 1. ∎
This means the reference trajectories generated from any augmented planning state within $\overline{\Omega}_{0}$ are guaranteed to reach $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$ with mode sequence $S$. This process, along with an expert plan, is illustrated in Fig. 2.
Traditionally, methods such as the reachableSet function in MPT3 [herceg2013multi] compute the full BRS of the PWA system instead of that of only a particular mode sequence. While computing the full BRS would return all possible reachable states and do not require knowledge of an expert plan, the computation time grows exponentially with the number of BRS steps and quickly becomes intractable due to the large number of polytopes, as shown in Table I. Instead, Proposition 10 boasts a much more reasonable computation time by isolating the affine dynamics of a single mode sequence and keeping only one polytope regardless of the number of BRS steps.
Figure 2: (Right) Visualization of the reach set computation in Section IV-B for the TurtleBot example (Section III-D), projected onto the planning state space $P$, when $\mathbf{p}_{0}=[-4,0,\pi/5]^{\intercal}$, $t_{\mathrm{\textnormal{f}}}=4(s)$. The goal set is the green polytope extended to all $\theta\in[-\pi,\pi]$. An expert plan, computed by solving a two-point boundary value problem, passes through two modes (blue and magenta line). We compute the corresponding $t_{\mathrm{\textnormal{f}}}$-time BRS $\overline{\Omega}_{0}$ (blue polytope with solid outline). Intermediate BRSs $\overline{\Omega}_{t}$ are the polytopes without outlines, colored by mode. (Left) Illustration of the trajectory parameter space; the grey region represents $K$, the green region represents the projection of $\overline{\Omega}_{0}$ on $K$. After reach-set computation, the state $\mathbf{x}_{\mathrm{\textnormal{sample}}}=[\mathbf{p}_{\mathrm{\textnormal{sample}}}^{\intercal},\mathbf{k}_{\mathrm{\textnormal{sample}}}^{\intercal}]^{\intercal}$ is sampled from $\overline{\Omega}_{0}$: black dot (right) denotes the $\mathbf{p}_{\mathrm{\textnormal{sample}}}$ and black star (left) denotes the $\mathbf{k}_{\mathrm{\textnormal{sample}}}$. The resulting trajectory is denoted as the black line which reaches the goal.
Computation Time (s) Number of Polytopes
BRS Steps MPT3 Prop. 10 MPT3 Prop. 10
1 0.03 0.0015 16 1
2 0.27 0.0022 100 1
3 1.45 0.0038 484 1
4 6.63 0.0036 2116 1
5 29.46 0.0027 8836 1
26 timeout 0.0155 1
TABLE I: Computation time and number of polytopes for 1, 2, 3, 4, 5, and 26-step BRS for the TurtleBot example (Section III-D) using the reachableSet function in MPT3 [herceg2013multi] and Proposition 10. Since MPT3 computes the full BRS of the PWA system, the computation time quickly grows intractable due to the large number of polytopes. In contrast, Proposition 10 maintains a reasonable computation time by keeping only one polytope at each step, characterized by the affine dynamics corresponding to the mode sequence of the expert plan.
IV-C PARC: Avoid Set without Tracking Error
Without accounting for tracking error, the avoid set is the set of pairs of initial planning model state and trajectory parameter such that any plans generated from outside the avoid set are guaranteed to not collide with the obstacles at all continuous time before $t_{\mathrm{\textnormal{f}}}$. Recall that, per (8), we model the continuous-time behavior as straight line segments between the discrete states of the PWA system.
Intuitively, the avoid set identifies all trajectories where the straight lines connecting the PWA planning model’s discrete states collide with any obstacles:
Proposition 11 (Unsafe Plans w/o Tracking Error).
Consider the obstacle $\mathcal{O}_{i}$, some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Suppose we compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ and the $(t_{\mathrm{\textnormal{f}}}-t-\Delta t)$-time BRS $\overline{\Omega}_{t+\Delta t}$ as in (23). If any trajectories between $\overline{\Omega}_{t}$ and $\overline{\Omega}_{t+\Delta t}$ collide with $\mathcal{O}_{i}$, that is, if $\exists\ \mathbf{x}(t)\in\overline{\Omega}_{t},\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\in\overline{\Omega}_{t+\Delta t}$ such that
[…]
then
$\mathcal{O}_{i}\cap\mathrm{\textnormal{conv}}\!\left(\overline{\Omega}_{t},\overline{\Omega}_{t+\Delta t}\right)\neq\emptyset.$ (25)
Proof.
By the definition of convex-hull, we have
[…]
$\mathbf{x}(t),\mathbf{x}(t+\Delta t)\in\overline{\Omega}_{t}\bigcup\overline{\Omega}_{t+\Delta t}\},$ (26a)
[…]
$\mathbf{x}(t)\in\overline{\Omega}_{t},\mathbf{x}(t+\Delta t)\in\overline{\Omega}_{t+\Delta t}\}.$ (26b)
Thus (24) necessarily implies (25). ∎
Next, before we can overapproximate the avoid set between two-time steps, we must adapt our notion of extended translation invariance to the PWA planning model:
Lemma 12 (PWA ETI).
Suppose $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in the workspace as in Assum. 2. Also suppose we convert it to a PWA system as in Section IV-A2 and (20). Then in mode $i$ at time $t$, each state $\mathbf{x}$ of the PWA system obeys
$\left\|\hat{\mathbf{C}}_{i,t}^{W}{(\hat{\mathbf{C}}_{i,t}\mathbf{x}+\mathbf{d}_{i,t})_{1:{n}_{w}}}\right\|_{\mathrm{\textnormal{2}}}=0\quad\forall\mathbf{x}\in{\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})}$ (27)
where $\hat{\mathbf{C}}_{i,t}=\mathbf{C}_{i,t}-\mathbf{I}_{{n}_{x}}$ and $\hat{\mathbf{C}}^{W}_{i,t}$ denotes the ${n}_{w}\times{n}_{w}$ upper-left submatrix of $\hat{\mathbf{C}}_{i,t}$, namely $\hat{\mathbf{C}}_{i,t}^{W}\triangleq(\hat{\mathbf{C}}_{i,t})_{1:{n}_{w},1:{n}_{w}}$
Proof.
This follows directly from Assum. 2 and (20). ∎
The definition of PWA-ETI (27) in other subspace is similarly defined by substituting $W$ to the subspace of interest. Note that $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in $W\times K$ since the trajectory parameter $\mathbf{k}\in K$ is constant.
Now, for avoid-set computation, it is helpful to relabel the ETI states of $\mathbf{x}(t)$, identified per Lemma 12:
$\mathbf{x}(t)$ $=\begin{bmatrix}\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\\ \mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\end{bmatrix},$ (28)
where $\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\in X_{\mathrm{\textnormal{ETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}$ are the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ states in $\mathbf{x}(t)$ that are ETI states, and $\mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\in X_{\mathrm{\textnormal{nonETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}$ are the remaining states in $\mathbf{x}(t)$. Thus ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}\geq{n}_{w}+{n}_{k}$. Note that all sets and matrices including $\mathcal{O}$, and $P_{\mathrm{\textnormal{G}}}$ are relabeled accordingly. This reordering is only to simplify the exposition.
We are now ready to state our first key contribution: overapproximating the avoid set without tracking error. The following theorem overapproximates the continuous-time BRS of an obstacle and intersects it with the discrete-time reach-set of the PWA planning model to identify unsafe plans.
Theorem 13 (Avoid Set w/o Tracking Error).
Consider the obstacle $\mathcal{O}_{i}$, some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and some mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Suppose we compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ as in (23). Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\mathcal{O}_{i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$, and define the intermediate avoid set $\overline{\Lambda}_{i,t,t}\subset\overline{\Omega}_{t}$ as:
$\overline{\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{nonETI}}},\mathcal{O}_{i}\right)\cap\overline{\Omega}_{t}$ (29)
Then, for any $\mathbf{x}(t)\in\overline{\Omega}_{t}\setminus\overline{\Lambda}_{i,t,t}$, $\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}$,
[…]
Proof.
See Appendix -D. ∎
In summary, Proposition 11 establishes a necessary condition for the collision to occur between two discrete timesteps, while Theorem 13 overapproximates the avoid sets contained in each intermediate BRS.
Remark 14.
We note that PARC can be readily extended to moving obstacles by replacing $\mathcal{O}_{i}$ in (29) with the over-approximated swept volume of the obstacle between $t$ and $\Delta t$, assuming its motion is known or can be over-approximated. This strategy has shown to be effective on other set-based planner-tracker frameworks [vaskov2019towards, vaskov2019not]. For ease of exposition, we consider only static obstacles in this paper, as many of the methods we are comparing PARC against do not handle moving obstacles [chen2021fastrack, dawson2022safe, verginis2022kdf], and reach-avoid for static obstacles in near-danger setups is already a difficult problem for existing literature.
We now propose a straightforward scheme to construct the avoid set using Proposition 11 and Theorem 13: reference trajectories generated from outside the avoid set are guaranteed to avoid the obstacle. First, for each timestep $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and for each obstacle $i=1,\cdots,{n}_{\mathcal{O}}$, we check if (25) is true; this reduces conservativeness because we identify obstacles for which we do not have to compute the avoid set $\overline{\Lambda}_{i,t,t}$ with Theorem 13, at the expense of additional convex hull computations. That is, one can speed up computation but potentially make the avoid set more conservative by skipping checking (25). Then, we apply the following corollary:
Corollary 15 (Avoid Set Computation).
We compute the avoid set $\overline{\Lambda}_{i,t,0}$ by computing the $t$-time BRS of $\overline{\Lambda}_{i,t,t}$. Mathematically,
$\overline{\Lambda}_{i,t,j}$ $=\mathcal{B}\!\left(\overline{\Lambda}_{i,t,j+\Delta t},\mathbf{C}_{s_{j},j},\mathbf{d}_{s_{j},j}\right)\ \forall j=0,\Delta t,\cdots,t-\Delta t.$ (31)
Then, the final, overapproximated avoid set $\overline{\Lambda}$ can be stored as a union of H-polytopes as:
[…]
Proof.
This follows as a consequence of Theorem 13 and since the BRS operation, $\mathcal{B}$ is exact for an affine system defined for a mode sequence (Lemma 1). ∎
To sample from within the reach set but not from the avoid set, one can either rejection-sample, sample via random-walk algorithms [kannan2009random, chen2018fast], or compute the set difference between the reach set and the avoid set [baotic2009polytopic], and then sample via solving an LP. An example of the reach set and the avoid set are shown in Fig. 3; notice that the avoid set is a nonconvex union of polytopes.
Figure 3: (Left) Visualization of the reach set (green) and the avoid set (red) computation from Section IV-C for the TurtleBot example (Section III-D), sliced in the first three dimensions with respect to the initial condition $\mathbf{p}_{0}=[-3.5,-0.5,\pi/5]^{\intercal}$. The black star denotes the expert plan with $\mathbf{k}=[-0.318,0.9]^{\intercal}$ and the blue star denotes a safe plan with $\mathbf{k}=[-0.320,1.1]^{\intercal}$. (Right) Illustration of the resulting trajectory in $p_{x}$ and $p_{y}$, with the expert plan shown as a dashed black line and the safe plan shown as a dashed blue line. While the expert plan collides with the obstacle (red), PARC is still able to compute a safe plan with the same mode sequence that safely reach the goal set (green).
Figure 4: Visualization of the reach set (green), the avoid set (red), and a grid-based over-approximation of the true avoid set (yellow) for the TurtleBot example (Section III-D) for different values of the timestep $\Delta t$, with the same problem setup and slicing as Fig. 3. The grid-based over-approximation was obtained by dividing the reach set into a $50\times 50$ grid, and checking if the trajectories from each grid collides with the obstacle using FRS computation [herceg2013multi] and Proposition 11. Both the reach set and the numerical overapproximation error gets smaller as $\Delta t$ decreases.
Disregarding tracking error and perception uncertainties, which all planning and control methods suffer from, PARC’s only source of conservativeness (from computing the BRAS with respect to the expert plan’s mode sequence) comes from the convex hull approximation for linearly interpolated trajectories and the projection operation to account for non-ETI states. We illustrate the extent of this overapproximation in Fig. 4. Note that the amount of overapproximation depends heavily on the system, the problem setup, and the expert plan, but can generally be alleviated by choosing smaller timesteps at the expense of computation time.
IV-D Estimating Tracking Error
Now that we have a basis for computing the BRAS, we will model tracking error such that, when incorporated into the PARC framework, we can produce plans that, when followed by the tracking model, guaranteed to reach the goal region while avoiding all obstacles.
Our method is based on [kousik2019safe, shao2021reachability]. We sample $(\mathbf{z}_{0,j},\mathbf{k}_{j})\in\tilde{Z}\times\tilde{K}\subset Z\times K$ for $j=1,\cdots,{n}_{\mathrm{\textnormal{sample}}}$ uniformly, where $\tilde{Z}\times\tilde{K}$ is a hyperrectangle created by partitioning $Z\times K$; we do not index this partition to ease notation. We then have the robot start at $\mathbf{z}_{0,j}$ to track the plans parameterized by initial planning state $\mathbf{p}_{0,j}\triangleq(\mathbf{z}_{0,j})_{1:{n}_{p}}$ and trajectory parameter $\mathbf{k}_{j}$, and compute the error between the planned and realized trajectories in the workspace states (recall that we assume the goal and obstacles are subsets of workspace).
We consider two types of errors: the maximum final error $e_{t_{\mathrm{\textnormal{f}}},i}\in\mathbb{R}$ for goal-reaching, and maximum interval error $\tilde{e}_{t,i}\in\mathbb{R}$ for obstacle avoidance, defined for the discrete-time at time $t$ and the $i^{\mathrm{\textnormal{th}}}$ dimension of $\mathbf{z}$ and $\mathbf{p}$. PARC require two types of errors because it provides continuous-time guarantees for obstacle avoidance, but gives guarantees at the final timestep only for goal-reaching.
For each workspace coordinate $i=1,\cdots,{n}_{w}$, we define the maximum final error $e_{t_{\mathrm{\textnormal{f}}},i}$ as:
$e_{t_{\mathrm{\textnormal{f}}},i}=\max_{j}\left|\mathbf{p}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{p}_{0,j},\mathbf{k}_{j})-\mathbf{z}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0,j},\mathbf{k}_{j})\right|,$ (33)
where $\mathbf{p}_{i}(\cdot)$ and $\mathbf{z}_{i}(\cdot)$ are the $i^{\mathrm{\textnormal{th}}}$ element of $\mathbf{p}(\cdot)$ and $\mathbf{z}(\cdot)$. In plain words, $e_{t,i}$ represents the maximum error observed at the final time $t_{\mathrm{\textnormal{f}}}$ for the $i^{\mathrm{\textnormal{th}}}$ workspace dimension.
For $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, $i=1,\cdots,{n}_{w}$, we define the maximum interval error $\tilde{e}_{t,i}$ as:
$\tilde{e}_{t,i}=\max_{j,\ t^{\prime}\in[t,t+\Delta t]}\left|\mathbf{p}_{i}(t^{\prime};\mathbf{p}_{0,j},\mathbf{k}_{j})-\mathbf{z}_{i}(t^{\prime};\mathbf{z}_{0,j},\mathbf{k}_{j})\right|$ (34)
which is the maximum error observed between each discrete timestep for the $i^{\mathrm{\textnormal{th}}}$ workspace dimension.
We assume this approach provides an upper bound for feasible tracking error, which can be proven for simple planning models [kousik2019safe, shao2021reachability].
Assumption 16 (Maximum Sampled Error Contains Maximum Error).
We assume that, for a large enough ${n}_{\mathrm{\textnormal{sample}}}$,
$e_{t_{\mathrm{\textnormal{f}}},i}$ $\geq\max_{\mathbf{z}_{0}\in\tilde{Z},\,\mathbf{k}\in\tilde{K}}\left|\mathbf{p}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{p}_{0},\mathbf{k})-\mathbf{z}_{i}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0},\mathbf{k})\right|,$ (35a)
$\tilde{e}_{t,i}$ $\geq\max_{\mathbf{z}_{0}\in\tilde{Z},\,\mathbf{k}\in\tilde{K},\,t^{\prime}\in[t,t+\Delta t]}\left|\mathbf{p}_{i}(t^{\prime};\mathbf{p}_{0},\mathbf{k})-\mathbf{z}_{i}(t^{\prime};\mathbf{z}_{0},\mathbf{k})\right|,$ (35b)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, where $\mathbf{p}_{0}=(\mathbf{z}_{0})_{1:{n}_{p}}$.
This assumption is clearly a potential weakness of the method, but we find that it holds in practice because the sampling process produces adversarial $(\mathbf{z}_{0},\mathbf{k})$ pairs. Furthermore, we can refine the tracking error estimate by increasing the fineness of the partition of $Z\times K$ into subsets $\tilde{Z}\times\tilde{K}$ [kousik2019safe]. Note there exist alternative methods to collect or construct tracking error functions in the literature [verginis2022kdf, chen2021fastrack]. If desired, they could be swapped out with (33) and (34) and PARC would provide guarantees with respect to the new tracking error functions instead. We demonstrate PARC’s modularity to the tracking error model in Section V-D.
To incorporate these errors into the geometric computation of PARC, we express them as the maximum final error set $E_{t_{\mathrm{\textnormal{f}}}}\subset\mathbb{R}^{{n}_{x}}$ and the maximum interval error set $\tilde{E}_{t}\subset\mathbb{R}^{{n}_{x}}$:
$E_{t_{\mathrm{\textnormal{f}}}}$ $=[-e_{t_{\mathrm{\textnormal{f}}},1},e_{t_{\mathrm{\textnormal{f}}},1}]\times\cdots\times[-e_{t_{\mathrm{\textnormal{f}}},{n}_{w}},e_{t_{\mathrm{\textnormal{f}}},{n}_{w}}]\times\underbrace{\{0\}\times\cdots\times\{0\}}_{({{n}_{p}}_{\mathrm{\textnormal{other}}}+{n}_{k})\mathrm{\textnormal{ times}}},$ (36)
$\tilde{E}_{t}$ $=[-\tilde{e}_{t,1},\tilde{e}_{t,1}]\times\cdots\times[-\tilde{e}_{t,{n}_{w}},\tilde{e}_{t,{n}_{w}}]\times\underbrace{\{0\}\times\cdots\times\{0\}}_{({{n}_{p}}_{\mathrm{\textnormal{other}}}+{n}_{k})\mathrm{\textnormal{ times}}}.$ (37)
It is straightforward to express these intervals as H-polytopes [herceg2013multi]. The Cartesian product with singletons of zeros is necessary to augment the states for Minkowski sum and Pontryagin difference operations with the obstacles and the goal.
We note that collection of the tracking error only requires forward simulation or rollouts of the system and knowledge of the input and output states, so our approach could still apply if the tracking model is a “black box” by adapting methods such as [selim2022safe, alanwar2023data]. In this paper, we consider “white-box” dynamics in our experiments to treat them as a control condition to compare against the literature. While the forward simulation and collision checking of trajectories may sometimes be computationally expensive, we consider the tradeoff acceptable as the tracking error of PARC is expected to be collected offline. An example of computed tracking error for a near-hover quadrotor dynamical model is shown in Fig. 9.
IV-E Incorporating Tracking Error into PARC
To conclude our proposed method, we incorporate tracking error into PARC’s BRAS computation, thereby transferring guarantees for the planning model to the tracking model. In short, we get reach-avoid guarantees that incorporate tracking error by simply Minkowski summing the error with obstacles and Pontryagin differencing from the goal set.
IV-E1 PARC: Reach Set with Tracking Error
From Assumption 16, we assume the sampled tracking error is valid only over the compact hyperrectangle $\tilde{Z}\times\tilde{K}$. Thus, we must define a valid input region $\tilde{X}\subset X$ for which the computed reach set will have guarantees over:
$\tilde{X}=\mathrm{\textnormal{proj}}_{1:{n}_{w}}\!\left(\tilde{Z}\right)\times\tilde{K}\times\mathrm{\textnormal{proj}}_{({n}_{w}+1):{n}_{p}}\!\left(\tilde{Z}\right).$ (38)
The projection and Cartesian product are necessary to reorder $\tilde{Z}$ and $\tilde{K}$ into that of (18).
To incorporate tracking error, we first tighten our assumption on the expert plan to find a mode sequence:
Assumption 17 (Expert Plan with Tracking Error).
We assume that a goal-reaching initial condition $\mathbf{x}_{0}$ that defines an expert plan is known and given for the PWA planning model in Section IV-A. Particularly, $\mathbf{x}_{0}$ is any point that satisfies:
$\mathbf{\psi}\!\left(\mathbf{x}_{0}\right)\in P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}}.$ (39)
Intuitively, $E_{t_{\mathrm{\textnormal{f}}},i}$ captures the maximum tracking error at time $t_{\mathrm{\textnormal{f}}}$: if the plans can arrive at the smaller goal region $P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}}$, then the robot should be guaranteed to arrive at the original, bigger goal region $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$.
With a given expert plan, we can obtain $\mathbf{x}_{0}$’s mode sequence using (22), then compute the underapproximated reach set $\Omega_{0}$ of the goal similar to Proposition 10.
Proposition 18 (Reach Set with Tracking Error).
Consider a mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$.. Then we can underapproximate the reach set $\Omega_{0}$ corresponding to this mode sequence by:
$\Omega_{t_{\mathrm{\textnormal{f}}}}=$ $P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}},$ (40a)
$\Omega_{t}=$ $\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$
$\forall t=\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t,$ (40b)
$\Omega_{0}=$ $\mathcal{B}\!\left(\overline{\Omega}_{\Delta t},\mathbf{C}_{s_{\Delta t},\Delta t},\mathbf{d}_{s_{\Delta t},\Delta t}\right)\cap\tilde{X},$ (40c)
Proof.
This follows from Lemma 1, Assumption 16, and Proposition 10. ∎
This extends Proposition 10 by shrinking the goal set and intersection with the tracking error data coverage domain to include tracking error. With $\Omega_{0}$ computed, if the robot tracks any plan generated from an initial condition within $\Omega_{0}$, it is guaranteed to reach the goal at time $t_{\mathrm{\textnormal{f}}}$, but may not be safe; next we guarantee safety by computing the avoid set.
IV-E2 PARC: Avoid Set with Tracking Error
Consider the obstacle $\mathcal{O}_{i}$ and some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$. We define the $i^{\mathrm{\textnormal{th}}}$ buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ between time $t$ and $t+\Delta t$ as:
$\tilde{\mathcal{O}}_{t,i}$ $=\mathcal{O}_{i}\oplus\tilde{E}_{t}.$ (41)
Assuming $\tilde{E}_{t}$ captures the maximum tracking error from $t$ to $t+\Delta t$, by buffering $\mathcal{O}_{i}$ with $\tilde{E}_{t}$, if the trajectory does not collide with $\tilde{\mathcal{O}}_{t,i}$, then the robot following the trajectory should never collide with $\mathcal{O}_{t,i}$. Thus, incorporating tracking errors into the PARC formulation for avoid set follows closely to that of Section IV-C, but with $\tilde{\mathcal{O}}_{t,i}$ instead of $\mathcal{O}_{t,i}$ as an obstacle.
Proposition 19 (Unsafe Plans w/ Tracking Error).
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ and the $(t_{\mathrm{\textnormal{f}}}-t-\Delta t)$-time BRS ${\Omega}_{t+\Delta t}$ using (40). If any planning trajectories between ${\Omega}_{t}$ and ${\Omega}_{t+\Delta t}$ collide with $\tilde{\mathcal{O}}_{t,i}$, that is, if $\exists\ \mathbf{x}(t)\in{\Omega}_{t},\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\in{\Omega}_{t+\Delta t}$ such that
[…]
then
$\tilde{\mathcal{O}}_{t,i}\cap\mathrm{\textnormal{conv}}\!\left({\Omega}_{t},{\Omega}_{t+\Delta t}\right)\neq\emptyset.$ (43)
Proof.
This follows directly from Proposition 18 and Proposition 11. ∎
Finally, we can state our second key contribution: overapproximating the avoid set with tracking error (i.e., computing the BRAS).
Theorem 20 (Avoid set w/ Tracking Error).
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and from (40), the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ of the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\tilde{\mathcal{O}}_{t,i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$. Define the intermediate avoid set ${\Lambda}_{i,t,t}\subset{\Omega}_{t}$ as:
${\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{nonETI}}},\tilde{\mathcal{O}}_{t,i}\right)\cap{\Omega}_{t}.$ (44)
Then, for any $\mathbf{x}(t)\in{\Omega}_{t}\setminus{\Lambda}_{i,t,t}$, $\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}$,
[…]
Proof.
Follows from Proposition 19 and Theorem 13. ∎
We can adapt Corollary 15 to compute the avoid set ${\Lambda}_{i,t,0}$ by computing the $t$-time BRS of ${\Lambda}_{i,t,t}$ as
${\Lambda}_{i,t,j}$ $=\mathcal{B}\!\left({\Lambda}_{i,t,j+1},\mathbf{C}_{s_{j},j},\mathbf{d}_{s_{j},j}\right)\ \forall j=0,\Delta t,\cdots,t-\Delta t.$ (46)
Then, the final overapproximated avoid set ${\Lambda}$ can be stored as a union of H-polytopes as:
[…]
With the avoid set and the reach set from Section IV-E1 computed, if the robot tracks any plans starting in $\Omega_{0}\setminus\Lambda$ it is guaranteed to reach the goal region while avoiding all obstacles. Thus, the set $\Omega_{0}\setminus\Lambda$ is the desired object $\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})$. Next, we demonstrate PARC’s BRAS computation on a variety of numerical examples.
V Experiments
We now assess the utility of our proposed PARC method for generating safe trajectory plans near danger. We seek to understand the impact of planning model design (Section V-A), cooperative vs. adversarial planner-trackers (Section V-B), tight approximation of planning model reachable sets (Section V-C), and respect of time-varying tracking error in planning level (Section V-D). We also assess the utility of learning-based reach-avoid in Sections V-A and V-B.
To ensure fair assessment, we apply PARC to robots from the safe motion planning and control literature: a near-hover quadrotor in 2-D [dawson2022safe] and 3-D [chen2021fastrack], and a more general quadrotor model in 3-D [kousik2019safe]. Implementation details are reported in Appendix -E. All experiments, and training of learning-based methods, are run on a desktop computer with a 24 core i9 CPU, 32 GB RAM, and an Nvidia RTX 4090 GPU.
Key result: PARC outperforms other methods in goal reaching and safety in near-danger scenarios, but requires careful choice of planning model. While it can still guarantee reach-avoid in non-near-danger scenarios, its performance is worse than existing methods. Most experiments are summarized in Table II.
V-A Impact of Planning Model Design
This experiment seeks to understand the impact of choosing a planning model for a given tracking model. We deploy PARC on a tracking model and environment that have been successfully solved by a Neural CLBF [dawson2022safe] approach.
Key result: Naïve design of PARC’s planning model can impact performance negatively. We confirm that a better choice of planning model solves the same task with higher performance in Appendix -G.
V-A1 Experiment Setup
To examine the importance of the planning model, we employ the same dynamics for both the planning and tracking model, namely 2-D near-hover quadrotor dynamics:
$\dot{\mathbf{p}}=\dot{\mathbf{z}}=\frac{d}{dt}\begin{bmatrix}p_{x}\\ p_{z}\\ \theta\\ v_{x}\\ v_{z}\\ \omega\end{bmatrix}=\left[\begin{array}[]{l}v_{x}\\ v_{z}\\ \omega\\ \tfrac{1}{m}\sin(\theta)({F_{\mathrm{\textnormal{l}}}}+{F_{\mathrm{\textnormal{r}}}})\\ \tfrac{1}{m}\cos(\theta)({F_{\mathrm{\textnormal{l}}}}+{F_{\mathrm{\textnormal{r}}}})-g\\ \tfrac{r}{I}({F_{\mathrm{\textnormal{l}}}}-{F_{\mathrm{\textnormal{r}}}})\end{array}\right]$ (54)
The workspace states are $\mathbf{w}=\begin{bmatrix}p_{x}&p_{y}\end{bmatrix}^{\intercal}$, and the inputs are left- and right-side propeller forces
$\mathbf{u}=\mathbf{k}$ $=\begin{bmatrix}{F_{\mathrm{\textnormal{l}}}}&{F_{\mathrm{\textnormal{r}}}}\end{bmatrix}^{\intercal},$
which are held constant for $[0,t_{\mathrm{\textnormal{f}}}]$ to generate plans. The feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ produces time-varying left- and right-side propeller forces using an iLQR controller [kong2021ilqr].
[…]
We use the obstacle and goal environment from [dawson2022safe, Sec. 6.3], wherein the quadrotor must navigate over an obstacle and through a small gap to a goal region (see the left side of Fig. 5 and Fig. 6).
We use PARC to compute a BRAS for this scenario with an expert plan, accounting and without accounting for tracking error. The planning model piece-wise affinize (54) with 80 PWA regions, timestep $\Delta t=0.1$units, and final time $t_{\mathrm{\textnormal{f}}}=2$units. We assess the computation time, as well as the size of the BRAS to investigate the impact of planning model design.
V-A2 Hypothesis
We expect PARC to be able to compute plans to the goal reliably, with no safety violations during $[0,t_{\mathrm{\textnormal{f}}}]$. However, due to the naïve choice of planning model, we expect the avoid set to be overly conservative, especially since (54) has three non-ETI states ($v_{x},v_{z},\omega$) in its PWA formulation.
V-A3 Results
[…]
Figure 5: Results of PARC on a naïve choice of planning model, without accounting for tracking error. Green indicates the goal region, red indicates the obstacles with the volume of the drone accounted for, and blue shows the reach set set-differenced with the avoid set without tracking error. Safe trajectories were successfully generated from the left of the leftmost obstacle.
Figure 6: Results of PARC on a naïve choice of planning model after accounting for tracking error. Green indicates the goal region, red indicates the obstacles with the volume of the drone accounted for, and blue shows the computed BRAS. There were no longer any safe trajectories generating from the left of the leftmost obstacle, which the Neural CLBF was able to accomplish.
V-A4 Discussion
The decrease in BRAS volume was expected due to the many non-ETI states, leading to overconservativeness in avoid set computation. Moreover, as the problem framework requires the trajectory parameters ${F_{\mathrm{\textnormal{l}}}},{F_{\mathrm{\textnormal{r}}}}$ to be constant, the quadrotor shall continue accelerating after reaching the goal region, leading to instability should ${F_{\mathrm{\textnormal{l}}}},{F_{\mathrm{\textnormal{r}}}}$ be tracked.
In Appendix -G and the next few examples, we show how a careful design of the planning model leads to much better performance on more complicated scenarios, as well as enabling stabilizing behavior such as braking without explicitly modeling velocity as a goal state.
V-B Impact of Planner-Tracker Cooperation
We now study cooperative vs. adversarial planning and tracking models. We also further study the Neural CLBF [dawson2022safe] learning-based reach-avoid approach. For this experiment, we use a 3D near-hover quadrotor introduced for FaSTrack [chen2021fastrack], a reachability-based planner tracker that uses adversarial planning and tracking models.
Key result: PARC takes advantage of cooperative planner-tracker behavior to reduce conservativeness.
V-B1 Experiment Setup
The tracking model and PARC’s planning model are listed in Appendix -E. The quadrotor must navigate a narrow gap (width 1.8 m, compared to drone width 0.54 m) to reach its goal as shown in Fig. 7. We simulate 8,100 trials from different initial conditions. We measure the number of trials that reach the goal, and the number that crashes. We also consider the total computation time of each method. Further experiment details are in Appendix -F.
[…]
V-B2 Hypotheses
We anticipate that our PARC planner always reach the goal safely. We also anticipate faster computation time of our reachability analysis versus the other two approaches, but note that FaSTrack does not need to recompute the reachable set for every obstacle. Since FaSTrack provides formal proof of safety, we anticipate it has no crashes. However, since it treats the planner and tracker as adversarial, we expect it to struggle to reach the goal. On the other hand, since CLBFs specifically balance safety and liveness, we expect it to reach the goal more often than FaSTrack but at the expense of crashing due to the high chance of learning an incorrect representation of safety when forced to operate near obstacles.
V-B3 Results
The overall results are reported in Table II. Fig. 7 illustrates results for all methods with an initial position of $[4,-1,3]^{\intercal}$. FaSTrack (orange) fails to find a feasible plan that reaches the goal, the trajectory rolled out by Neural CLBF (pink) violates safety, while the trajectory rolled out by PARC (blue) finds a plan that reaches the goal safely despite tracking error.
Fig. 8 illustrates the trajectories simulated on the grid of 8,100 initial positions using each method. Neural CLBF found 576 successful trajectories (7.1 %), while safety was violated for the rest. FaSTrack found 0 successful trajectories but never violated safety. However, PARC found the plan to reach the goal without collision for 2,592 initial states (32 %).
[…]
Figure 8: Simulated trajectories of 10-D quadrotor for 8,100 initial states using three methods: PARC, FaSTrack [chen2021fastrack], Neural CLBF [dawson2022safe]. The green box is the goal set and the pink boxes are obstacles.
Figure 9: Comparison of tracking error function between FaSTrack and PARC for planning horizon $t_{\mathrm{\textnormal{f}}}=10$units. FaSTrack computes a fixed, worst-case tracking error bound for all time because it assumes an adversarial planner, whereas PARC takes advantage of the smaller, time-varying tracking error that results from a planner and controller working together. The pale pink trajectories are samples used to compute the error as in Section IV-D.
V-B4 Discussion
The performance of FaSTrack is notably compromised by the large, conservatively computed TEB, specified as a $2.47\times 2.47\times 0.1$ box. This often results in the inability to compute feasible plans in spaces tighter than the TEB. However, as seen in Fig. 9, sampled tracking error under parameterized plan, the tracking error is much smaller than TEB, enabling PARC to find a plan in challenging scenarios.
Although Neural CLBF succeeds in capturing most of the safe, unsafe, goal regions, the synthesized control policy steers some of the initial states to violate safety. Thus, the learning-based method sacrifices guarantees.
PARC has a strong safety guarantee, so every plan PARC found did not collide with the obstacles. We note that a better choice of planning model could increase PARC’s success rate; the simplicity of the single-integrator planning model, employed mainly for fair comparison with FaSTrack, reduced the flexibility of possible plans. Indeed we leverage a better choice of trajectory model next.
TABLE II: Performance Evaluation of PARC and Other Reach-Avoid Frameworks: This table compares success and safety rates across various systems and scenarios. The success rate is calculated as the percentage of simulations achieving their goals without collisions, while the safety rate measures the avoidance of obstacles. Computation times are categorized into offline and online phases, with the offline phase not involving known goals, obstacles, or initial states. We report the mean and the standard deviation for time.
[…]
10D Quadrotor PARC 0.32 1.00 $15.34\pm 1.44$ 115 V-B
FaSTrack 0.00 1.00 Timeout 3,600
Neural CLBF 0.07 0.07 18,000 0
PARC-TEF 0.07 1.00 $65.42\pm 9.73$ 2,880 V-C
13D Quadrotor RTD 0.00 1.00 Timeout 2,880
(Narrow Gap) PARC-Funnel 0.08 1.00 $12.20\pm 1.70$ 0 V-D
KDF 0.00 1.00 Timeout 0
13D Quadrotor PARC-Funnel 1.00 1.00 $3.07\pm 0.06$ 00 V-D
(Wide Gap) KDF 1.00 1.00 $0.44\pm 0.27$ 0
V-C Impact of Tight Approximation for Planning Reachability
We seek to understand the importance of computing a tighter approximation of the reachable set for the planning model. To do this, we compare against RTD implemented on a general 3-D quadrotor [kousik2019safe] using zonotope reachability [althoff2015introduction]. This allows us to use the same tracking error model but employs a different approach to planning model reachability analysis.
V-C1 Experiment Setup
[…]
We use the same planning model as quadrotor RTD [kousik2019safe], which generates time-varying polynomial positions in 3-D [mueller2015computationally] with $\Delta t=0.02$units with $t_{\mathrm{\textnormal{f}}}=3$units; this formulation is affine time-varying in the trajectory parameters with no non-ETI states, so reachability can be tightly approximated by PARC. Both PARC and RTD use the same tracking error that is precomputed in [kousik2019safe].
[…]
We declare safe if the quadrotor does not crash, and success if the quadrotor reaches the goal region without colliding with any obstacles.
V-C2 Hypothesis
We anticipate that PARC will always safely find a way to reach the goal since we know the narrow gap is large enough to do so, and PARC’s primary source of conservativeness is its tracking error estimation (which is identical to that of RTD). We anticipate that RTD will reach the goal much less often because it suffers an additional course of conservativeness from using a polynomial planning model with zonotopic reachable sets.
V-C3 Results
[…]
V-C4 Discussion
As expected, both RTD and PARC guaranteed safety. However, only PARC was able to traverse through the narrow gap. Thus, this example clearly illustrates the utility of PARC’s tightly approximating reachable set computation in comparison to RTD’s approximation. Although PARC is significantly slower in planning when compared to RTD, once a BRAS is computed, the robot obtains a continuum of safe plans.
[…]
V-D Impact of Time-Varying Tracking Error
Figure 11: Comparison of PARC (a-b) and KDF (c-d) motion planning for a 3-D quadrotor across two scenarios: a narrow gap of $0.85m$ (left) and a wide gap of $3m$ (right), projected onto the $p_{x}$-$p_{y}$ plane. The initial planning state is $\mathbf{p}_{0}=[4.8,1,4]^{{}^{\top}}$ with a planning horizon of $t_{\mathrm{\textnormal{f}}}=15$units. Goals are in the green box, the dense pink boxes indicate the actual obstacles and the lighter pink boxes show the obstacles buffered by the drone’s volume. Pink trajectories within blue volumes represent computed motion plans and bounds of realized trajectories when tracking the plan (i.e. funnels), illustrating how each method handles tracking errors at the motion-planning level.
Finally, we explore the importance of accounting for the temporal correlation of tracking errors. In this experiment, we compare our approach with the KDF framework that combines sampling-based planning with funnel-based feedback control [verginis2022kdf]. This control mechanism prescribes that tracking error remain in funnels described by a performance function, $\rho(t)$, which decays exponentially over time [bechlioulis2014low]. Employing the same funnel-based control across both the KDF and PARC methodologies allows us to maintain a consistent tracking error model. This setup confirms that any observed differences in performance are solely attributed to the planner’s strategy of using tracking errors.
Key result: By considering time-varying as opposed to worst-case tracking error, PARC significantly reduces conservativeness of motion planning. Furthermore, PARC’s use of parameterized trajectories enabled the design of a funnel controller that maintained the prescribed tracking error bounds, whereas the same controller could not reliably track the smoothed output of a sampling-based motion planner.
Remark 21.
As shown in the other experiments, PARC can leverage a data-driven tracking error function that is potentially less conservative than the prescribed $\rho(t)$. In this experiment, we use $\rho(t)$ to ensure a fair comparison between PARC and KDF.
V-D1 Experiment Setup
We extend the setup from Section V-B to compare PARC and KDF using gap widths of 0.85 m, and 3 m, relative to the quadrotor’s 0.54 m of body width. We evaluate success and safety ratios, along with computation times, by simulating the quadrotor from zero velocity at 3,700 and 2,400 uniformly distributed initial positions across each scenario.
For the planner, PARC employs the polynomial planning model from Section V-C, but with $t_{\mathrm{\textnormal{f}}}=15\sec$. $\Delta t$ is $0.2\sec$ for the narrow gap and $0.5\sec$ for the wide gap. KDF employs RRT [lavalle1998rapidly] as its motion planner.
For the tracker, both PARC and KDF use a specialized funnel controller adapted for an under-actuated 3-D quadrotor proposed in [lapandic2022robust], as the conventional KDF funnel-based feedback controller is unsuitable for such systems [lapandic2023kinodynamic]. This controller conforms to the performance function $\rho(t)=\rho_{0}+(\rho_{0}-\rho_{\infty})e^{-\lambda t}$ where we choose $\rho_{0}=0.5,\rho_{\infty}=0.1,\lambda=0.5$. We note that this controller does not respect input limits. The 48 hyperparameters for the funnel controllers are reported in Appendix -F.
Despite extensive tuning efforts, a universal hyperparameter set for KDF could not be determined that respected the prescribed tracking error bounds; thus, we assume the ideal performance of the funnel controller, and base our statistics on reference trajectories ignoring tracking error. In other words, the comparison is biased towards benefitting KDF and disadvantaging PARC. We note that we designed controllers that function for a selected few initial states with KDF, the outcomes of which are demonstrated in Appendix -F and Fig. 11. In contrast, since PARC uses parameterized reference trajectories, we were able to succesfully identify a universal set of hyperparameters that enabled us to base the statistics on realized trajectories that included tracking error.
V-D2 Hypothesis
In the wide-gap scenario, both PARC and KDF are hypothesized to successfully execute reach-avoid plans. We anticipate KDF to demonstrate faster computation times, benefiting from its sampling-based planning approach. In the narrow-gap scenario, KDF is expected to struggle due to its reliance on maximum funnel values for free space construction in RRT, whereas PARC is predicted to succeed by leveraging the time-varying nature of tracking errors. We expect no safety violations for all experiments.
V-D3 Results
[…]
V-D4 Discussion
As anticipated, PARC demonstrates superior performance over KDF in narrow-gap scenarios due to its ability to utilize the time-varying nature of tracking errors effectively; thus taking advantage of a low tracking-error regime. This advantage is visually depicted in Fig. 11 a, where PARC waits for the funnel controller to reduce the tracking error before navigating through narrow gaps. In contrast, for wider gaps shown in Fig. 11 b, the quadrotor proceeds immediately. Conversely, KDF struggles in narrow spaces, as depicted in Fig. 11 c, because its planner computes paths based on a static maximum tracking error that exceeds the actual gap width (i.e., $\rho_{0}=0.5>0.31$), but performs well in wider spaces where this limitation does not impede planning as depicted in Fig. 11 d.
Notably, the difficulty in tuning the funnel controller to follow the smoothed reference trajectory proposed by RRT has led us to assume perfect funnel control in reporting KDF’s statistics. This aligns with observations in Section V-B that parameterized trajectories foster better cooperation with controllers compared to trajectories derived from smoothing splines over RRT waypoints, which can create adversarial tracking conditions.
PARC achieves reach-avoid plans in narrow-gap scenarios but faces lengthy computation times due to detailed sampling from BRAS and smaller $\Delta t$ steps. In contrast, KDF uses sampling-based motion planning (SBMP) effectively in scenarios with adequate gap widths, swiftly creating viable plans. This indicates a potential research direction: integrating SBMP within PARC could shorten computation times and facilitate the generation of expert trajectories that adeptly handle near-danger scenarios, surpassing conventional kinodynamic SBMP methods such as KDF.
V-E Overall Discussion
Through the above experiments, we have shown how our proposed PARC approach confers advantages in computing and using the Backward Reach-Avoid Set (BRAS). In near-danger scenarios, PARC was less conservative than FaSTrack and KDF since the former uses a time-variant tracking error function, whereas the latter two assumes constant, worst-case tracking errors. While RTD can similarly handle time-variant tracking error functions, PARC still outperforms it due to the lower numerical approximation error from representing the planning model as a PWA system instead of a polynomial and using H-polytopes instead of zonotopes. Finally, Neural CLBF cannot provide hard reach-avoid guarantees, resulting in many crashes with obstacles.
However, PARC still has room for improvement. Most obviously, the safe plan computation time is on the order of seconds for just a few obstacles; this is too slow for real-time operation. In contrast, RTD [kousik2019safe], although more conservative and failed to reach the goal, can compute a new plan about every 51 ms in Section V-C. Typically, we expect motion planners with real-time capabilities to compute new plans for at least every 0.5 s [kousik2020bridging]. To achieve this for PARC, parallelization needs to be implemented as a future work. Furthermore, the planning model requires careful hand-crafting as shown in our experiments.
VI Drift Vehicle Demonstration
We now demonstrate the utility of PARC in safely planning extreme near-danger vehicle drifting maneuvers, where the robot loses controllability.
Key result: PARC enables computation of provably safe drift parking trajectories for the first time, extending the state of the art.
VI-A Background
VI-A1 Drifting
Drifting is when a vehicle turns with a high sideslip angle, causing tire saturation and losing direct control of its heading [goh2016simultaneous]. Stable drifting can be achieved by finding “drift equilibria” with constant yaw rate and sideslip angle of the highly nonlinear dynamics [goh2016simultaneous, goh2020toward, goel2020opening, weber2023modeling]. Drift parking, however, remains challenging due to the highly coupled inputs [leng2023deep] and reduced controllability as the vehicle specifically must not maintain a drift equilibrium [kolter2010probabilistic].
VI-A2 Reach-Avoid Method Comparison
To the best of our knowledge, no other method can guarantee safety and liveness for drift parking.
Our drifting tracking model dynamics (see (96) below) are not control affine, so directly applying a CBF or CLBF control approach is not tractable [ames2016control, dawson2023safe]; future work may apply results for non-control-affine systems [son2019safety, seo2022safety]. Note, [dawson2022safe] poses a control affine vehicle dynamics model but does not include tire saturation, so cannot drift. Furthermore, our planning and tracking models are too high-dimensional for standard HJB value function computation [chen2021fastrack, mitchell2008flexible]; future work will compare against learning-based HJB [bansal2021deepreach]. The most similar work is RTD [kousik2020bridging, liu2022refine], which has not been applied to these extreme dynamics. As a future implementation, RTD’s reachable set can use PARC’s formulation to enable a direct drifting comparison; then, the only difference with PARC is that RTD implicitly represents the avoid set. Due to this similarity, we focus on demonstrating PARC’s capabilities as opposed to a head-to-head comparison.
VI-B Demo Details
Extensive details are presented in Appendix -H. Here, we briefly present setup, simulation results, and a discussion.
VI-B1 Planning and Tracking Models
As we saw in Section V-A, carefully planning model design is critical to PARC’s reach-avoid. In this case, we fit an affine time-varying model to drifting data collected by rolling out open-loop drifting maneuvers. This not only guarantees a reasonable planning model (i.e., it successfully parameterizes drifting maneuvers) but also demonstrates the flexibility of using PWA models for PARC since they can readily be learned from expert demonstrations.
We use a tracking model based on [jelavic2017autonomous, goh2020toward, weber2023modeling, goh2016simultaneous], detailed in Appendix -H. Our tracking controller is based on [jelavic2017autonomous]: we first use nonlinear MPC to bring the car into a drifting state, then switch to open-loop control to complete the parking action (since the vehicle loses controllability when sliding sideways to a stop).
VI-B2 Environment and Demo Setup
The objective of this demo is to drift safely into a parallel parked state between two closely parked cars, as shown in Fig. 1 and Fig. 12. Our environment is inspired by a real-world demo video [guinness2014parallel].
We assess PARC’s ability to compute the BRAS for this maneuver. We further evaluate safety trajectory rollouts using initial conditions and trajectory parameters sampled from the BRAS to check for collision.
VI-B3 Results
Figure 12: Example planned drifting trajectories sampled from the Backward Reach-Avoid Set using PARC. The two red cars are obstacles. The goal set for the center of mass is in green. The trajectory with a very low tracking error is shown in (a), whereas a high tracking error is shown in (b). Critically neither collide with the obstacles, and both are represented in the single BRAS.
[…]
VI-B4 Discussion
This demo showcased (i) PARC’s capability to generate safe and goal-reaching trajectories with challenging near-danger tasks such as reach-avoid drift parking, and (ii) the flexibility of PARC in terms of handling a suitable planning model from data. We notice that the BRAS is relatively small (see Fig. 1(b)) in the state space, indicating that drift parking is indeed a difficult maneuver. This emphasizes the importance of PARC minimizing conservativeness in the BRAS computation. However, it also suggests that this maneuver is brittle to large changes in initial conditions, and PARC will struggle to find drifting maneuvers for different goal sets or obstacle configurations. In future work, we plan to test more robust drifting controllers to produce a larger BRAS. With that in mind, our ultimate goal is physical hardware experiments on a platform such as the F1:10 car in Fig. 17.
VII Hardware Demonstration
Thus far, PARC has only been shown to work in simulation. In reality, the tracking model will have model mismatch with the actual hardware due to disturbances and perception errors. As such, we now present preliminary results for PARC’s reach-avoid guarantees applied on mobile robot hardware navigating a narrow gap, as shown in Fig. 13. Specifically, we use a TurtleBot3 Burger Model [grossimplementation] differential drive robot running ROS 2 [macenski2022robot].
VII-A Environment and Demo Setup
[…]
Figure 13: Problem setup of the hardware demo. A snapshot of the robot’s realized motion and the tracked trajectory (gradient line) from one run of the experiment was shown. The robot successfully avoided the obstacles (red) and reached the goal set (green).
The robot is equipped with IMU sensors and wheel encoders, which gives estimation of its planning states $p_{x}$, $p_{y}$, and $\theta$. As future work, these states can be more accurately estimates using an onboard lidar sensor.
VII-B Planning and Tracking Models
[…]
VII-C Results
[…]
VII-D Discussion
[…]
VIII Conclusion
This paper has presented an approach to use Piecewise Affine Reach-avoid Computation (PARC) for safe goal-reaching robot trajectory planning near danger. PARC enables goal-reaching through narrow gaps and in tight obstacle configurations, even with extreme dynamics. The method outperforms a variety of state-of-the-art reach-avoid methods and establishes a new benchmark for safe, extreme vehicle motion planning. Future work will improve PARC’s computation time, consider more types of uncertainty as opposed to lumped tracking error, and extend beyond mobile robots.
Limitations: PARC has several key limitations. First and foremost, it requires an assumption that the data coverage of sampled tracking error is large enough. While this assumption holds in practice (over hundreds of collision-free trials planned in this paper), it means that further mathematical work is necessary for true safety guarantees. Second, if the planning of the robot takes place at the joint level, which is common for manipulator and legged robots, augmentation of the obstacles with tracking error is not straight-forward. Though this has been successfully done in other set-based planner-tracker reach-avoid methods such as ARMTD [holmes2020reachable] and ARMOUR [michaux2023can] by expressing H-polytopes in configuration space as polynomial zonotopes [kochdumper2020reachability] in workspace, we opted to leave its extension to PARC as future work, as the focus of this paper is just computing the BRAS in the first place. Third, although PARC is faster than compared methods, it is still too slow for real-time; the present implementation has not been parallelized and may benefit from alternative representations to H-polytopes. Fourth, PARC has only been compared on hand-crafted examples of goal reaching near danger, and hardware validation has only been done on a simple robot. The experiments have only considered tracking error, with the hardware validation considering error from localization drift of the IMU and encoders in addition. It is critical to validate the method across randomly-generated, realistic environments on a variety of hardware with multiple significant sources of uncertainties, such as perception error. To this end, the authors are preparing a small robotic car to perform drift parking experiments in the coming months.
[…]
-A Related Work
We now discuss methods that guarantee safety (avoiding dangerous states) and liveness (reaching a goal) for control systems and robot motion planning. In particular, we discuss optimal control methods, safety filters, and planner-tracker frameworks. We also note connections between our approach and sampling-based motion planning. Finally, we discuss piecewise affine (PWA) system models, which we leverage to overcome limitations of conservativeness and computational tractability for providing safety and liveness guarantees with state-of-the-art methods.
-A1 Optimal Control Methods
Many safety methods focus on synthesizing or analyzing controllers. There has been extensive development of techniques such as Control Barrier Functions (CBFs) [ames2016control, agrawal2017discrete, rosolia2022unified, castaneda2022probabilistic], Hamilton-Jacobi (HJ) reachability [herbert2021scalable, seo2022real], sums-of-squares reachability [majumdar2017funnel, tedrake2010lqr], robust model predictive control (MPC) [rakovic2012parameterized, villanueva2017robust, wabersich2021predictive], and graphs of convex sets (GCS) [marcucci2023motion, cohn2023non]. Yet, only a minority of these methods address the dual objectives of liveness and safety [hsu2021safety, so2023solving, borquez2023safety, fridovich2019safely]. Some methods seek a balance between performance and safety by implicitly integrating each constraint into the cost function of optimal control problem, solved through methods like MPC [williams2016aggressive] or reinforcement learning [srinivasan2020learning], but can fail to offer hard satisfaction of such constraints.
The key challenge with most optimal control methods is to synthesize a controller that accounts for arbitrary bounded disturbances to a model of a robot. By contrast, in this work, we seek to synthesize a planner that cooperates with, and leverages knowledge about, an imperfect controller.
-A2 Safety Filters
A related strategy prioritizes safety through the use of safety filters [wabersich2021probabilistic, wabersich2023data]; indeed, many optimal control methods can also be cast as safety filters [hsu2023safety, ames2016control, agrawal2017discrete]. In this case, a performance-oriented controller is minimally altered by a safety-oriented controller when the system is about to violate the safety constraints, mostly building on the safe invariant sets. Since this approach prioritizes safety over performance, thereby enforcing the hard satisfaction of safety in any case, the safety filter might produce conservative behavior. One way to mitigate conservativeness is to leverage reinforcement learning to exploit the safety filter [thumm2022provably, shao2021reachability].
The key challenge with safety filters is that they intervene at the control level, and typically treat upstream planning as a disturbance [chen2021fastrack, singh2020robust]. This can severely impact liveness, because the planner and controller are treated as adversarial, as opposed to cooperative. By contrast, we adopt a cooperative framework.
Another notable challenge, which we do not address, is extending to systems with very high-dimensional states or input/observation spaces. While some traditional approaches have tried a divide-and-conquer approach for high dimensions [chen2016fast], the most promising path forward appears to be learning-based methods [bansal2021deepreach, michaux2023reachability, dawson2022safe, xiao2023barriernet, selim2022safe]. However, by relying on deep or machine learning, these methods essentially lose all safety guarantees above 12-D. We look forward to potentially attacking high-dimensional systems by fusing our proposed method with learning in the future.
-A3 Planner-Tracker Methods
The final category of safe planing and control methods that we discuss is the planner-tracker framework [chen2021fastrack, kousik2019safe, kousik2020bridging, michaux2023reachability]. These methods use a simplified model to enable fast plan generation; plans are then tracked by a controller that uses a high-fidelity tracking model of a robot.
Two representative approaches are FaSTrack [chen2021fastrack] and Reachability-Based Trajectory Design (RTD) [kousik2019safe, kousik2020bridging]. FaSTrack employs Hamilton-Jacobi (HJ) reachability to precompute a worst-case error bound between the planning and tracking models, then dilates obstacles by this bound to ensure safe planning online. RTD precomputes a forward reachable set (FRS) of a parameterized set of planning model trajectories and associated (time-varying) tracking error, then uses the FRS to find safe plans at runtime. We compare our proposed method against both FaSTrack and RTD.
The primary challenge for planner-tracker methods is that the computational advantage of the planning model comes at the cost of model discrepancy [chen2021fastrack, singh2020robust, kousik2017safe, kousik2019safe]. An additional key challenge is that the numerical representation of safety can be overly conservative even when just representing the planning model and ignoring model error. In this work, we design a planner-tracker framework that exactly represents trajectories of the planning model using piecewise affine systems, meaning that we only need to account for model error. We find that this significantly reduces conservativeness over related methods.
-A4 Sampling-Based Motion Planning
The challenge of planning near danger, or through a narrow gap, has been well-studied with classical sampling-based motion planning techniques [elbanhawi2014sampling] and is still an active area of research (e.g., [orthey2024multilevel, liu2023simultaneous, yu2023gaussian]). When paths or trajectories must pass near obstacles, it becomes critical to accurately represent dynamics and account for error in modeling and control. To this end, methods such as kinodynamic RRT* [webb2013kinodynamic] require repeatedly rolling out and collision-checking the trajectories, as well as solving two-point boundary value problems of nonlinear dynamical systems (i.e., steer function) online, which can be computationally expensive. If in addition the dynamical system is high-dimensional and/or the set of true feasible reach-avoid set is small (which is the case for “near danger” problem setups), it could take a very long time for a solution to be found [berenson2009manipulation].
In this paper, instead of sampling to find paths or trajectories, we directly approximate a continuum of all feasible trajectories through a narrow gap. That said, we do not view our approach as competing with sampling-based motion planning. By contrast, our PARC planner is complementary to sampling in three ways. First, PARC requires expert plans that seed our reachability analysis; in this paper, we generate the plans by approximating the solution to a two-point boundary value problem, but we could leverage a sampling approach instead. Second, as we show in our experiments, PARC is only necessary when a goal is specifically near obstacles; a sampling-based approach that ignores dynamics could be used to find goals and identify these scenarios (e.g., if the sampling-based approach fails, then one should switch to using PARC). Third, our approach can be used to robustify a sampling-based approach by computing safety funnels around a sampled plan, similar to LQR trees [tedrake2010lqr] or funnel libraries [majumdar2017funnel].
-A5 Piecewise-Affine Systems
In this paper, we propose a parameterized, time-variant piecewise-affine (PWA) system as our planning model. PWA systems, which evolve through different affine models based on state regions (modes), adeptly capture nonlinear dynamics. Their capacity for modeling non-smooth dynamics through hybridization is well-documented [asarin2007hybridization, dang2010accurate] and they are prevalently used to approximate complex nonlinear hybrid systems, including legged locomotion [marcucci2017approximate], traffic systems [mehr2017stochastic], and neural network [vincent2021reachable].
However, planning with PWA systems, which involves solving their control problem, is recognized as challenging due to the necessity of determining both input and mode sequences [sadraddini2019sampling]. Standard solutions, including mixed-integer convex programming (MICP) [marcucci2019mixed] and multiparametric programming [sakizlis2007linear, baotic2005optimal], face computational challenges, particularly with increasing time steps exacerbating the number of integer decision variables.
To handle this challenge, we take advantage of mode sequences; it is well-established that a known mode sequence greatly simplifies computation [thomas2006robust, kerrigan2002optimal, rakovic2005robust]. By leveraging sampling to synthesize likely mode sequences, our proposed approach significantly improves computational efficiency to enable successful computation of reach-avoid motion plans.
-B Operations on H-Polytopes
The intersection of H-polytopes is an H-polytope:
$\begin{split}\mathcal{P}(\mathbf{A}_{1},\mathbf{b}_{1})\cap\mathcal{P}(\mathbf{A}_{2},\mathbf{b}_{2})&=\left\{\mathbf{x}\ |\ \mathbf{A}_{1}\mathbf{x}\leq\mathbf{b}_{1},\mathbf{A}_{2}\mathbf{x}\leq\mathbf{b}_{2}\right\},\\ &=\mathcal{P}\left(\begin{bmatrix}\mathbf{A}_{1}\\ \mathbf{A}_{2}\end{bmatrix},\begin{bmatrix}\mathbf{b}_{1}\\ \mathbf{b}_{2}\end{bmatrix}\right).\end{split}$ (55)
The Minkowski sum $\oplus$ is
$P_{1}\oplus P_{2}$ =$=\left\{\mathbf{x}+\mathbf{y}\ |\ \mathbf{x}\in P_{1},\mathbf{y}\in P_{2}\right\},$ (56)
which can be computed as one H-polytope by solving a linear program (LP) [herceg2013multi]. The Minkowski sum essentially adds a “buffer” to the object being summed, which we use to pad the obstacles to account for tracking error in our reachability analysis.
Similarly, the Pontryagin Difference $\ominus$ is
$P_{1}\ominus P_{2}$ =$=\left\{\mathbf{x}\in P_{1}\ |\ \mathbf{x}+\mathbf{y}\in P_{1}\ \forall\ \mathbf{y}\in P_{2}\right\},$ (57)
which can be expressed as one H-polytope by subtracting the support of $P_{2}$ for each inequality of $P_{1}$ [herceg2013multi]. This essentially inverts the Minkowski sum, reducing the volume of an H-polytope, which we use to shrink a goal set to account for tracking error.
The Cartesian product $\times$ is exactly an H-polytope:
$P_{1}\times P_{2}$ =$=\left\{\begin{bmatrix}\mathbf{x}\\ \mathbf{y}\end{bmatrix}\ \Bigr{|}\ \mathbf{x}\in\mathcal{P}(\mathbf{A}_{1},\mathbf{b}_{1}),\mathbf{y}\in\mathcal{P}(\mathbf{A}_{2},\mathbf{b}_{2})\right\},$ (58a)
$=\mathcal{P}\left(\begin{bmatrix}\mathbf{A}_{1}&\mathbf{0}\\ \mathbf{0}&\mathbf{A}_{2}\end{bmatrix},\begin{bmatrix}\mathbf{b}_{1}\\ \mathbf{b}_{2}\end{bmatrix}\right).$ (58b)
We use this to combine low-dimensional reachable sets into higher dimensions, and to ensure that H-polytopes have equal dimensions for operations such as Minkowski sums and Pontryagin differences.
The convex hull is
[…]
which can be computed as one H-polytope by vertex enumeration, computing the convex hull of the vertices, and expressing the result in H-polytope form [fukuda2003cddlib]. The convex hull contains all straight lines between two sets, which we use to overapproximate trajectories between discrete time points to secure continuous time safety guarantees.
For an ${n}$-dimensional H-polytope $\mathcal{P}(\mathbf{A},\mathbf{b})$, the projection from the ${p}^{\mathrm{\textnormal{th}}}$ dimension to the ${q}^{\mathrm{\textnormal{th}}}$ dimension for some ${p}\leq{q}\leq{n}$ is defined as:
$\begin{split}\mathrm{\textnormal{proj}}_{{p}:{q}}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b})\right)&=\left\{\begin{bmatrix}\mathbf{0}_{a},\ \mathbf{I}_{b},\ \mathbf{0}_{c}\end{bmatrix}\mathbf{x}\ |\ \mathbf{A}\mathbf{x}\leq\mathbf{b}\right\},\quad\mathrm{\textnormal{where}}\\ a&=({q}-{p}+1)\times({p}-1),\\ b&=({q}-{p}+1)\ \mathrm{\textnormal{and}}\\ c&=({q}-{p}+1)\times({n}-{q})\end{split}$ (60)
The right-hand side of (60) is exactly an AH-polytope, which can be converted to an H-polytope using block elimination, Fourier-Motzkin elimination, parametric linear programming, or vertex enumeration [fukuda2003cddlib, marechal2017scalable]. Somewhat the opposite of Cartesian product, the projection operation lowers the dimensions of the input polytope, which will be useful for retrieving low-dimensional information if only a subset of the original states is of interest.
Note that convex hull and projection are NP-hard for H-polytopes, exponentially scaling in computational time and complexity with the dimension of the polytope [althoff2021set]. However, since PARC performs polytope computation only for low-dimensional ($\leq$ 6-D) planning models, the computational time and complexity remain reasonable, as shown in Section V.
Finally, slicing [herceg2013multi] an ${n}$-dimensional H-polytope $\mathcal{P}(\mathbf{A},\mathbf{b})$ from the ${p}^{\mathrm{\textnormal{th}}}$ dimension to the ${q}^{\mathrm{\textnormal{th}}}$ dimension with respect to a constant $\mathbf{x}_{0}\in\mathbb{R}^{{q}-{p}+1}$, where ${p}\leq{q}\leq{n}$, is defined as:
$\mathrm{\textnormal{slice}}_{{p}:{q}}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{x}_{0}\right)$ =$=\left\{\begin{bmatrix}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{bmatrix}\ |\ \mathbf{A}\begin{bmatrix}\mathbf{y}_{1}\\ \mathbf{x}_{0}\\ \mathbf{y}_{2}\end{bmatrix}\leq\mathbf{b}\right\},$ (61a)
=$=\left\{\mathbf{y}\ |\ \begin{bmatrix}\mathbf{A}_{1},\ \mathbf{A}_{3}\end{bmatrix}\mathbf{y}\leq\mathbf{b}-\mathbf{A}_{2}\mathbf{x}_{0}\right\},$ (61b)
$=\mathcal{P}(\begin{bmatrix}\mathbf{A}_{1},\ \mathbf{A}_{3}\end{bmatrix},\mathbf{b}-\mathbf{A}_{2}\mathbf{x}_{0}),$ (61c)
where $\mathbf{A}=[\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3}]$, $\mathbf{A}_{1}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}_{{p}-1}}$, $\mathbf{A}_{2}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}_{{q}-{p}+1}}$, $\mathbf{A}_{3}\in\mathbb{R}^{{n}_{\mathrm{\textnormal{h}}}\times{n}_{{n}-{q}}}$. Thus, slicing results in exactly an H-polytope. We use slicing to aid in some visualization and mathematical proofs.
-C Method for Affinizing a Nonlinear System
We convert a nonlinear system into a PWA system via sampling in its state space, then computing Voronoi regions and linearizing the dynamics in each region, similar to in [casselman2009new]. In particular, we apply the following:
Proposition 22.
Express the domain $X$ as a H-polytope $\mathcal{P}(\mathbf{A}_{X},\mathbf{b}_{X})$. Consider linearization points $\mathbf{x}^{*}_{t,i}\in X$, $i=1,\cdots,{n}^{*}_{t}$. Then for $i=1,\cdots,{n}^{*}_{t}$, the Voronoi cells $V_{i,t}$ given by
$V_{i,t}$ $=\bar{V}_{i,t}\cap X,$ (62a)
$\bar{V}_{i,t}$ =$=\left\{\mathbf{x}\ |\ \left\|\mathbf{x}-\mathbf{x}^{*}_{t,i}\right\|^{2}_{2}\leq\left\|\mathbf{x}-\mathbf{x}^{*}_{t,j}\right\|^{2}_{2},j=1,\cdots,{n}^{*}_{t}\right\},$ (62b)
are compact H-polytopes that fulfill (6), and can therefore be the PWA regions of a PWA system.
Proof.
It was proven in [boyd2004convex] that $\bar{V}_{i,t}$ is a closed, but not necessarily bounded H-polytope. In fact,
$\bar{V}_{i,t}=$$\Biggl{\{}\mathbf{x}\ \Biggr{|}\ \begin{bmatrix}(\mathbf{x}-\mathbf{x}^{*}_{t,i})\cdot(\mathbf{x}-\mathbf{x}^{*}_{t,i})\\ \vdots\\ (\mathbf{x}-\mathbf{x}^{*}_{t,i})\cdot(\mathbf{x}-\mathbf{x}^{*}_{t,i})\end{bmatrix}\leq$
$\begin{bmatrix}(\mathbf{x}-\mathbf{x}^{*}_{t,1})\cdot(\mathbf{x}-\mathbf{x}^{*}_{t,1})\\ \vdots\\ (\mathbf{x}-\mathbf{x}^{*}_{t,{n}^{*}(t)})\cdot(\mathbf{x}-\mathbf{x}^{*}_{t,{n}^{*}(t)})\end{bmatrix}\Biggr{\}},$ (63a)
$=$ $\Biggl{\{}\mathbf{x}\ \Biggr{|}\ \begin{bmatrix}(2\mathbf{x}^{*}_{t,1}-2\mathbf{x}^{*}_{t,i})^{\intercal}\\ \vdots\\ (2\mathbf{x}^{*}_{t,{n}^{*}(t)}-2\mathbf{x}^{*}_{t,i})^{\intercal}\end{bmatrix}\mathbf{x}\leq$
$\begin{bmatrix}\left\|\mathbf{x}^{*}_{t,1}\right\|_{2}^{2}-\left\|\mathbf{x}^{*}_{t,i}\right\|_{2}^{2}\\ \vdots\\ \left\|\mathbf{x}^{*}_{t,{n}^{*}(t)}\right\|_{2}^{2}-\left\|\mathbf{x}^{*}_{t,i}\right\|_{2}^{2}\end{bmatrix}\Biggr{\}},$ (63b)
$=$ $\mathcal{P}\left(\begin{bmatrix}(2\mathbf{x}^{*}_{t,1}-2\mathbf{x}^{*}_{t,i})^{\intercal}\\ \vdots\\ (2\mathbf{x}^{*}_{t,{n}^{*}(t)}-2\mathbf{x}^{*}_{t,i})^{\intercal}\end{bmatrix},\begin{bmatrix}\left\|\mathbf{x}^{*}_{t,1}\right\|_{2}^{2}-\left\|\mathbf{x}^{*}_{t,i}\right\|_{2}^{2}\\ \vdots\\ \left\|\mathbf{x}^{*}_{t,{n}^{*}(t)}\right\|_{2}^{2}-\left\|\mathbf{x}^{*}_{t,i}\right\|_{2}^{2}\end{bmatrix}\right).$ (63c)
Since H-polytopes are closed under intersection (as shown in (55)), and the intersection of a closed set and a compact set is compact, $V_{i,t}$ is a compact H-polytope from (62a).
Further, for any $i,j\in\{1,\cdots,{n}^{*}_{t}\}$, $i\neq j$, the intersection of the interior of $V_{i,t}$ with $V_{j,t}$ is given by:
$\mathrm{\textnormal{int}}\!\left(V_{i,t}\right)\cap V_{j,t}=\Bigl{\{}\mathbf{x}\ |\ \left\|\mathbf{x}-\mathbf{x}^{*}_{t,i}\right\|^{2}_{2}<\left\|\mathbf{x}-\mathbf{x}^{*}_{t,k}\right\|^{2}_{2},$
$\left\|\mathbf{x}-\mathbf{x}^{*}_{t,j}\right\|^{2}_{2}\leq\left\|\mathbf{x}-\mathbf{x}^{*}_{t,k}\right\|^{2}_{2},k=1,\cdots,{n}^{*}_{t}\Bigr{\}}\cap\mathrm{\textnormal{int}}\!\left(X\right),$ (64)
where $\mathrm{\textnormal{int}}\!\left(\cdot\right)$ is the interior of a set. However, notice that $\left\|\mathbf{x}-\mathbf{x}^{*}_{t,i}\right\|^{2}_{2}<\left\|\mathbf{x}-\mathbf{x}^{*}_{t,k}\right\|^{2}_{2}$ and $\left\|\mathbf{x}-\mathbf{x}^{*}_{t,j}\right\|^{2}_{2}\leq\left\|\mathbf{x}-\mathbf{x}^{*}_{t,k}\right\|^{2}_{2}$ for $k=1,\cdots,{n}^{*}(t)$ implies $\left\|\mathbf{x}-\mathbf{x}^{*}_{t,i}\right\|^{2}_{2}<\left\|\mathbf{x}-\mathbf{x}^{*}_{t,j}\right\|^{2}_{2}$ and $\left\|\mathbf{x}-\mathbf{x}^{*}_{t,j}\right\|^{2}_{2}\leq\left\|\mathbf{x}-\mathbf{x}^{*}_{t,i}\right\|^{2}_{2}$, which is impossible. Thus,
$\mathrm{\textnormal{int}}\!\left(V_{i,t}\right)\cap V_{j,t}=\emptyset,$ (65)
which is exactly the condition (6).
Therefore, $V_{i,t}$ defines the PWA regions of a PWA system, with each of the ${n}_{\mathrm{\textnormal{PWA}},t}={n}^{*}_{t}$ regions given by:
$\mathbf{A}_{i,t}$ $=\begin{bmatrix}(2\mathbf{x}^{*}_{t,1}-2\mathbf{x}^{*}_{t,i})^{\intercal}\\ \vdots\\ (2\mathbf{x}^{*}_{t,{n}^{*}(t)}-2\mathbf{x}^{*}_{t,i})^{\intercal}\\ \mathbf{A}_{X}\end{bmatrix},$ (66)
$\mathbf{b}_{i,t}$ $=\begin{bmatrix}\left\|\mathbf{x}^{*}_{t,1}\right\|_{2}^{2}-\left\|\mathbf{x}^{*}_{t,i}\right\|_{2}^{2}\\ \vdots\\ \left\|\mathbf{x}^{*}_{t,{n}^{*}(t)}\right\|_{2}^{2}-\left\|\mathbf{x}^{*}_{t,i}\right\|_{2}^{2}\\ \mathbf{b}_{X}\end{bmatrix},$ (67)
for $i=1,\cdots,{n}^{*}_{t}$. ∎
-D Proof of Theorem 13: Avoid Set without Tracking Error
Here, we prove Theorem 13, which is our paper’s main result. We first introduce the notations used throughout the proof, then provide a concise proof sketch, before finally elaborating the detailed proof.
Proof.
Notation. For clarity, we introduce the block matrix representation of $\mathbf{C}_{s_{t},t}-\mathbf{I}_{{n}_{x}}$ and $\mathbf{d}_{s_{t},t}$:
$\mathbf{C}_{s_{t},t}-\mathbf{I}_{{n}_{x}}\triangleq\left[\begin{array}[]{c|c}\mathbf{C}_{1}&\mathbf{C}_{2}\\ \hline\cr\mathbf{C}_{3}&\mathbf{C}_{4}\end{array}\right],\quad\mathbf{d}_{s_{t},t}\triangleq\left[\begin{array}[]{c}\mathbf{d}_{1}\\ \hline\cr\mathbf{d}_{2}\end{array}\right]$ (68)
where
$\mathbf{C}_{1}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}\times{{n}_{x}}_{\mathrm{\textnormal{ETI}}}},$ $\mathbf{C}_{2}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}\times{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}},$
$\mathbf{C}_{3}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}\times{{n}_{x}}_{\mathrm{\textnormal{ETI}}}},$ $\mathbf{C}_{4}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}\times{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}},$
$\mathbf{d}_{1}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}},$ $\mathbf{d}_{2}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}.$
We denote the projection of obstacle $\mathcal{O}_{i}$ and $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ onto the subspace $X_{\mathrm{\textnormal{ETI}}}$ by $\mathcal{O}_{\mathrm{\textnormal{ETI}}}\triangleq\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\mathcal{O}_{i}\right)$ and $\overline{\Omega}_{\mathrm{\textnormal{ETI}}}\triangleq\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\overline{\Omega}_{t}\right)$, respectively. Similarly, the projection of $\overline{\Omega}_{t}$ onto the subspace $X_{\mathrm{\textnormal{nonETI}}}$ is denoted as $\overline{\Omega}_{\mathrm{\textnormal{nonETI}}}$. We also denote $\mathcal{O}_{+}\triangleq\mathcal{O}_{\mathrm{\textnormal{ETI}}}\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}$ for brevity.
Proof sketch. We reinterpret the theorem as a set-inclusion problem between a true avoid set $P$ and over-approximated avoid set $\overline{\Lambda}_{i,t,t}$. Then, a conservative avoid set $Q$ is defined, from which we derive the sufficient condition for set-inclusion. Using the Extended-Translation-Invariance (ETI) assumption (Assumption 2) and Lemma 12, we prove the satisfaction of the sufficient condition.
Detailed Proof. We first define the true avoid set $P$, a set of states that satisfy the equation (29) as follows:
$\begin{split}P=\{\mathbf{x}\in\overline{\Omega}_{t}&\mid~{}\exists(\hat{\mathbf{x}},\gamma)\in\mathcal{O}_{i}\times[0,1]\quad\\ &\mathrm{\textnormal{s.t.}}\quad\mathbf{x}+\gamma(\mathbf{C}_{s_{t},t}\mathbf{x}+\mathbf{d}_{s_{t},t}-\mathbf{x})=\hat{\mathbf{x}}\}\end{split}$ (69)
Note that the theorem holds if and only if $P\subset\overline{\Lambda}_{i,t,t}$.
Next, to derive the sufficient condition for $P\subset\overline{\Lambda}_{i,t,t}$, we define a conservative avoid set $Q$:
$\begin{split}Q=\{\mathbf{x}_{\mathrm{\textnormal{ETI}}}&\in\overline{\Omega}_{\mathrm{\textnormal{ETI}}}\mid~{}\exists(\mathbf{x}_{\mathrm{\textnormal{nonETI}}},\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}},\gamma)\in\overline{\Omega}_{\mathrm{\textnormal{nonETI}}}\times\mathcal{O}_{\mathrm{\textnormal{ETI}}}\times[0,1]\quad\\ &\mathrm{\textnormal{s.t.}}\quad\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\gamma(\mathbf{C}_{1}\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\mathbf{x}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\}\end{split}$ (70)
Note that $\mathbf{x}=[\mathbf{x}_{\mathrm{\textnormal{ETI}}}^{\intercal},\mathbf{x}_{\mathrm{\textnormal{nonETI}}}^{\intercal}]^{\intercal}\in P$ implies $\mathbf{x}_{\mathrm{\textnormal{ETI}}}\in Q$, $\mathbf{x}_{\mathrm{\textnormal{nonETI}}}\in X_{\mathrm{\textnormal{nonETI}}}$, $\mathbf{x}\in\overline{\Omega}_{t}$, which leads to:
$P\subset(Q\times X_{\mathrm{\textnormal{nonETI}}})\cap\overline{\Omega}_{t}$ (71)
Building on this relation, we derive the sufficient condition for the theorem as follows:
$Q\subset\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right),\mathcal{O}_{\mathrm{\textnormal{ETI}}}\right)$ (72)
since satisfaction of (72) and (71) implies $P\subset\overline{\Lambda}_{i,t,t}$.
Now we prove the sufficient condition (72) using the ETI assumption. Fix $\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in Q$. Then by definition of $Q$, there exists $(\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}},\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}},\tilde{\gamma})\in\overline{\Omega}_{\mathrm{\textnormal{nonETI}}}\times\mathcal{O}_{\mathrm{\textnormal{ETI}}}\times[0,1]$ that satisfies:
$\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\tilde{\gamma}(\mathbf{C}_{1}\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}$ (73)
By rearranging the equation, we get:
$\begin{split}\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}&=\tilde{\gamma}\left(\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})\right)+(1-\tilde{\gamma})\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\\ &=\tilde{\gamma}\mathbf{y}+(1-\tilde{\gamma})\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\end{split}$ (74)
where noting $\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{\mathrm{\textnormal{ETI}}}$, it suffices to prove $\mathbf{y}\in\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)=\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\mathcal{B}\!\left(\mathcal{O}_{+},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)\right)$ to prove (72). This is shown by proving $[\mathbf{y}^{\intercal},\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}^{\intercal}]^{\intercal}\in\mathcal{B}\!\left(\mathcal{O}_{+},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$. Noting $\mathcal{O}_{+}=\mathcal{O}_{\mathrm{\textnormal{ETI}}}\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}$, we have to show
$\mathbf{y}+\mathbf{C}_{1}\mathbf{y}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}$ $\in\mathcal{O}_{\mathrm{\textnormal{ETI}}}$ (75)
$\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{C}_{3}\mathbf{y}+\mathbf{C}_{4}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{2}$ $\in\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}.$ (76)
Equation (76) is immediate, while (75) is true by
$\mathbf{y}+\mathbf{C}_{1}\mathbf{y}+$ $\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}$ (77)
$\begin{split}=\hat{\mathbf{x}}&{}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})\\ &+\mathbf{C}_{1}(\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}))\\ &+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}\\ \end{split}$ (78)
$=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}^{2}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}-\mathbf{C}_{1}\mathbf{d}_{1}$ (79)
$=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})$ (80)
$=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{\mathrm{\textnormal{ETI}}}$ (81)
The equation (81) follows from (80) using ETI Assumption 2 and the result of Lemma 12, namely:
$\begin{split}\mathbf{C}_{X_{\mathrm{\textnormal{ETI}}}}{(\mathbf{C}_{s_{t},t}\mathbf{x}+\mathbf{d}_{s_{t},t})}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}=\mathbf{C}_{1}(\mathbf{C}_{1}\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\mathbf{x}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})&=0\\ \forall\mathbf{x}&\in X\end{split}$ (82)
Since $\mathbf{y}\in\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)$ and $\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{\mathrm{\textnormal{ETI}}}$, by (74), $\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right),\mathcal{O}_{\mathrm{\textnormal{ETI}}}\right)$. Hence the sufficient condition (72) is proved, and the theorem is proved accordingly. ∎
-E Experiment Details
-E1 Near-Hover Quadrotor in 3-D
For the planning model, we employ the 3D single integrator, described as:
$\begin{split}\dot{p}_{x}&=\hat{v_{x}}\\ \dot{p}_{y}&=\hat{v_{y}}\\ \dot{p}_{z}&=\hat{v_{z}}\\ \end{split}$ (83)
where the planning states $\mathbf{p}=[p_{x},p_{y},p_{z}]^{\intercal}$ represents the position of the robot, and trajectory parameter $\mathbf{k}=[\hat{v_{x}},\hat{v_{y}},\hat{v_{z}}]^{\intercal}$ denotes the desired velocity of the robot. The trajectory parameter domain is chosen to be $K=[-0.5,0.5]^{3}$.
For the tracking model, we adopt a simplified 10-D quadrotor assuming near-hover conditions (i.e. small pitch and roll) [chen2021fastrack], expressed by:
$\begin{split}\dot{p}_{x}&=v_{x}\\ \dot{p}_{y}&=v_{y}\\ \dot{p}_{z}&=v_{z}\\ \dot{v}_{x}&=g\tan(\theta_{x})\\ \dot{v}_{y}&=g\tan(\theta_{y})\\ \dot{v}_{z}&=k_{T}\alpha_{z}-g\\ \dot{\theta}_{x}&=-d_{1}\theta_{x}+\omega_{x}\\ \dot{\theta}_{y}&=-d_{1}\theta_{y}+\omega_{y}\\ \dot{\omega}_{x}&=-d_{0}\theta_{x}+n_{0}\alpha_{x}\\ \dot{\omega}_{y}&=-d_{0}\theta_{y}+n_{0}\alpha_{y}\end{split}$ (84)
where the states $(p_{x},p_{y},p_{z})$ denote the position, $(v_{x},v_{y},v_{z})$ are velocities, $(\theta_{x},\theta_{y})$ are pitch and roll, and $(\omega_{x},\omega_{y})$ are pitch and roll rates. The control input $\mathbf{u}=(\alpha_{x},\alpha_{y},\alpha_{z})$ is the desired pitch and roll $(\alpha_{x},\alpha_{y})$ and the vertical thrust $\alpha_{z}$. The model parameters $(d_{0},d_{1},n_{0},k_{T})$ and control input bound $U$ are referenced from FaSTrack [chen2021fastrack]. A simple Linear Quadratic Regulator (LQR) linearized around the tracking error $e=0$ is employed for the feedback controller.
-E2 General Quadrotor in 3-D
We consider the time-switched polynomial planning model defined in [kousik2019safe, mueller2015computationally], where, for $0\leq t<t_{\mathrm{\textnormal{pk}}}$,
$\dot{p}_{x}=\frac{c_{1}(k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x})}{6}t^{3}+\frac{c_{2}(k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x})}{2}t^{2}+k_{ax}t+k_{vx},$ (85a)
$\dot{p}_{y}=\frac{c_{1}(k_{vy},k_{ay},k_{\mathrm{\textnormal{pk}}y})}{6}t^{3}+\frac{c_{2}(k_{vy},k_{ay},k_{\mathrm{\textnormal{pk}}y})}{2}t^{2}+k_{ay}t+k_{vy},$ (85b)
$\dot{p}_{z}=\frac{c_{1}(k_{vz},k_{az},k_{\mathrm{\textnormal{pk}}z})}{6}t^{3}+\frac{c_{2}(k_{vz},k_{az},k_{\mathrm{\textnormal{pk}}z})}{2}t^{2}+k_{az}t+k_{vz},$ (85c)
$c_{1}(k_{v},k_{a},k_{pk})=\frac{12}{t_{\mathrm{\textnormal{pk}}}^{3}}k_{v}+\frac{6}{t_{\mathrm{\textnormal{pk}}}^{2}}k_{a}-\frac{12}{t_{\mathrm{\textnormal{pk}}}^{3}}k_{pk},$ (85d)
$c_{2}(k_{v},k_{a},k_{pk})=-\frac{6}{t_{\mathrm{\textnormal{pk}}}^{2}}k_{v}-\frac{4}{t_{\mathrm{\textnormal{pk}}}}k_{a}+\frac{6}{t_{\mathrm{\textnormal{pk}}}^{2}}k_{pk},$ (85e)
and for $t_{\mathrm{\textnormal{pk}}}\leq t\leq t_{\mathrm{\textnormal{f}}}$,
$\dot{p}_{x}$ $=\frac{c_{3}(k_{\mathrm{\textnormal{pk}}x})}{6}(t-t_{\mathrm{\textnormal{pk}}})^{3}+\frac{c_{4}(k_{\mathrm{\textnormal{pk}}x})}{2}(t-t_{\mathrm{\textnormal{pk}}})^{2}+k_{\mathrm{\textnormal{pk}}x},$ (86a)
$\dot{p}_{y}$ $=\frac{c_{3}(k_{\mathrm{\textnormal{pk}}y})}{6}(t-t_{\mathrm{\textnormal{pk}}})^{3}+\frac{c_{4}(k_{\mathrm{\textnormal{pk}}y})}{2}(t-t_{\mathrm{\textnormal{pk}}})^{2}+k_{\mathrm{\textnormal{pk}}y},$ (86b)
$\dot{p}_{z}$ $=\frac{c_{3}(k_{\mathrm{\textnormal{pk}}z})}{6}(t-t_{\mathrm{\textnormal{pk}}})^{3}+\frac{c_{4}(k_{\mathrm{\textnormal{pk}}z})}{2}(t-t_{\mathrm{\textnormal{pk}}})^{2}+k_{\mathrm{\textnormal{pk}}z},$ (86c)
$c_{3}(k_{pk})$ $=\frac{12}{(t_{\mathrm{\textnormal{f}}}-t_{\mathrm{\textnormal{pk}}})^{3}}k_{pk},$ (86d)
$c_{4}(k_{pk})$ $=\frac{-6}{(t_{\mathrm{\textnormal{f}}}-t_{\mathrm{\textnormal{pk}}})^{2}}k_{pk},$ (86e)
where $t_{\mathrm{\textnormal{f}}}=3$, $t_{\mathrm{\textnormal{pk}}}=1$, the planning states and workspace $\mathbf{p}=\mathbf{w}=[p_{x},p_{y},p_{z}]^{\intercal}$ that represents the position of the quadrotor, and trajectory parameters $\mathbf{k}=[k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x},k_{vy},k_{ay},k_{\mathrm{\textnormal{pk}}y},k_{vz},k_{az},k_{\mathrm{\textnormal{pk}}z}]^{\intercal}$, where $k_{vx},k_{vy},k_{vz}$ represents the desired initial speed, $k_{ax},k_{ay},k_{az}$ represents the desired initial acceleration, and $k_{\mathrm{\textnormal{pk}}x},k_{\mathrm{\textnormal{pk}}y},k_{\mathrm{\textnormal{pk}}z}$ represents the desired speed at $t=t_{\mathrm{\textnormal{pk}}}$.
This planning model can be separated into three low-dimensional models for each of the three workspace dimensions. That is, we can consider the planning states and workspace $\mathbf{p}_{x}=\mathbf{w}_{x}=p_{x}$ and trajectory parameters $\mathbf{k}_{x}=[k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x}]$, defined by the dynamics (85a) and (86a), compute the reach set $\Omega_{0,x}\subset\mathbb{R}^{4}$ and avoid set ${\Lambda}_{i,t,0,x}\subset\mathbb{R}^{4}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and each $i=1,\cdots,{n}_{\mathcal{O}}$ using PARC, and repeat for $p_{y}$ and $p_{z}$.
We recover the 12-D reach set $\Omega_{0}$ and avoid set $\Lambda$ by:
$\Omega_{0}$ $=\Omega_{0,x}\times\Omega_{0,y}\times\Omega_{0,z},$ (87)
$\Lambda$ $=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\bigcup_{t=0}^{t_{\mathrm{\textnormal{f}}}-\Delta t}{\Lambda}_{i,t,0,x}\times{\Lambda}_{i,t,0,y}\times{\Lambda}_{i,t,0,z}.$ (88)
Further, we note that for each of the low-dimensional model, the intermediate avoid sets without tracking error $\overline{\Lambda}_{i,t,t,x}$, $\overline{\Lambda}_{i,t,t,y}$, and $\overline{\Lambda}_{i,t,t,z}$ and hence the avoid set $\Lambda$ can be computed without using convex hull or projection according to the following theorem:
Theorem 23 (Intermediate Avoid Set w/o Convex Hull or Projection).
Consider a planning model where the augmented planning state $\mathbf{x}(t)$ is:
$\mathbf{x}(t)=\begin{bmatrix}\mathbf{w}(t)\\ \mathbf{k}\end{bmatrix},$ (89)
where $\mathbf{w}(t)\in W\subset\mathbb{R}$ and $\mathbf{k}=[k_{1},\cdots,k_{{n}_{k}}]\in K\subset\mathbb{R}^{{n}_{k}}$. If
$\mathrm{\textnormal{slicedTraj}}\!\left(\mathcal{O}_{i},A,\mathbf{k}^{*}\right)=$ $\mathrm{\textnormal{slice}}_{2:({n}_{k}+1)}\!\left(\mathcal{O}_{i},\mathbf{k}^{*}\right)\cup$
$\mathrm{\textnormal{slice}}_{2:({n}_{k}+1)}\!\left(A,\mathbf{k}^{*}\right)$ (90)
is connected for all $\mathbf{k}^{*}\in\{k_{1,\mathrm{\textnormal{min}}},k_{1,\mathrm{\textnormal{max}}}\}\times\cdots\times\{k_{{n}_{k},\mathrm{\textnormal{min}}},k_{{n}_{k},\mathrm{\textnormal{max}}}\}$, where
$k_{i,\mathrm{\textnormal{min}}}$ $=\min_{\mathbf{k}}\{(\mathbf{k})_{i:i}\ |\ \mathbf{k}\in K\},$ (91)
$k_{i,\mathrm{\textnormal{max}}}$ $=\max_{\mathbf{k}}\{(\mathbf{k})_{i:i}\ |\ \mathbf{k}\in K\},$ (92)
for $i=1,\cdots,{n}_{k}$, then
$\overline{\Lambda}_{i,t,t}=(\mathcal{O}_{i}\cup A)\cap\overline{\Omega}_{t},$ (93)
can replace (29) to fulfill the condition in (30).
Proof.
We follow the proof strategy outlined in Appendix -D. Since all states in $\mathbf{x}(t)$ are ETI, $\mathbf{w}(t)$ is 1-D, and all $\mathbf{k}\in K$ are constant, Theorem 23 is proven if $\mathbf{y}\in A$ and $\hat{\mathbf{x}}=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{i}$ implies $\tilde{\mathbf{x}}=\tilde{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{i}\cup A$ by (74). This is true if $\mathrm{\textnormal{slicedTraj}}\!\left(\mathcal{O}_{i},A,\mathbf{k}\right)$ are connected for all $\mathbf{k}\in K$. But if (23) is connected for all $\mathbf{k}^{*}$, by convexity of $\mathcal{O}_{i}$ and $A$, $\mathrm{\textnormal{slicedTraj}}\!\left(\mathcal{O}_{i},A,\mathbf{k}\right)$ must be connected for all $\mathbf{k}\in K$, since $[k_{1,\mathrm{\textnormal{min}}},\cdots,k_{{n}_{k},\mathrm{\textnormal{min}}}]^{\intercal}\leq\mathbf{k}\leq[k_{1,\mathrm{\textnormal{max}}},\cdots,k_{{n}_{k},\mathrm{\textnormal{max}}}]^{\intercal}\ \forall\mathbf{k}\in K$. ∎
The tracking model for the system is from [kousik2019safe, lee2010control], defined as:
$\begin{bmatrix}\dot{p}_{x}\\ \dot{p}_{y}\\ \dot{p}_{z}\end{bmatrix}$ $=\mathbf{v},$ (94a)
$\dot{\mathbf{v}}$ $=\tau\mathbf{R}\mathbf{e}_{3}-mg\mathbf{e}_{3},$ (94b)
$\dot{\bm{\omega}}$ $=\mathbf{J}^{-1}\left(\begin{bmatrix}\mu_{x}\\ \mu_{y}\\ \mu_{z}\end{bmatrix}-\bm{\omega}\times\mathbf{J}\bm{\omega}\right),$ (94c)
$\dot{\mathbf{R}}$ $=\mathbf{R}\ \mathrm{\textnormal{hat}}(\bm{\omega}),$ (94d)
where $\mathbf{v}\in\mathbb{R}^{3}$ is velocity, $\mathbf{R}\in\mathrm{\textnormal{SO}}(3)$ is the attitude, $\mathbf{e}_{3}\in\mathbb{R}^{3}$ is the unit vector in the inertial frame that points “up” relative to the ground, $m=0.547$, $g=9.81$, $\bm{\omega}\in\mathbb{R}^{3}$ is the angular velocity, $\mathbf{J}=\mathrm{\textnormal{diag}}(0.0033,0.0033,0.0058)$, $\mathrm{\textnormal{hat}}:\mathbb{R}^{3}\to\mathrm{\textnormal{SO}}(3)$ is the hat map operator [lee2010control], and the control inputs $\mathbf{u}=[\tau,\mu_{x},\mu_{y},\mu_{z}]^{\intercal}$ are net thrust and body moments.
-E3 Comparison Methods
FaSTrack
FaSTrack is a planner-tracker framework that precomputes a uniform Tracking Error Bound (TEB) (maximum possible tracking error) using Hamilton-Jacobi (HJ) reachability analysis. Obstacles are padded by the TEB; then, as long as the planner only plans outside the padded obstacles the robot is safe. During the TEB computation, FaSTrack also synthesizes a controller to guarantee the bound. We use RRT as the planning model as per [chen2021fastrack].
Neural CLBF
Neural CLBF approximates a safety filter value function as a neural network, thereby segmenting the environment into goal, safe, and unsafe regions. The value function, trained with data from the entire state space, enables the generation of safe, goal-directed control inputs through Quadratic Programming (QP).
Reachability-based Trajectory Design
RTD is a planner-tracker framework that is very similar to PARC in that it uses a planner-tracker framework and compensates for tracking errors in a time-varying way. However, RTD leverages forward reachable sets (FRS). RTD numerically approximates the FRS either with sums-of-squares [kousik2020bridging] or zonotope [kousik2019safe, althoff2015introduction] reachability. At runtime, RTD performs receding-horizon trajectory planning with nonlinear, nonconvex constraints derived from the FRS intersecting with obstacles to eliminate unsafe motion plans. Each parameterized plan contains a failsafe maneuver such that, if replanning fails, the previous (safe) maneuver can be executed.
-F Additional Experiment Details
-F1 Impact of Planner-Tracker Cooperation
Environment
The environment is a $[0,10]\times[-10,10]\times[0,10]$ cuboid (all lengths are in meters). The goal $\overline{P_{\mathrm{\textnormal{G}}}}$ is a cube centered at $(p_{x},p_{y},p_{z})=(9,0,5)$ with a side of $1$unitm. Obstacles $\overline{\mathcal{O}}$ include the floor, ceiling, and walls, and two additional cuboid obstacles of $[6.5,7.5]\times[-8.2,-\Delta w/2]\times[0,10]$ and $[6.5,7.5]\times[\Delta w/2,8.2]\times[0,10]$ with varying $\Delta w$ that denotes the gap between two obstacles. These are shown in Fig. 7 with a gap width $\Delta w=1.8$. The obstacles are padded to accommodate the quadrotor’s body, approximated by a cuboid of $0.54\times 0.54\times 0.05$ using the specifications of Hummingbird [dong2015development]. The planning horizon is $t_{\mathrm{\textnormal{f}}}=10$units with time discretization $\Delta t=0.1$units each. The frequency of the tracking controller is $1000$unitHz for all methods.
-F2 Impact of Tighter Planning Reachability
Environment
The general 3-D quadrotor environment is similar to that of the 10-D near-hover quadrotor environment but with a narrower gap to the goal. The obstacles are defined as $\overline{\mathcal{O}}_{1}=[3.23,6.77]\times[0.23,9.27]\times[0.73,9.27]$, $\overline{\mathcal{O}}_{2}=[3.23,6.77]\times[-9.27,-0.23]\times[0.73,9.27]$ (after accounting for an overapproximated volume of the quadrotor), that is, two walls with a clearance of 0.46, the goal region is defined as $\overline{P_{\mathrm{\textnormal{G}}}}=[7.44,9.56]\times[-1.06,1.06]\times[3.94,6.06]$, and the domains are defined as $P=[0,10]\times[-10,10]\times[0,10]$ and $K=[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]\times[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]\times[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]$. The quadrotor starts stationary at $p_{x}=1,p_{y}=0,p_{z}=5$. This means the quadrotor begins at the side of the wall opposite to the goal region, but the $p_{y}$ coordinate is at the middle of the gap.
-F3 Impact of Proper Use of Time-varying Tracking Error
Environment
In the 3-D quadrotor environment, similar to the one described in Section V-C, we evaluate two scenarios differentiated by obstacle gap widths. The obstacles are specified for both scenarios as $\overline{\mathcal{O}}_{1}=[5.75,7.25]\times[\triangle w/2,10]\times[1,9]$ and $\overline{\mathcal{O}}_{2}=[5.75,7.25]\times[-10,-\triangle w/2]\times[1,9]$, using $\triangle w=0.85$ for the narrow gap and $\triangle w=3.0$ for the wide gap. The goal region is set as $\overline{P_{\mathrm{\textnormal{G}}}}=[7,9]\times[-1,1]\times[4,6]$, with planning domains defined by $P=[0,10]\times[-10,10]\times[0,10]$. The trajectory parameter space, $K=[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]\times[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]\times[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]$, was selected to simplify the tuning of a funnel-based feedback controller. Initial states for simulations are sampled from a $(25,25,3)$ grid spanning $[0.1,4.8]\times[-9.9,9.9]\times[3,7]$. Due to uniform obstacle distribution in the $p_{z}$ direction, sampling is denser in the $p_{x},p_{y}$ plane, since we expect varying initial states in $p_{z}$ direction would not be a significant influence. Each simulation starts with the quadrotor stationary, following reference trajectories determined by each methodology.
Implementation of Comparison Methods
In this section, we detail the implementation of KDF [verginis2022kdf], and the tracking controller we used to compare PARC and KDF. The originally proposed controller from KDF [verginis2022kdf] was unsuitable for under-actuated quadrotors; thus we adopted the Prescribed Performance Controller (PPC) from [lapandic2022robust]. This PPC ensures that position and yaw angle tracking errors stay within prescribed performance funnels, allowing it to be compatible with the KDF framework where it uses maximal tracking error in motion planning level to define extended free space.
For PARC, we successfully tuned 48 hyper-parameters and defined 12 funnel functions for each quadrotor state, with corresponding control gains as follows:
  • $\rho_{p_{i}}(t)=0.1+(0.5-0.1)e^{-0.5t}$, $i=\{x,y,z\}$,
  • $\rho_{v_{x}}(t)=0.5+(5-0.5)e^{-0.2t}$, $\rho_{v_{y}}(t)=0.5+(5-0.5)e^{-0.2t}$, $\rho_{v_{z}}(t)=0.2+(5-0.2)e^{-1.5t}$
  • $\rho_{\phi\theta_{1}}(t)=0.1+(0.5-0.1)e^{-0.5t}$, $\rho_{\phi\theta_{2}}(t)=0.1+(0.5-0.1)e^{-0.5t}$, $\rho_{\psi}(t)=0.1+(0.4-0.1)e^{-0.05t}$
  • $\rho_{w_{i}}(t)=0.3+(0.5-0.3)e^{-0.5t}$, $i=\{x,y,z\}$
  • $K_{p}=\mathrm{\textnormal{diag}}\!\left(1.25,1.25,12.5\right)$
  • $K_{v}=\mathrm{\textnormal{diag}}\!\left(20,20,5\right)$
  • $K_{\phi\theta}=\mathrm{\textnormal{diag}}\!\left(3,1.5\right)$, $K_{\psi}=1$,
  • $K_{w}=10\cdot\mathbf{I}_{3}$.
We refer to [lapandic2022robust] for details of the role of each hyper-parameter. This hyper-parameter effectively tracked 2,400 reference trajectories from PARC in the wide gap scenario and all computed reach-avoid plans in the narrow gap scenario, as detailed in Table II and Fig. 15.
However, despite extensive efforts, we could not find a universal hyper-parameter that effectively tracked KDF’s trajectories, which may be due to RRT-generated adversarial trajectories being difficult for the controller to follow. To address this, we used smoothing splines to fit the RRT plans to adhere to the twice-differentiability required by funnel controllers but still encountered challenges with universal hyper-parameter settings. Therefore, assuming the existence of an effective funnel controller for KDF, we evaluated its performance based on planned trajectories, and padding obstacles by $\rho_{p}(0)$ as per the MATLAB Navigation Toolbox’s basic RRT implementation. We also present a selection of successfully tracked trajectories by KDF in Fig. 11 to validate our implementation of KDF.
Figure 15: The plot of realized trajectory in narrow-gap scenario, clearance of $0.31m$ for PARC. The 306 reach-avoid plans were computed out of 3,750 initial states.
-G Additional Experiments
We provide a further comparison to the experiment in V-A by using a different planning model in the same environmental setting. The experiment seeks to show how PARC could improve using a better planning model under the same experimental conditions.
-G1 Planning and Tracking Model
We consider the time-switched polynomial model defined by (85a), (85c), (86a), and (86c), with $t_{\mathrm{\textnormal{f}}}=2$, $t_{\mathrm{\textnormal{pk}}}=1.5$, the planning states and workspace $\mathbf{p}=\mathbf{w}=[p_{x},p_{z}]^{\intercal}$, and trajectory parameters $\mathbf{k}=[k_{vx},k_{ax},k_{\mathrm{\textnormal{pk}}x},k_{vz},k_{az},k_{\mathrm{\textnormal{pk}}z}]^{\intercal}$. Just like in Section -E2, we separated the planning model into low-dimensional models for each of the two workspace dimensions, and recovered the high-dimensional reach set, avoid set, and BRAS by (87) and (88).
We used the same tracking model defined in (54) for a fair comparison.
-G2 Experimental Setup
Our experimental setup for the obstacles, goal region, and workspace domain are identical to that in Section V-A and [dawson2022safe]. We also use the same iLQR controlling scheme as in Section V-A but with the gains tuned to the new planning model. We define the domain for trajectory parameters with $K=[-1.5,1.5]\times\{0\}\times[-0.1,0.1]\times\{0\}\times[-1,1]\times[-1.5,1.5]$.
-G3 Hypothesis
We anticipate PARC to be able to compute a less conservative, but still guaranteed safe BRAS compared to Section V-A due to the reduced number (three vs. zero) of non-ETI states in the planning model. Moreover, we also expect PARC’s computation time to be faster than the naïve model choice, as the dimension for polytopic computation has been reduced from 8 to 4.
Since the planning model in Appendix -G1 also incorporates stabilizing behavior for the quadrotor near the goal region [kousik2019safe, mueller2015computationally], we expect to produce similar reach-avoid results to [dawson2022safe], that is, the quadrotor would be able to safely navigate between the two obstacles while stopping at the goal region.
-G4 Results
[…]
Trajectories were generated by uniformly sampling the BRAS. No crashes were reported, with the agent successfully stopping at the goal region in all simulations. Comparing Fig. 16 with Fig. 6, safe trajectories that travel in-between two obstacles can be generated, reproducing similar results to [dawson2022safe].
-G5 Discussion
The time-switched polynomial planning model has comparable tracking error with the naïve planing model in Section V-A. Thus, the reduced conservativness in the computed BRAS for the polynomial model is largely due to the reduced number of non-ETI states in the PWA system, leading to a tighter approximation of the avoid sets. Note that the BRAS in Fig. 6 extends more to the right compared with Fig. 16 because the naïve planning model does not encode any stabilizing behavior, with the trajectory leading the agent to continuously accelerating towards the goal. Thus, within the same $t_{\mathrm{\textnormal{f}}}$, it was able to reach the goal from a further distance compared with the stabilizing trajectories.
We further note that, while the chosen domain for the time-switched polynomial planning model was not able to generate very parabolic trajectories that enable the agent to reach the goal from behind the left obstacle, this can be easily fixed by receding-horizon planning into the BRAS, or by performing PARC across multiple domains of trajectory parameters, both of which remains as future work.
In conclusion, these experiments show the importance of choosing a good planning model for PARC with minimal dimensions and non-ETI states without losing representation power in the trajectories. Should good planning models be unavailable, Section VI shows how one can construct planning models from data while using PARC to compute their corresponding BRAS.
Figure 16: Results of PARC on a time-switched polynomial planning model with stabilizing end-behavior, after accounting for tracking error. Green indicates the goal region, red indicates the obstacles with the volume of the drone accounted for, and blue shows the computed BRAS. Safe trajectories that pass between the two obstacles can now be generated compared to the results in Section V-A.
-H Drift Vehicle Demo Details
We now define our tracking model, drift parking controller, and the time-variant affine planning model. We then present results and a brief discussion.
-H1 Tracking Model
We use a tracking model based on [jelavic2017autonomous, goh2020toward, weber2023modeling, goh2016simultaneous]:
$\mathbf{z}=\begin{bmatrix}p_{x}&p_{y}&\theta&\omega&v&\beta\end{bmatrix}^{\intercal}$ (95)
with its dynamics defined as:
$\begin{split}\dot{p}_{x}&=v\cos(\theta+\beta)\\ \dot{p}_{y}&=v\sin(\theta+\beta)\\ \dot{\theta}&=\omega\\ \dot{\omega}&=\frac{1}{I_{z}}(L_{a}{f_{y,\mathrm{\textnormal{f}}}}\cos(\delta)+L_{a}{f_{x,\mathrm{\textnormal{f}}}}\sin(\delta)-L_{b}{f_{y,\mathrm{\textnormal{r}}}})\\ \dot{v}&=\frac{1}{m}({f_{x,\mathrm{\textnormal{f}}}}\cos(\delta-\beta)-{f_{y,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)\\ &\qquad+{f_{x,\mathrm{\textnormal{r}}}}\cos(\beta)+{f_{y,\mathrm{\textnormal{r}}}}\sin(\beta))\\ \dot{\beta}&=\frac{1}{mv}({f_{x,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)-{f_{y,\mathrm{\textnormal{f}}}}\sin(\delta-\beta)\\ &\qquad+{f_{x,\mathrm{\textnormal{r}}}}\cos(\beta)+{f_{y,\mathrm{\textnormal{r}}}}\sin(\beta))-\omega\end{split}$ (96)
[…]
The lateral forces on both the front and the rear tires,
${f_{y,\mathrm{\textnormal{f}}}}$ $=f_{\text{front}}(\omega,v,\beta,{f_{x,\mathrm{\textnormal{f}}}},\delta)\quad\mathrm{\textnormal{and}}$ (97)
${f_{y,\mathrm{\textnormal{r}}}}$ $=f_{\text{rear}}(\omega,v,\beta,{f_{x,\mathrm{\textnormal{r}}}}),$ (98)
are formulated using the Fiala brush tire model [fiala1954lateral]. Lastly, the car’s physical parameters ($I_{z},m,L_{a},L_{b}$) as well as other constants used in the Fiala tire model are adopted from the DeLorean car specs from [goh2016simultaneous].
-H2 Feedback Controller
Literature in drift parking controllers falls into two categories. [lau2011learning, leng2023deep] approached this problem with learning-based controllers. [kolter2010probabilistic, jelavic2017autonomous] designed mixed open and closed-loop control strategies with the former using the probabilistic method and the latter deterministic.
In this demo, we adopt the control scheme described in [jelavic2017autonomous]. We first used a nonlinear MPC tracking controller to bring the car into a drift state then switched to an open-loop control scheme to complete the drift parking action.
[…]
-H3 Planning Model
Different from the experiment examples, the mixed open and closed loop drifting controller cannot track explicit trajectories when the car enters the drifting regime. Therefore, the construction of an analytical model that describes the entire drift parking trajectory is difficult or even impossible just like many robotics applications [nguyen2011model].
In this demo, we showcase the flexibility of PARC by constructing the planning model using data. We are directly constructing the planning model as a time-variant affine system with 3-D planning space and 2-D trajectory parameters, so PARC can be directly applied without Section IV-A. Specifically, we wish to design the affine dynamics in the form of:
$\begin{bmatrix}p_{x,t+\Delta t}\\ p_{y,t+\Delta t}\\ \theta_{t+\Delta t}\\ {v}_{t+\Delta t}\\ {\beta}_{t+\Delta t}\end{bmatrix}=\begin{bmatrix}1&0&0&c_{vx,t}&c_{\beta x,t}\\ 0&1&0&c_{vy,t}&c_{\beta y,t}\\ 0&0&1&c_{v\theta,t}&c_{\beta\theta,t}\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}\begin{bmatrix}p_{x,t}\\ p_{y,t}\\ \theta_{t}\\ {v}_{t}\\ {\beta}_{t}\end{bmatrix}+\begin{bmatrix}d_{x,t}\\ d_{y,t}\\ d_{\theta,t}\\ 0\\ 0\end{bmatrix}$ (99)
where the planning states $\mathbf{p}=[p_{x},p_{y},\theta]^{\intercal}$ are the world-frame coordinates and heading angle of the car, $\mathbf{w}=[p_{x},p_{y}]^{\intercal}$ are the workspace states, trajectory parameters $\mathbf{k}=[v,\beta]^{\intercal}$ are the drifting velocity and the side-slip angle for controller switching, the domain for all timesteps of the PWA function is $P\times K$, and we wish to design $c_{vx,t},c_{\beta x,t},d_{x,t},c_{vy,t},c_{\beta y,t},d_{y,t},c_{v\theta,t},c_{\beta\theta,t},$ and $d_{\theta,t}$ for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ to fit (99) to the collected data.
For this example, as we mentioned in Remark 6, we are extending the dimensions of the goal region to all of $\mathbf{p}$ instead of just $\mathbf{w}$, so $\overline{\mathcal{O}}\subset P$ and $\mathcal{O}_{i}=\overline{\mathcal{O}}_{i}\times K$ for all $i=1,\cdots,{n}_{\mathcal{O}}.$ This is such that we can enforce the $\theta$ of the car at $t_{\mathrm{\textnormal{f}}}$ so the parking is “parallel”. The extension of the goal dimension does not change the other parts of PARC’s formulation.
Now we design the parameters in (99). Unlike our other system models, due to the lack of controllability during drifting, the behavior of the realized drift parking trajectories are “open-loop” in that they are uniquely and entirely determined given $\mathbf{z}_{0}$ and $\mathbf{k}$ without needing to provide the planning model, as illustrated in Section -H2. As such, we fit a planning model to the data collected in open-loop simulations and use PARC to compute the BRAS of the initial states.
We first collect data by sampling from the pairs $(\mathbf{z}_{0,i},\mathbf{k}_{i})\in Z_{\mathrm{\textnormal{data}}}\times K_{\mathrm{\textnormal{data}}}\subset Z\times K$ for $i=1,\cdots,{n}_{\mathrm{\textnormal{data}}}$. For the $i^{\mathrm{\textnormal{th}}}$ data, we denote the realized trajectories $\mathbf{z}_{i}(t;\mathbf{z}_{0,i},\mathbf{k}_{i})=[{p_{x}}_{i}(t),{p_{y}}_{i}(t),{\theta}_{i}(t),\cdots]^{\intercal}$ and the sampled trajectory parameters $\mathbf{k}_{i}=[{v}_{i},{\beta}_{i}]^{\intercal}$. Then, we can fit the parameters in (99) by solving the least squares problems:
$\min_{c_{vx,t},c_{\beta x,t},d_{x,t}}$ $\Bigg{\|}\begin{bmatrix}{v}_{1}&{\beta}_{1}&1\\ \vdots&\vdots&\vdots\\ {v}_{{n}_{\mathrm{\textnormal{data}}}}&{\beta}_{{n}_{\mathrm{\textnormal{data}}}}&1\end{bmatrix}\begin{bmatrix}c_{vx,t}\\ c_{\beta x,t}\\ d_{x,t}\end{bmatrix}$
$-\begin{bmatrix}{p_{x}}_{1}(t+\Delta t)-{p_{x}}_{1}(t)\\ \vdots\\ {p_{x}}_{{n}_{\mathrm{\textnormal{data}}}}(t+\Delta t)-{p_{x}}_{{n}_{\mathrm{\textnormal{data}}}}(t)\end{bmatrix}\Bigg{\|}^{2}_{2},$ (100a)
$\min_{c_{vy,t},c_{\beta y,t},d_{y,t}}$ $\Bigg{\|}\begin{bmatrix}{v}_{1}&{\beta}_{1}&1\\ \vdots&\vdots&\vdots\\ {v}_{{n}_{\mathrm{\textnormal{data}}}}&{\beta}_{{n}_{\mathrm{\textnormal{data}}}}&1\end{bmatrix}\begin{bmatrix}c_{vy,t}\\ c_{\beta y,t}\\ d_{y,t}\end{bmatrix}$
$-\begin{bmatrix}{p_{y}}_{1}(t+\Delta t)-{p_{y}}_{1}(t)\\ \vdots\\ {p_{y}}_{{n}_{\mathrm{\textnormal{data}}}}(t+\Delta t)-{p_{y}}_{{n}_{\mathrm{\textnormal{data}}}}(t)\end{bmatrix}\Bigg{\|}^{2}_{2},$ (100b)
$\min_{c_{v\theta,t},c_{\beta\theta,t},d_{\theta,t}}$ $\Bigg{\|}\begin{bmatrix}{v}_{1}&{\beta}_{1}&1\\ \vdots&\vdots&\vdots\\ {v}_{{n}_{\mathrm{\textnormal{data}}}}&{\beta}_{{n}_{\mathrm{\textnormal{data}}}}&1\end{bmatrix}\begin{bmatrix}c_{v\theta,t}\\ c_{\beta\theta,t}\\ d_{\theta,t}\end{bmatrix}$
$-\begin{bmatrix}{\theta}_{1}(t+\Delta t)-{\theta}_{1}(t)\\ \vdots\\ {\theta}_{{n}_{\mathrm{\textnormal{data}}}}(t+\Delta t)-{\theta}_{{n}_{\mathrm{\textnormal{data}}}}(t)\end{bmatrix}\Bigg{\|}^{2}_{2},$ (100c)
repeated for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$. (100) can be solved by any least squares solver. Once (100) is computed, (99) is now a fully defined time-variant affine function that can be used in the PARC framework.
Note that, since (99) is generated only from data with initial conditions in $Z_{\mathrm{\textnormal{data}}}\times K_{\mathrm{\textnormal{data}}}$, the BRAS is also only valid from within that region. Specifically, when sampling for tracking error à la Section IV-D, one should have $\tilde{Z}\times\tilde{K}\subset Z_{\mathrm{\textnormal{data}}}\times K_{\mathrm{\textnormal{data}}}$.
For the planning model, we set the timestep $\Delta t=0.1$units and final time $t_{\mathrm{\textnormal{f}}}=7.8$units.
-H4 Environment Details
In our environment, we define the tracking model state and input bounds to be:
$\mathbf{z}_{\mathrm{\textnormal{bound}}}=\begin{bmatrix}-5,&40\\ -20,&5\\ -1.5\pi,&0.5\pi\\ -2,&1\\ -0.5,&20\\ -\pi,&\pi\end{bmatrix}\quad\mathbf{u}_{\mathrm{\textnormal{bound}}}=\begin{bmatrix}0,&7684\\ -0.6632,&0.6632\\ -5163,&5163\end{bmatrix}$ (101)
We use the geometry specs of the 1981 DeLorean ($l=4.267,w=1.988$unitm) [goh2020toward, goh2016simultaneous] to represent the car agent as a 2D rectangular H-Polytope $\overline{B}$. The obstacles ${\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{1},{\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{2}$ are two parked car obstacles centered at $[24,-16,-\pi]^{\intercal}$ and $[37,-16,-\pi]^{\intercal}$ each with the same geometry as $\overline{B}$.
Next, to ensure that the agent doesn’t collide with the obstacles, we create a circular over approximation of $\overline{B}$ as $B$ with a radius of $r=\sqrt{l^{2}+w^{2}}/2$. We then dilate the obstacle into:
${\overline{\mathcal{O}}_{1}}$ $={\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{1}\oplus B$
${\overline{\mathcal{O}}_{2}}$ $={\overline{\mathcal{O}}_{\mathrm{\textnormal{body}}}}_{2}\oplus B$
so that as long as the geometrical center of the car agent $\mathbf{w}$ does not enter $\overline{\mathcal{O}}_{1}$ and $\overline{\mathcal{O}}_{2}$ as formulated in Problem 4, the trajectory is safe. Note that in this demo, the geometric center and the center of mass coincide.
In this demo, our desired state for the car includes position and orientation states. Therefore we expand the goal set dimension from the previously used workspace dimension $\mathbb{R}^{{n}_{w}}$ to the planning state dimension $\mathbb{R}^{{n}_{p}}$ by including a range of $\theta$ as a goal state. So, we slightly modify Problem 4’s goal reaching condition to
$\mathrm{\textnormal{proj}}_{1:{n}_{p}}\!\left(\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})\right)\in\overline{P_{\mathrm{\textnormal{G}}}}$
We place the center of the goal in between $\overline{\mathcal{O}}_{1}$ and $\overline{\mathcal{O}}_{2}$ at
$\mathbf{p}_{\mathrm{\textnormal{goal}}}=\begin{bmatrix}30.5&-16&-\pi\end{bmatrix}^{\intercal}$
The goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ is then defined as a 3D cuboid H-Polytope centered at $\mathbf{p}_{\mathrm{\textnormal{goal}}}$ with its length in $[p_{x},p_{y},\theta]$ as $[3.7,1.6,\pi/3]$.
Using the drift parking test environment, we collect 10,000 expert trajectories generated by feeding samples of $\mathbf{k}\in K=[9,11]\times[0.6109,0.7854]$ into the drift parking controller described above. The initial condition of all trajectories $\mathbf{z}_{0}$ is all zeros except for the velocity state $v_{0}=0.01$, as the tracking model experiences numerical instability at low velocities and sideslip angles. The entire simulation process lasts for 2 hours. With the expert trajectories ready, we create the time-variant affine planning model using the formulation above.
Figure 17: We are preparing an F1:10 class robotic vehicle to implement and deploy our safe drifting maneuvers. We aim to have a working demo by May 2024.